### v1

# Bayesian Concept Bottleneck Models with LLM Priors

Abstract
Concept Bottleneck Models (CBMs) have been proposed as a compromise between white-box and black-box models, aiming to achieve interpretability without sacrificing accuracy. The standard training procedure for CBMs is to predefine a candidate set of human-interpretable concepts, extract their values from the training data, and identify a sparse subset as inputs to a transparent prediction model. However, such approaches are often hampered by the tradeoff between enumerating a sufficiently large set of concepts to include those that are truly relevant versus controlling the cost of obtaining concept extractions. This work investigates a novel approach that sidesteps these challenges: BC-LLM iteratively searches over a potentially infinite set of concepts within a Bayesian framework, in which Large Language Models (LLMs) serve as both a concept extraction mechanism and prior. BC-LLM is broadly applicable and multi-modal. Despite imperfections in LLMs, we prove that BC-LLM can provide rigorous statistical inference and uncertainty quantification. In experiments, it outperforms comparator methods including black-box models, converges more rapidly towards relevant concepts and away from spuriously correlated ones, and is more robust to out-of-distribution samples.
1 Introduction
Although machine learning (ML) algorithms have demonstrated remarkable predictive performance, many lack the interpretability and transparency necessary for human experts to audit their accuracy, fairness, and safety (Dwork et al.,, 2012; Rudin,, 2018). This has limited their adoption in high-stakes applications such as medicine (Thirunavukarasu et al.,, 2023) and settings where regulatory agencies require algorithms to be explainable (Meskó and Topol,, 2023).
Recent works have explored Concept Bottleneck Models (CBMs) (Koh et al.,, 2020; Kim et al.,, 2023) as a potential solution: these methods leverage black-box algorithms to extract a small number of interpretable concepts, which are subsequently processed by a fully transparent tabular model to make predictions. This allows users to understand how exactly each concept contributes to the final prediction and, if necessary, modify the extracted concepts to fix model predictions.
However, there are major limitations in the way CBMs are currently trained. They require human experts to identify a large pool of concepts a priori and annotate the concepts present in each observation. This heavy dependence on human experts is expensive and often impractical. More importantly, it is difficult to ensure this pool of concepts includes the truly relevant concepts. Consequently, resulting CBMs are often outperformed by their black-box counterparts (Yuksekgonul et al.,, 2022) and may even mislead users as to which concepts are truly relevant (Ramaswamy et al.,, 2023).
Consider the task of predicting hospital readmission from patient notes. It is impossible to prespecify all possible related concepts. Not only are there many different concepts, but any single concept can also be refined. For instance, for the concept regarding the “patient’s smoking status,” there is the more general concept regarding “substance use,” the correlated concept of “whether a patient has quit smoking,” and many more. While iterative refinement of candidate concepts may help, it would make learning a CBM even more burdensome.
LLMs have the potential to address these challenges. They are cheaper than human experts, can provide (relatively) high-quality concept annotations, and have sufficient world knowledge to hypothesize useful concepts. However, applying LLMs as a simple stand-in for a human expert does not fully address the challenges of concept discovery. First, LLM queries, though cheaper than human experts, still have a non-zero cost, which practically limits the number of concepts one can explore. Second, searching over concepts using LLMs introduces a “garden of forking paths” and may lead to the selection of spuriously correlated concepts (Gelman and Loken,, 2013). Finally, LLMs are imperfect query engines that can hallucinate, suggest bad candidate concepts, provide incorrect concept annotations, and may not even be self-consistent in their prior beliefs (Wang et al.,, 2023). As such, current methods that rely on LLMs to learn CBMs have struggled to achieve good performance, particularly in areas where the LLM’s prior knowledge is not well-aligned with the truth (Ludan et al.,, 2023; Benara et al.,, 2024).
Figure 1: Overview of Bayesian inference for learning concept bottleneck models with LLM priors (BC-LLM).
To address the aforementioned challenges, we show how wrapping LLMs within a Bayesian framework allows one to iteratively explore concepts in a way that is statistically principled and provides uncertainty quantification. An LLM can serve multiple purposes in this framework: it defines a prior over a potentially infinite set of candidate concepts, proposes concepts to explore during posterior sampling, and annotates observations with the candidate concepts. The procedure, which we refer to as BC-LLM, is designed to be computationally efficient; at each iteration, the LLM proposes a batch of candidate concepts so their respective annotations for each observation can be batched into a single LLM query. We prove that asymptotically, BC-LLM converges to the correct concepts even when the LLM prior is inconsistent or incompatible. Moreover, BC-LLM is highly generalizable and works across multiple modalities (text, images, and tabular data). Experiments across multiple datasets show that BC-LLM outperforms comparator methods and often even black-box models, converges more rapidly towards the true concepts and away from spuriously correlated ones, and is more robust to out-of-distribution (OOD) samples. Finally, when BC-LLM assisted a real-world hospital’s data science team to revise an existing tabular ML model, clinicians found the resulting CBM to be substantially more interpretable and actionable than the prior model.
2 Related work
Prior CBM methods train on observations with hand-annotated (Koh et al.,, 2020; Kim et al.,, 2023; Yuksekgonul et al.,, 2023) and, more recently, LLM-annotated concepts (Patel et al.,, 2023; McInerney et al.,, 2023; Benara et al.,, 2024) for a set of prespecified concepts by human experts. These methods are known to be highly sensitive to the selected set of concepts (Ramaswamy et al.,, 2023). More recent works have suggested using LLMs to prespecify the list of concepts instead (Oikarinen et al.,, 2023; Yang et al.,, 2023), but this requires the LLM to have accurate prior knowledge about the supervised learning task.
Consequently, recent works have hypothesized that taking an iterative approach can more effectively find a set of relevant concepts (Ludan et al.,, 2023; Liu et al., 2024a, ). These methods employ a boosting-like approach, where the LLM is provided a set of misclassified examples and asked to add/revise concepts to improve performance. However, the LLM can typically only review a few observations due to limits on context length, so the resulting concept proposals are far from optimal.
An alternative strategy is to conduct post-hoc distillation of a black-box model into interpretable concepts (Kim et al.,, 2018; Yeh et al.,, 2020; Zhao et al.,, 2024; Jiang et al.,, 2023), which can be used to learn an interpretable prediction model. Nevertheless, prior work has shown that the generated explanations are often not faithful to the black-box model, leading to poorer performance from the interpretable model (Doshi-Velez and Kim,, 2017).
Another related line of work uses LLMs to describe conceptual differences between pairs of datasets  (Zhong et al.,, 2023, 2022; Dunlap et al.,, 2024). However, existing methods are primarily suited for simple classification tasks, as they essentially learn a CBM with a single concept rather than a set of concepts.
Finally, recent works have considered combining LLMs with Bayesian techniques (Yang et al.,, 2024; Liu et al., 2024b, ; Liu et al., 2024c, ). Some have suggested that In-Context Learning (ICL) can be viewed as LLMs conducting Bayesian inference (Xie et al.,, 2021; Ye et al.,, 2024), though recent work has proven that ICL is not fully consistent with Bayesian inference (Falck et al.,, 2024). To our knowledge, this work is the first to leverage LLMs in a Bayesian framework for learning CBMs and show that rigorous statistical inference is feasible despite imperfections in LLM reasoning.
3 Method
Consider a dataset $\mathcal{D}=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$, where $x_{i}$ are model inputs and $y_{i}$ are labels. Let $\mathbf{X}\coloneq(x_{1},x_{2},\ldots,x_{n})$ and $\mathbf{y}\coloneq(y_{1},\ldots,y_{n})$. A concept $c$ defines a binary- or real-valued function $\phi_{c}$ of a model input, e.g. “Does the doctor note describe an elderly patient?” where $1$=yes and 00=no. A CBM selects a fixed number of $K$ concepts $\vec{c}=(c_{1},\ldots,c_{K})$ and fits a parametric model for $y$ using the extractions $(\phi_{c_{1}}(x),\ldots,\phi_{c_{K}}(x))$. For simplicity of presentation, our discussion will focus on logistic regression (LR)-based CBMs, although other choices are possible. In this case, the data is modeled by
$p\left(Y=1|X=x,\vec{\theta},\vec{c}\right)=\sigma\left(\sum_{k=1}^{K}\theta_{k}\phi_{c_{k}}(x)+\theta_{0}\right)$ (1)
where $\sigma$ is the sigmoid function and $\vec{\theta}=(\theta_{0},\cdots,\theta_{K})$ are the parameters. We refer to $\vec{c}$ as the “support” of the model. Throughout, we use capital and lowercase letters to differentiate between random variables and realized values, respectively. For instance, $C$ refers to a random concept whereas $c$ refers to a particular one.
3.1 Review: Bayesian variable selection
Learning CBMs shares many similarities with the more standard problem of learning sparse models in high-dimensional settings, where one must select among a finite number of features that have already been tabulated (Ishwaran et al.,, 2008; Castillo et al.,, 2015; Biswas et al.,, 2022). Here, we review the Bayesian solution for this more standard setting. In later sections, we describe extensions to the CBM setting, where there are a potentially infinite number of concepts and a cost associated with extracting each.
[…]
Inference for $p(\vec{c}|\mathbf{y},\mathbf{X})$ is typically achieved through Gibbs sampling. For each epoch, Gibbs rotates through indices $k=1,\ldots,K$, during which it replaces the $k$th concept in the current iterate $\vec{c}$ by drawing from the posterior distribution for $C_{k}$ conditional on the other concepts $\vec{c}_{-k}$. When one is unable to sample from the conditional posterior, Metropolis-within-Gibbs sampling can be used instead (Gilks et al.,, 1995). Rather than replacing the $k$th concept, it proposes a candidate concept $\check{c}_{k}$ given the other concepts per some distribution $Q(C_{k};\vec{C}_{-k}=\vec{c}_{-k})$ and accepts the candidate concept with probability $\alpha$. This acceptance probability $\alpha$, delineated in Algorithm 1, depends on the relative posterior probabilities of the candidate and existing concepts and is carefully designed so that the posterior distribution is stationary with respect to the sampling procedure.
Algorithm 1 Metropolis-within-Gibbs
1:Initialize concept set $\vec{c}=(c_{1},\cdots,c_{K})$
2:List of concept sets $L=[]$
3:for $t=1,2,\ldots,T$ do
4:     for $k=1,2,\ldots,K$ do
5:         $c_{k}\leftarrow$ MH-Update($\vec{c}$, $k$) //Update $k$-th concept
6:         Append the current $\vec{c}$ to $L$
7:return $L$
8:function MH-Update($\vec{c}$, $k$)
9:     Propose concept $\check{c}_{k}\sim Q(C_{k};\vec{c}_{-k})$
10:     $\alpha\leftarrow\min\left\{\frac{p((\vec{c}_{-k},\check{c}_{k})|\mathbf{y},\mathbf{X})Q(c_{k};\vec{c}_{-k})}{p(\vec{c}|\mathbf{y},\mathbf{X})Q(\check{c}_{k};\vec{c}_{-k})},1\right\}$
11:     if Accept with probability $\alpha$ then return $\check{c}_{k}$
12:     else return $c_{k}$
3.2 Bayesian variable selection with LLMs
Given the Gibbs procedures described above, there are various ways to integrate LLMs for learning CBMs. To begin, we describe two naïve approaches:
Approach 1: Propose with LLM “prior.” Let the LLM define a conditional prior over concepts, i.e. $p(C_{k}|\vec{c}_{-k})$, and propose concepts from this conditional prior, i.e. $Q(C_{k};\vec{c}_{-k})=p(C_{k}|\vec{c}_{-k})$. As concepts are generated in a data-agnostic fashion, the LLM may produce many irrelevant concepts that have low acceptance probabilities. So while Approach 1 may provide valid inference, it is inefficient and costly.
Approach 2: Propose with LLM “posterior.” Let the LLM conduct posterior inference for the $k$th concept given all the data and all but the $k$th concept through ICL and let the LLM propose from this conditional posterior, i.e. $Q(C_{k};\vec{c}_{-k})=p(C_{k}|\vec{c}_{-k},\mathbf{y},\mathbf{X})$. This simplifies Metropolis-within-Gibbs to Gibbs, which always accepts the proposed concept. As LLMs are not perfect Bayesian inference engines (Falck et al.,, 2024) and may be miscalibrated, this may lead to the inclusion of inappropriate concepts. While efficient, Approach 2 leads to biased sampling.
Instead, we propose to take the middle road, which we argue gives the best of both worlds. At each iteration, we first randomly split the data to construct subset $S$ of size $\lfloor\omega n\rfloor$ and its complement $S^{c}$, for some fraction $\omega$ (e.g. half). Then the LLM is asked to propose a candidate concept based on $\vec{c}_{-k}$ and subset $S$, which can be viewed as approximately sampling from the “partial posterior”
$Q(C_{k};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})\approx p(C_{k}|\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X}).$ (2)
Assuming (2) holds with equality, the Metropolis acceptance probability simplifies to
$\alpha\leftarrow\min\left\{\frac{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},(\vec{c}_{-k},\check{c}),\mathbf{X})}{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})},1\right\}.$
This essentially compares the likelihood of the held-out data with respect to the candidate and existing concepts, much like sample-splitting in frequentist settings. If the CBM with the candidate concept is better, it is accepted; otherwise it is accepted with probability $\alpha$. This procedure, which we refer to as Split-sample Metropolis-within-Gibbs (Algorithm 2) can learn from the data to override erroneous priors and concepts from the LLM, unlike Approach 2. In addition, it is efficient because the LLM will propose relevant concepts based on some of (though not all of) the data, unlike Approach 1.
Algorithm 2 Split-sample Metropolis update
1:function SS-MH-Update($\vec{c}$, $k$)
2:     Sample subset $S$ of size $\lfloor\omega n\rfloor$.
3:     Propose candidate $\check{c}\sim Q(C_{k};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})$
4:     $\alpha\leftarrow\min\left\{\frac{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},(\vec{c}_{-k},\check{c}),\mathbf{X})}{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})},1\right\}$
5:     if Accept with probability $\alpha$ then return $\check{c}$
6:     else return $c_{k}$
In practice, rather than sampling a single candidate concept, we use a version of multiple-try Metropolis-Hastings (Algorithm 3) (Liu et al.,, 2000). Here, the LLM proposes a batch of candidate candidates, and then we sample a candidate from this batch based on their posterior probabilities and consider the sampled candidate for acceptance. This is more cost-effective as it lets us batch concept annotations, as described later.
Algorithm 3 Multiple-try Split-sample Metropolis update
1:function Multi-SS-MH-Update($\vec{c}$, $k$)
2:     Sample subset $S$ of size $\lfloor\omega n\rfloor$.
3:     Propose $\check{c}^{(1)}_{k},...,\check{c}^{(M)}_{k}\sim Q(C_{k};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})$
4:     Sample $\check{m}\in\{1,\ldots,M\}$ with probabilities $\propto w_{m}\leftarrow p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},(\vec{c}_{-k},\check{c}^{(m)}),\mathbf{X})Q(\check{c}^{(m)};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})$
5:     $w_{0}\leftarrow p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})Q(c_{k};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})$
6:     $\alpha\leftarrow\min\left\{\frac{Q(c_{k};\vec{c}_{-k},\mathbf{y}_{S}\mathbf{X})\sum_{m=1}^{M}w_{m}}{Q(\check{c}_{k}^{(\check{m})};\vec{c}_{-k},\mathbf{y}_{S}\mathbf{X})\sum_{0\leq m\leq M,m\neq\check{m}}w_{m}},1\right\}$
7:     if Accept with probability $\alpha$ then return $\check{c}_{k}^{(\check{m})}$
8:     else return $c_{k}$
Assuming that the LLM’s partial posterior across all iterations is consistent with some prior distribution $p(\vec{C})$, we can show that the Markov chain defined by this procedure has the posterior $p(\vec{c}|\mathbf{y},\mathbf{X})$ as its stationary distribution (see Proposition I.1 in the Appendix). More importantly, we can prove that the procedure still converges to the true concepts even in the likely scenario where there is no valid prior.
Theorem 3.1.
Suppose the data is IID. Let $L(\vec{c})\coloneqq\max_{\vec{\theta}}\mathbb{E}\{\log p(Y|X,\vec{\theta},\vec{c})\}$ and $\mathcal{C}^{*}\coloneqq\operatorname{argmax}_{\vec{c}}L(\vec{c}).$ For sample size $n$, let $\Pi_{n}$ denote the set of stationary distributions of the Markov chain defined by running Algorithm 1 with SS-MH-Update or Multi-SS-MH-Update instead of MH-Update. Assuming a Gaussian prior for $\vec{\theta}$ and under regularity conditions (see Assumption H.1 in the Supplement), then $\inf\{\pi(\mathcal{C}^{*})\colon\pi\in\Pi_{n}\}\to 1$ in probability as $n\to\infty$.
3.3 BC-LLM
We now describe the main steps in BC-LLM (Fig 1): the procedure is initialized in Step 0, and each iteration (Steps 1 to 4) samples a CBM. The samples then describe the posterior distribution over CBMs and form an ensemble prediction model (after removing those from the burn-in period). Depending on one’s use case, the search over concepts can be guided by the LLM’s prior knowledge about the target $Y$ by injecting this information into the prompts (e.g. “we are trying to predict readmission…”) or in an “unbiased” fashion by masking the meaning of $Y$ (e.g. “we are trying to predict some label $Y$…”). Example prompts and other implementation details are provided in the Appendix.
Step 0: Initialization and keyphrase extraction
We first use an LLM to extract “keyphrases” describing each observation $x$, which will later be used by the LLM to brainstorm candidate concepts. Here, we use “keyphrases” to refer to short phrases that describe the observation: keyphrases for images describe what’s in the image (e.g. blue eyes), while keyphrases for text data can be phrases that appear in the text or summarize the main ideas (e.g. diabetes) (Papagiannopoulou and Tsoumakas,, 2020). Unlike concepts, keyphrases do not define a function that outputs a value, though they can help one come up with a concept (e.g. “diabetes” is a keyphrase whereas the binary questions “Does the patient currently have diabetes?” or “Does the patient have a family history of diabetes?” are concepts). In addition, this step does not restrict the set of keyphrases one can extract to a prespecified list, which allows for a wide exploration of keyphrases. The tradeoff is that the extracted keyphrases may be, and likely are, an incomplete summary of each observation.
To initialize the CBM sampling procedure, these keyphrases are converted into numeric vectors to fit a (penalized) LR model for the target $Y$, similar to the process for fitting a bag-of-words model. An LLM analyzes the top keyphrases in this LR model to generate an initial set $\vec{c}$ of $K$ concepts. The LLM also extracts these concepts from each observation.
Step 1: Drop a concept & fit a keyphrase model
Each iteration proposes candidate concepts based on some data subset $S$ to replace the $k$th concept for $k\in\{1,\cdots,K\}$. Given the context limits of LLMs, it is unlikely that an LLM can generate a candidate concept from the partial posterior distribution by reviewing each observation in $S$ individually. As such, we will generate a summary by highlighting the top keyphrases in $S$ for predicting $Y$, on top of the existing $K-1$ concepts $\vec{c}_{-k}$.
For continuous outcomes, this can be accomplished by finding the top keyphrases associated with the residuals of a model that predicts $Y$ given $\vec{c}_{-k}$. As the residuals are noisy for binary/categorical outcomes, we fit a (multinomial) LR model of $Y$ instead, where the input features are the keyphrases (again encoded using a bag-of-words approach) and annotated concepts $\vec{c}_{-k}$. The parameters of this “keyphrase model” $f$ are the coefficients for the keyphrases and concepts, denoted $\beta_{W}$ and $\beta_{\vec{c}_{-k}}$, respectively. Given the high-dimensionality of the keyphrases, we penalize $\beta_{W}$ using some penalty function $J$ and estimate $f$ by solving
$\min_{\beta_{\vec{c}_{-k}},\beta_{W}}\frac{1}{n}\sum_{i=1}^{n}\ell\left(y_{i},f\left(x_{i}|\beta_{\vec{c}_{-k}},\beta_{W}\right)\right)+\lambda J(\beta_{W}),$
where $\ell$ is the logistic loss and penalty parameter $\lambda\geq 0$ is selected by cross-validation. We found the ridge penalty (i.e., $J(\beta_{W})=\|\beta_{W}\|_{2}^{2}$) to work well in practice, though other choices for $J$ may be used instead.
Step 2: Query the LLM for candidate concepts Given the summary of the top keyphrases generated from Step 1, the LLM is asked to propose $M$ candidates for the $k$-th concept from the partial posterior $p(C_{k}|\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})$. In addition, we ask the LLM to provide the value of the partial posterior probability for each candidate concept. For instance, if the top keyphrases include “chronic kidney disease” and “acute kidney injury,” the LLM may propose a concept like “Does the patient have a history of kidney disease?” with an assigned partial posterior probability of 0.4.
Step 3: Concept annotation Given the batch of candidate concepts, we use the LLM to extract their values for each observation $x$, i.e. $\phi_{c}(x)$ for $c\in\{\check{c}^{(1)}_{k},\ldots,\check{c}^{(M)}_{k}\}$. Because multiple-try Metropolis-within-Gibbs proposes a batch of candidate concepts, we can batch their extractions into a single LLM query for each observation. This makes BC-LLM more cost/query-efficient than ordinary Metropolis-within-Gibbs, which outputs only a single concept per iteration and thus cannot batch extractions.
Step 4: Accept/reject step The final step is to compute the sampling probabilities for the candidate concepts and the acceptance ratio, which involve comparing the proposal probabilities for the existing and candidate concept sets $\vec{c}$ as well as their split-sample posteriors
[…]
To estimate the integral (3), one approach is to sample from the posterior on $\vec{\theta}$, but this can be computationally expensive since one would need to implement posterior sampling in inner and outer loops. Alternatively, in the case of LR-based CBMs, one can use a Laplace-like approximation for (3) (Lewis and Raftery,, 1997), which is discussed in more detail in the Appendix. This essentially approximates the partial and full posterior distributions with Gaussian distributions. As the posterior distribution of LR parameters are asymptotically normal, this approximation becomes increasingly accurate as $n\rightarrow\infty$.
3.4 Computational cost
The key computational cost of BC-LLM is from querying the LLM. Running BC-LLM requires $O(nTK)$ queries, where $T$ is the number of epochs. $K$ is typically small by design, because large $K$ leads to less interpretable CBMs. One may be concerned by $T$, as standard Bayesian inference procedures typically draw thousands of posterior samples. Nevertheless, combined with a greedy warm-start procedure, we found that a small $T$ to be quite effective; our experiments go as low as $T=5$. Generally speaking, if the goal is solely prediction accuracy, a small $T$ may suffice. If the goal is uncertainty quantification for relevant concepts, a larger $T$ (e.g., 10) is helpful for more complete exploration of concepts.
Figure 2: MIMIC evaluation: (a) Comparison of BC-LLM and existing methods in terms of performance and recovery of true concepts. Error bars indicate the 95% CI. (b) Dendrograms of concepts learned by BC-LLM with 100 and 800 training observations (left and right, respectively). Labels are shortened concept questions, which generally have the format “Does the note mention the patient having…?” Highlighted labels correspond to the true, color-coded concepts in Section 4.2. Numbers in parentheses describe the proportion of posterior samples with the concept.
4 Experimental results
4.1 Experimental setup
We compare BC-LLM against One-pass summarization, where an LLM performs post-hoc distillation of the keyphrase model into a CBM (Pham et al.,, 2023), and Boosting, where an LLM iteratively adds concepts to a CBM by analyzing observations with large residuals (Ludan et al.,, 2023; Liu et al., 2024a, ). We also compare it to black-box and semi-interpretable models to assess whether BC-LLM can close the performance gap. When the true concepts are known, we also fit an Oracle CBM.
Methods were evaluated with respect to predictive performance as measured by AUC and uncertainty quantification as measured by log likelihood (higher is better). When true concepts are known, we also quantify concept selection rates using “concept precision,” as defined by $\frac{1}{K}\sum_{k=1}^{K}p(\hat{C}_{k}\notin\vec{c}^{*}\mid\mathbf{y},\mathbf{X})$, and “concept recall,” as defined by $\frac{1}{K}\sum_{k=1}^{K}p(c_{k}^{*}\notin\hat{\vec{C}}\mid\mathbf{y},\mathbf{X})$, where $\hat{\vec{C}}$ is the posterior distribution over concept sets from BC-LLM or a single set of concepts learned by a non-Bayesian procedure. Quantifying a “correct” concept match requires going beyond exact string matching, as there are many possible refinements for a single concept. For instance, for the concept “Does the note mention the patient smoking in the present or the past?”, there are rephrasings (“Does this note mention the patient currently smoking or having smoked?”), polar opposites (“Does the note not mention the currently smoking or having previously smoked?”), or closely overlapping concepts (“Does the note mention the patient currently smokes?”). To quantify if a learned concept was correct, we used the following rule: if the absolute value of the Pearson correlation for LLM annotations of two concepts exceeds 50%, the two concepts are sufficiently similar. We manually confirmed that any pair of learned and true concepts that passed this rule were always closely related.
Experiments used GPT-4o-mini (OpenAI,, 2023) and a protected health information (PHI)-compliant version of GPT-4o for real-world clinical notes in Section 4.4. Additional implementation details and experimental settings are in the Appendix.
4.2 Simulation study: Learning CBMs for clinical notes from MIMIC
We first compared BC-LLM against existing methods in terms of predictive performance and concept selection rates. To assess how well the methods perform even when the LLM has no prior knowledge about the target, we simulated $Y$ using a LR model with five Social Determinants of Health as inputs, which were annotated from real-world patient notes in MIMIC-IV (Johnson et al.,, 2023; Ahsan et al.,, 2021):
1. Does the note mention the patient consuming alcohol in the present or the past?
2. Does the note mention the patient smoking in the present or the past?
3. Does the note mention the patient using recreational drugs in the present or the past?
4. Does the note imply the patient is unemployed?
5. Does the note imply the patient is retired?
CBMs with $K=6$ were trained on 100, 200, 400, and 800 observations and evaluated on 500 held-out observations. This was done for 4 replicates, where the data were randomly selected from the full dataset of 7000 observations. An Oracle CBM was also fit as the concepts are known.
BC-LLM outperforms the comparator methods both in predictive performance and uncertainty quantification (Figure 2a) and, with only 400 observations, performs as well as Oracle. The performance gap between BC-LLM and the comparator methods widens with more training observations because BC-LLM can iteratively refine its concepts. Although Boosting is also an iterative procedure, it struggles to hypothesize relevant concepts because it reviews only a few misclassified observations each iteration and cannot revise concepts added in previous iterations. Additionally, BC-LLM was better at recovering the true concepts compared to existing methods.
Table 1: AUCs of CBMs and black-box models (ResNet) for classifying bird subcategories.
Bird # classes BC-LLM One-pass Boosting ResNet
Albatross 3 0.837 (0.771, 0.894) 0.833 (0.771, 0.890) 0.775 (0.709, 0.842) 0.837 (0.767, 0.897)
Auklet 4 0.884 (0.830, 0.928) 0.865 (0.813, 0.914) 0.733 (0.671, 0.792) 0.831 (0.767, 0.893)
Blackbird 4 0.869 (0.822, 0.912) 0.719 (0.671, 0.771) 0.749 (0.689, 0.801) 0.857 (0.811, 0.900)
Bunting 3 0.932 (0.880, 0.972) 0.629 (0.585, 0.672) 0.681 (0.593, 0.762) 0.919 (0.869, 0.962)
Flycatcher 7 0.862 (0.835, 0.887) 0.800 (0.766, 0.833) 0.784 (0.753, 0.813) 0.817 (0.781, 0.849)
Sparrow 21 0.803 (0.785, 0.819) 0.723 (0.704, 0.739) 0.618 (0.597, 0.640) 0.812 (0.794, 0.830)
Tern 7 0.817 (0.784, 0.849) 0.705 (0.667, 0.740) 0.609 (0.567, 0.651) 0.714 (0.674, 0.751)
The posterior distribution over concepts can be visualized using hierarchical clustering (Figure 2b). Comparing concepts learned given 100 versus 800 observations, we find that the posterior learned by BC-LLM notably contracts towards the true concepts with increasing data.
4.3 Learning to classify birds
To demonstrate the multimodality of BC-LLM, we evaluate how well BC-LLM and other CBM methods can differentiate between bird categories in the CUB-birds image dataset (Welinder et al.,, 2010). [^1] We classify bird categories rather than all subtypes because there is likely a parsimonious CBM for classifying between categories but not between all subtypes. While the exact set of $K$ concepts needed to classify bird categories are unknown, we know certain features that are certainly spurious (e.g. features in the background are irrelevant).
Similar to the prior experiment, we do not tell the LLM that we are predicting bird types. By replacing the bird labels with one-hot encodings, the supervised learning task is more difficult, as the LLM cannot rely on prior knowledge about birds and must conduct an unbiased exploration over concepts. The data is split 50/50 between training and testing (so 30 bird images per category are available for training). Due to cost limitations, seven bird classification tasks were considered. The number of concepts $K$ learned for each task was set to the number of classes, but no smaller than 4 and no greater than 10. For the black-box comparator, we fine-tuned the last layer of ResNet50 pre-trained on ImageNetV2 (He et al.,, 2016; Recht et al.,, 2019).
BC-LLM outperforms all the other CBM learning procedures for every bird type (Table 1). Moreover, it outperforms ResNet50 model in all but one bird. Whereas ResNet50 easily overtrained on the small bird datasets, BC-LLM identified relevant concepts as inputs to a simple parametric model, leading to faster model convergence.
As an example, we show concepts learned by BC-LLM for classifying Bunting birds in Figure 3b. The concepts learned by BC-LLM tended to be more relevant than those from comparator methods. For instance, given all the training data, One-pass summarization checks for a “bird sitting sideways”, “bird on a perch”, and “bird feeding or foraging,” whereas BC-LLM assigned bird position a low posterior weight. Moreover, visualizing the learned concepts helps us understand the learning procedure itself. For instance, we see that BC-LLM learns to downweight spurious features, i.e. concepts relating to backgrounds or bird positions, with accumulating data.
We also assess how robust BC-LLM is in out-of-distribution (OOD) settings, as the Bayesian framing should improve uncertainty quantification. We apply the Bunting CBMs to classify images of the first author’s dog (Fig 3c). Compared to other CBM methods, BC-LLM was more appropriately unsure which Bunting bird to classify the dog as. Interestingly, when the dog is wearing a rainbow unicorn costume (C2), One-pass and Boosting overconfidently classify the dog as the colorful Bunting category (A3) with probabilities greater than 95%. Likewise, when these CBMs were used to classify images of finches (also small colorful birds), the average entropy of BC-LLM was again highest.
Method C1 C2 Finches (e.g., C3)
BC-LLM 0.91 0.90 0.70 (0.53 -0.88)
One-pass 0.06 0.20 0.27 (0.00-0.66)
Boosting 0.00 0.00 0.15 (0.00-0.57)
Figure 3: Bunting bird classification across categories A1-A3. (b) Highlighted labels in the dendrogram correspond to relevant features. Labels are shortened concept questions, generally in the format “Does the image depict…?” Numbers in parentheses describe the proportion of posterior samples with the concept. Left and right dendrograms correspond to learned posteriors for 1/3 and 3/3 of the training data, respectively. (c) To test robustness in OOD settings, the Bunting bird CBMs are asked to classify two images of the first author’s dog and 30 images of finches. Entropies of the model predictions (a measure of model uncertainty) are shown for all the CBM methods. 95% CI is shown for the Finch dataset.
Method AUC (95% CI) Log Lik (95% CI)
BC-LLM 0.64 (0.58, 0.70) -0.45 (-0.50, -0.40)
One-pass 0.59 (0.52, 0.65) -0.46 (-0.51, -0.41)
Boosting 0.59 (0.52, 0.66) -0.46 (-0.51, -0.41)
Bag-of-words 0.52 (0.46, 0.58) -0.47 (-0.52, -0.41)
Figure 4: Learning to augment a readmission risk prediction model for heart failure patients. Labels are shortened versions of concept questions of the format “Does the note mention the patient having…?”
4.4 Augmenting an existing tabular model with real-world clinical notes
Finally, we demonstrate how BC-LLM can help ML developers integrate different data modalities to improve model performance. In particular, the data science team at the Zuckerberg San Francisco General Hospital (ZSFG) has both tabular data and unstructured clinical notes from the electronic health record (EHR). The team has so far fit a model using the tabular data to predict 30-day unplanned readmission risk for patients with congestive heart failure (CHF). Using CBMs, the team wants to assess if clinical notes contain valuable information to extract as additional features. Moreover, the learned model should be interpretable and easy to audit.
To address this, we extend BC-LLM by having the CBM take as inputs (i) the risk prediction from the original tabular model and (ii) $K$ concepts extracted from the clinical notes. This can be viewed as revising the existing model by including additional features. We set to $K=4$. The CBMs were trained on 1000 patients and evaluated on 500 held-out patients. Data in this experiment contained PHI, for which IRB approval was obtained.
The original tabular model, trained on lab and flowsheet values, achieved an AUC of 0.60. BC-LLM revised this model to perform substantially better, achieving an AUC of 0.64 (Figure 4). Moreover, it outperforms comparator methods with respect to both AUC and log likelihood. When the CBMs were shown to four clinicians, [^2] All clinicians work or have worked in ZSFG. Three are in the hospital’s data science team, one of whom (LZ) is a coauthor on this paper. The fourth leads a separate clinical data science research group. They were all interviewed separately. they found that the CBM learned by BC-LLM to be substantially more interpretable and meaningful than the existing tabular model. The clinicians from the hospital’s data science team noted that they had not considered engineering many of the learned concepts on the list. All clinicians noted that many of the learned concepts are known to be causally relevant and would have been difficult to learn from existing tabular data. For example, “substance dependence” is rarely documented directly in the tabular data, and information relevant to determining its value is often missing. Finally, the clinicians highly valued the fact that many of the learned concepts were clinically actionable, such as helping patients better control their diabetes or opting for medication regimes that are easier for patients to follow.
5 Discussion
BC-LLM is a new method for learning CBMs that iteratively proposes concepts using an LLM within a Bayesian framework, which allows for rigorous uncertainty quantification despite LLMs being prone to error and hallucinations. The method is highly general: it is compatible with various data modalities (text, images, and tabular data) and can be extended beyond the settings of binary and multiclass classification. The empirical results show that BC-LLM outperforms existing methods, learns to recover relevant concepts and downweight spurious ones, and can suggest useful features for ML developers to engineer. Code for running BC-LLM and reproducing results in the paper are available at https://github.com/jjfeng/bc-llm.
Acknowledgements
JF was funded through a Patient-Centered Outcomes Research Institute® (PCORI®) Award (ME-2022C1-25619). YT gratefully acknowledges support from NUS Start-up Grant A-8000448-00-00 and MOE AcRF Tier 1 Grant A-8002498-00-00. The authors thank the UCSF AI Tiger Team, Academic Research Services, Research Information Technology, and the Chancellor’s Task Force for Generative AI for their technical support related to the use of Versa API gateway. The authors thank Yifan Mai, Seth Goldman, Andrew Bishara, and Jim Marks for their invaluable feedback on the work. The views presented in this work are solely the responsibility of the author(s) and do not necessarily represent the views of the PCORI®, its Board of Governors or Methodology Committee.
Appendix
Appendix A Multiple-Try Metropolis-Hastings
In this section, we provide intuition for the Multiple-Try Metropolis-Hastings partial posterior method (Algorithm 3). To draw connections with the original Multiple-Try Metropolis-Hastings method (Liu et al.,, 2000), we first work in a completely abstract setting.
Given a state space $\Omega$, stationary distribution $\pi$, and proposal transition kernel $T(y;x)$, define weights $w(x,y)=\pi(x)T(y;x)$. Note that under regular Metropolis-Hastings, the acceptance ratio for proposing the state $y$ given a state $x$ is $\min\{w(y,x)/w(x,y),1\}$.
Classical version.
We briefly describe Liu et al., (2000)’s original method. Suppose the current state is $x$.
  1. 1.
    Draw $z_{1},z_{2},\ldots,z_{M}\sim T(Z;x)$.
  2. 2.
    Sample $y$ from a distribution with probabilities
    $\mathbb{P}\{Y=z_{j}\}=\frac{w(z_{j},x)}{\sum_{i=1}^{M}w(z_{i},x)}.$ (4)
    Let $\check{m}$ denote the sampled index.
  3. 3.
    Sample $z_{1}^{*},\ldots,z_{M-1}^{*}\sim T(Z;y)$, and set $z_{M}^{*}=x$.
  4. 4.
    Accept $y$ with probability
    $\beta(x,y;\mathbf{z},\mathbf{z}^{*})=\min\left\{\frac{\sum_{i=1}^{M}w(z_{i},x)}{\sum_{i=1}^{M}w(z_{i}^{*},y)},1\right\}.$ (5)
One can show that this Markov chain satisfies the detailed balance equations when marginalizing out the intermediate states $z_{1},z_{2},\ldots,z_{M},z_{1}^{*},\ldots,z_{M}$ (see Theorem 1 in Liu et al., (2000)). We would like, however, to minimize the number of proposals necessary, so it is natural to ask whether the algorithm can be modified such that the proposals needed for the backward transitions, $z_{1}^{*},\ldots,z_{M-1}^{*}$, can be omitted. This is indeed possible.
Modified version.
  1. 1.
    Draw $z_{1},z_{2},\ldots,z_{M}\sim T(Z;x)$.
  2. 2.
    Sample $y$ from a distribution with probabilities
    $\mathbb{P}\{Y=z_{j}\}=\frac{w(z_{j},x)}{\sum_{i=1}^{M}w(z_{i},x)}.$ (6)
    Let $\check{m}$ denote the sampled index, $\mathbf{z}=(z_{1},z_{2},\ldots,z_{M})$, $\mathbf{z}^{*}=(z_{1},\ldots,z_{\check{m}-1},x,z_{\check{m}+1},\ldots,z_{M})$, and set $q(y;\mathbf{z})=\mathbb{P}\{Y=y|z_{1},\ldots,z_{M}\}$.
  3. 3.
    Accept $y$ with probability
    $\begin{split}\alpha(x,y;\mathbf{z}_{-\check{m}})&=\min\left\{\frac{p(y\rightarrow\mathbf{z}^{*}\rightarrow x)}{p(x\rightarrow\mathbf{z}\rightarrow y)},1\right\}\\ &=\min\left\{\frac{\pi(y)\prod_{i=1}^{M}T(z_{i};y)q(x;\mathbf{z}^{*})}{\pi(x)\prod_{i=1}^{M}T(z_{i}^{*};x)q(y;\mathbf{z})},1\right\}.\end{split}$ (7)
Proposition A.1.
The modified Multiple-Try MH sampler is reversible.
Proof.
Let $\Phi(y;x)$ denote the actual transition matrix. We want to show
$\pi(x)\Phi(y;x)=\pi(y)\Phi(x;y).$ (8)
To see this, we compute
[…]
Connections between modified and classical versions.
The acceptance probability in the modified version of Multiple-Try MH can be expanded as follows:
$\begin{split}\min\left\{\frac{\pi(y)\prod_{i=1}^{M}T(z_{i};y)q(x;\mathbf{z}^{*})}{\pi(x)\prod_{i=1}^{M}T(z_{i}^{*};x)q(y;\mathbf{z})},1\right\}&=\min\left\{\frac{\pi(y)\prod_{i=1}^{M}T(z_{i};y)\frac{w(x,y)}{\sum_{i=1}^{M}w(z_{i}^{*},y)}}{\pi(x)\prod_{i=1}^{M}T(z_{i}^{*};x)\frac{w(y,x)}{\sum_{i=1}^{M}w(z_{i},x)}},1\right\}\\ &=\min\left\{\frac{\sum_{i\neq\check{m}}w(z_{i},x)+w(y,x)}{\sum_{i\neq\check{m}}w(z_{i},y)+w(x,y)}\cdot\prod_{i=2}^{M}\frac{T(z_{i};y)}{T(z_{i};x)},1\right\}.\end{split}$ (10)
If $T(-,-)$ is invariant in the first argument (as in our application to BC-LLM), then this last formula is exactly equal to $\beta(x,y;\mathbf{z},\mathbf{z}^{*})$, the acceptance probability for the classical version, but if we were to use the same points for both the non-realized forward and backward proposals.
Equivalence with Multi-SS-MH-Update.
Finally, we show that the modified version of Multiple-Try MH described in the previous sections is equivalent to what is implemented in Multi-SS-MH-Update, for a fixed $S$. To see this, we make the replacements:
  • $x\leftarrow\vec{c}$
  • $z_{i}\leftarrow(\check{c}_{k}^{(i)},\vec{c}_{-k})$ for $i=1,\ldots,M$
  • $y\leftarrow(\check{c}_{k}^{(\check{m})},\vec{c}_{-k})$
  • $T(-;-)\leftarrow Q(-;-,\mathbf{y}_{S},\mathbf{X})$
  • $\pi(-)\leftarrow p(-|\mathbf{y},\mathbf{X})$
Plugging these into (10), the acceptance ratio becomes
$\begin{split}&\min\left\{\frac{Q(c_{k};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})\sum_{m=1}^{M}p((\check{c}^{(m)},\vec{c}_{-k})|\mathbf{y},\mathbf{X})}{Q(\check{c}_{k}^{(\check{m})};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})\sum_{m=1}^{M}p((\check{c}^{(m)},\vec{c}_{-k})|\mathbf{y},\mathbf{X})},1\right\}\\ =~{}&\min\left\{\frac{Q(c_{k};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})\sum_{m=1}^{M}p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},(\check{c}_{k}^{(m)},\vec{c}_{-k}),\mathbf{X})Q(\check{c}_{k}^{(m)};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})}{Q(\check{c}_{k}^{(\check{m})};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})\sum_{0\leq m\leq M,m\neq\check{m}}p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},(\check{c}_{k}^{(m)},\vec{c}_{-k}),\mathbf{X})Q(\check{c}_{k}^{(m)};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})},1\right\},\end{split}$ (11)
where the equality comes from (48) and we set $\check{c}_{k}^{(0)}=c_{k}$ for convenience of notation. Observe that this is exactly the formula in Line 6 of Algorithm 3.
Meanwhile, plugging into (6) and using (48), we see that the sampling weights used in the modified version of Multiple-Try MH are equivalent to those in Line 3 of Algorithm 3. This completes the proof of the equivalence.
Appendix B Implementation details for BC-LLM
Here we discuss how the hyperparameters for BC-LLM should be selected:
  • Fraction $\omega$ of data used for partial posterior: Choosing a small $\omega$ may lead to the LLM proposing less relevant concepts, but tends to lead to more diverse proposals. In contrast, a large $\omega$ tends to lead to less diverse proposals because the LLM is encouraged to propose concepts that are relevant to the dataset $\mathcal{D}$, which may not necessarily generalize. In experiments, we found that $\omega=0.5$ provided good results.
  • Number of candidate concepts $M$: More candidate proposals per iteration can allow for more efficient exploration of concepts. The number of candidates is limited by the number of concepts that the LLM can reliably extract in a single batch. There tends to also be diminishing returns, as the first few candidates generated by the LLM based on the top keyphrases tend to be highest quality and most relevant. We found that setting $M=10$ provided good performance.
  • Warm-start and Burn-in: Since Gibbs sampling can be slow to converge, we precede it with a warm-start, which we obtain by updating concepts greedily. That is, we select the concept that maximizes $\operatorname{argmax}p(\gamma|\vec{c}_{-k},\mathbf{y}_{S},\boldsymbol{\Phi})$, instead of sampling from the distribution. In experiments, we run warm-start for one epoch and stored the last 20 iterates as posterior samples; the rest of the samples were treated as burn-in.
  • Number of iterations $T$: Although Monte Carlo procedures for Bayesian inference typically have thousands of samples, this is cost-prohibitive when an LLM must query each observation for a new concept per iteration. In experiments, we found that even setting $T$ as low as $4$ to be quite effective. Generally speaking, if the goal is solely prediction accuracy, a small $T$ may suffice. On the other hand, if the goal is uncertainty quantification for the relevant concepts, one may prefer $T$ on the higher end (e.g. 10) for more complete exploration of concepts.
Appendix C Example prompt for keyphrase extraction (Step 0)
Here is a clinical note: $\langle$ note $\rangle$
Output a list of descriptors that summarizes the patient case (such as aspects on demographics, diagnoses, social determinants of health, etc). For each descriptor, also list as many descriptors that mean the same thing or generalizations of the descriptor. All descriptors, synonyms, and generalizations cannot be more than two words. Output as a JSON in the following format. Do not output more than 30 entries in the JSON:
Appendix D Example prompt for proposing concepts (Step 2)
The goal is to come up with a concept bottleneck model (CBM) that only extracts 5 meta-concepts from patient notes to predict some outcome Y with maximum accuracy. A meta-concept is a binary feature extractor defined by a yes/no question. We have 4 meta-concepts so far:
  1. 1.
    Does the note mention the patient having social support?
  2. 2.
    Does the note mention the patient having stable blood pressure?
  3. 3.
    Does the note mention the patient not experiencing respiratory distress?
  4. 4.
    Does the note mention the patient experiencing substance abuse or dependence?
To come up with the 5th meta-concept, I have done the following: I first fit a CBM on the 4 existing meta-concepts. Then to figure out how to improve this 4-concept CBM, I first asked an LLM to extract a list of concepts that are present in each note, and then fit a linear regression model on the extracted concepts to predict the residuals of the 4-concept CBM. These are the top extracted concepts in the resulting residual model, in descending order of importance:
$\langle$ top words from residual model $\rangle$
To interpret this residual model, a general rule of thumb is that an extracted concept with a large positive coefficient means that the outcome is positively associated with the concept being mentioned in the note, and an extracted concept with a negative coefficient means that the outcome is negatively associated with the concept being mentioned in the note. Remember that notes typically discuss the current state as well as the past medical history of a patient, so a concept being mentioned is in reference to the present OR past state of the patient.
Given the residual model, create cohesive candidates for the 5th meta-concept. Be systematic and consider all the listed concepts in the residual model. Start from the most to the least predictive concept. For each concept, check if it matches an existing meta-concept or create a new candidate meta-concept. Work down the list, iterating through each concept. Clearly state each candidate meta-concept as a yes/no question.
Suggestions for generating candidate meta-concepts: Do not propose meta-concepts that are simply a union of two different concepts (e.g. “Does the note mention this patient experiencing stomach pain or being female?” is not allowed), questions with answers that are almost always a yes (e.g. the answer to “Does the note mention this patient being sick?” is almost always yes), or questions where the yes/no options are not clearly defined (e.g. “Does the note mention this patient experiencing difficulty?” is not clearly defined because difficulty may mean financial difficulty, physical difficulties, etc). Do not propose meta-concepts where you would expect over 95% agreement or disagreement with the 4 existing meta-concepts (e.g. “Does the note mention the patient having high blood pressure?” overlaps too much with “Does the note mention the patient having hypertension?”).
Finally, summarize all the generated candidates for the 5th meta-concept in a JSON. Merge any candidate meta-concepts that are essentially the same (where you would expect over 95% agreement) or essentially opposites (you would expect over 95% disagreement). In the JSON, include a list of comma-separated list of phrases that mean the same or opposite of each candidate meta-concept.
Appendix E Example prompt for annotations for candidate concepts (Step 3)
A simple concept extraction procedure is to obtain binary or categorical concept extractions. In line with findings in Kim et al., (2023), we found that BC-LLM performed better when the LLM quantified its uncertainty when extracting concepts. So if the LLM was unsure about the concept’s value, it was allowed to return a probability. As such, we use prompts like the following:
You will be given a clinical note. I will give you a series of questions. Your task is answer each question with 1 for yes or 0 for no. If the answer to the question is not clearly yes or no, you may answer with the probability that the answer is a yes. Respond with a JSON that includes your answer to all of the questions. Questions:
  1. 1.
    Does the note mention the patient having social support?
  2. 2.
    Does the note mention the patient having stable blood pressure?
  3. 3.
    Does the note mention the patient not experiencing respiratory distress?
  4. 4.
    Does the note mention the patient experiencing substance abuse or dependence?
  5. 5.
    Does the note mention the patient being uninsured?
clinical note: $\langle$ note $\rangle$
Appendix F Experiment details
Note on computation time: When running BC-LLM with GPT-4o-mini-2024-07-18 and GPT-4o, each iteration completed within a few minutes.
F.1 MIMIC
The entire dataset consists of 7043 observations of which we trained on 100, 200, 400, and 800 randomly selected observations and evaluated 500 held-out observations. We used the “Chief Complaint” and “Social History” sections of the notes. The label $Y$ was generated per a LR model, in which
[…]
One-pass summarization and BC-LLM were run to fit CBMs with $K=6$ concepts. BC-LLM was run for $T=5$ epochs. Boosting ran for 10 iterations, in which at most 1 new candidate was added each iteration.
F.2 CUB-Birds dataset
To train a bird classifier for $R$ subtypes, One-pass summarization and BC-LLM were run to fit CBMs with $K=\min(10,\max(4,R))$ concepts. BC-LLM was run for a maximum of $T=5$ epochs or 25 iterations, whichever was reached first. Boosting ran for 10 iterations, in which at most 1 new candidate was added each iteration.
F.3 Clinical notes and tabular data from ZSFG
To predict readmission risk, the models were trained to analyze the sections “Brief history leading to hospitalization” and the “Summary of hospitalization” in the discharge summary for each patient. One-pass summarization and BC-LLM were run to fit CBMs with $K=4$ concepts. BC-LLM was run for $T=5$ epochs. Boosting ran for 10 iterations, in which at most 1 new candidate was added each iteration. For the black-box comparator, we used a ResNet50 model pre-trained on ImageNetV2, which ensures ResNet’s training set does not overlap with the CUB-Birds dataset.
For the OOD experiment, the Bunting CBM was applied to 30 images of finches.
Appendix G Laplace Approximation of Split-Sample Posterior
In this section, we describe how to perform Laplace approximation of the split-sample posterior $p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})$ when the likelihood model is logistic and the prior on the coefficient vector $\vec{\theta}$ is a standard normal $\mathcal{N}(0,\gamma^{2}I)$. For simplicity, we omit discussing the constant term $\theta_{0}$. Fixing the list of concepts $\vec{c}$, we let $\Phi$ denote the $n\times K$ matrix whose $(i,j)$-th entry is given by $\Phi_{ij}=\phi_{c_{j}}(x_{i})$. Given any subset of example indices $T\subset[n]$, we can write the joint likelihood for these examples as
$p(\mathbf{y}_{T}|\vec{\theta},\vec{c},\mathbf{X})=\exp\left(\sum_{i\in T}\left(y_{i}\vec{\theta}^{T}\Phi_{i\cdot}-\log(1+\exp(\vec{\theta}^{T}\Phi_{i\cdot}))\right)\right).$ (12)
Multiplying by the prior for $\vec{\theta}$, the joint conditional distribution for $\mathbf{y}_{T}$ and $\vec{\theta}$ is then:
$p(\vec{\theta},\mathbf{y}_{T}|\vec{c},\mathbf{X})=(2\pi\gamma^{2})^{-1/2}\exp\left(\sum_{i\in T}\left(y_{i}\vec{\theta}^{T}\Phi_{i\cdot}-\log(1+\exp(\vec{\theta}^{T}\Phi_{i\cdot}))\right)-\frac{\|\vec{\theta}\|_{2}^{2}}{2\gamma^{2}}\right).$ (13)
To form the Laplace approximation to (13), we first denote
$\begin{split}g_{T}(\vec{\theta})&\coloneqq-\log p(\vec{\theta},\mathbf{y}_{S^{c}}|\vec{c},\mathbf{X})\\ &=\sum_{i\in T}\left(-y_{i}\vec{\theta}^{T}\Phi_{i\cdot}+\log(1+\exp(\vec{\theta}^{T}\Phi_{i\cdot}))\right)+\frac{\|\vec{\theta}\|_{2}^{2}}{2\gamma^{2}},\end{split}$ (14)
and its Hessian by
$H_{T}(\vec{\theta})\coloneqq\nabla^{2}g_{T}(\vec{\theta})=\sum_{i\in T}\frac{e^{\vec{\theta}^{T}\Phi_{i\cdot}}}{(1+e^{\vec{\theta}^{T}\Phi_{i\cdot}})^{2}}\Phi_{i\cdot}\Phi_{i\cdot}^{T}+\gamma^{-2}I.$ (15)
Let
$\vec{\theta}_{\operatorname{\operatorname{MAP}},T}\coloneqq\operatorname{argmin}g_{T}(\vec{\theta})$ (16)
be the maximum a posteriori (MAP) estimate. We can then perform a quadratic approximation of the exponent around $\vec{\theta}_{\operatorname{MAP},T}$ to get
$p(\vec{\theta},\mathbf{y}_{T}|\vec{c},\mathbf{X})\approx p(\vec{\theta}_{\operatorname{MAP},T},\mathbf{y}_{T}|\vec{c},\mathbf{X})\exp\left(-\frac{1}{2}(\vec{\theta}-\vec{\theta}_{\operatorname{MAP},T})^{T}H_{T}(\vec{\theta}_{\operatorname{MAP},T})(\vec{\theta}-\vec{\theta}_{\operatorname{MAP},T})\right).$ (17)
Plugging in (13) evaluated at $\vec{\theta}_{\operatorname{MAP},T}$ and integrating out with respect to $\vec{\theta}$ gives
$\begin{split}p(\mathbf{y}_{T}|\vec{c},\mathbf{X})&\approx\exp\left(\sum_{i\in T}\left(y_{i}\vec{\theta}_{\operatorname{MAP},T}^{T}\Phi_{i\cdot}-\log(1+\exp(\vec{\theta}_{\operatorname{MAP},T}^{T}\Phi_{i\cdot}))\right)+\frac{\|\vec{\theta}_{\operatorname{MAP},T}\|_{2}^{2}}{2\gamma^{2}}\right)\\ &\quad\cdot(2\pi)^{(K-1)/2}\gamma^{-1}\det(H_{T}(\vec{\theta}_{\operatorname{MAP},T}))^{-1/2}.\end{split}$ (18)
We finish by applying the above formulas with respect to $T=[n]$ and $T=S$ and taking ratios:
$\begin{split}&p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})\\ &=\frac{p(\mathbf{y}|\vec{c},\mathbf{X})}{p(\mathbf{y}_{S}|\vec{c},\mathbf{X})}\\ &\approx\exp\left(\sum_{i=1}^{n}\left(y_{i}\vec{\theta}_{\operatorname{MAP}}^{T}\Phi_{i\cdot}-\log(1+\exp(\vec{\theta}_{\operatorname{MAP}}^{T}\Phi_{i\cdot}))\right)-\sum_{i\in S}\left(y_{i}\vec{\theta}_{\operatorname{MAP},S}^{T}\Phi_{i\cdot}-\log(1+\exp(\vec{\theta}_{\operatorname{MAP},S}^{T}\Phi_{i\cdot}))\right)\right)\\ &\quad\cdot\exp\left(\frac{\|\vec{\theta}_{\operatorname{MAP},S}\|_{2}^{2}-\|\vec{\theta}_{\operatorname{MAP}}\|_{2}^{2}}{2\gamma^{2}}\right)\left(\frac{\det(H_{S}(\vec{\theta}_{\operatorname{MAP},S}))}{\det(H(\vec{\theta}_{\operatorname{MAP}}))}\right)^{1/2}.\end{split}$ (19)
Remark G.1.
The MAP estimate $\vec{\theta}_{\operatorname{MAP}}$ can be computed by a call to scikit-learn’s LogisticRegression() model, setting C $=2\gamma^{2}$.
Remark G.2.
In the current implementation of our procedure, we use a slightly different method to perform Laplace approximation, which instead approximates the integral in (3).
Appendix H Proof of Theorem 3.1
Recall the definition
$L(\vec{c})\coloneqq\max_{\vec{\theta}}\mathbb{E}_{(X,Y)\sim\nu}\{\log p(Y|X,\vec{\theta},\vec{c})\}.$ (20)
This quantifies how relevant the concepts in $\vec{c}$ are to the prediction of the response $Y$. Note that the definition does not require that the model class is correctly specified (i.e. there does not have to be some $\vec{c}$ and $\vec{\theta}$ so that (1) holds). Before presenting the proof, we first detail the assumptions we make.
Assumption H.1.
  1. (i)
    $Y\in\{0,1\}$ is a binary response;
  2. (ii)
    The data $\mathcal{D}$ is generated i.i.d. from some distribution $\nu$;
  3. (iii)
    The prior on $\vec{\theta}$ is $\mathcal{N}(0,\gamma^{2}I)$;
  4. (iv)
    $\mathcal{C}^{*}\coloneqq\operatorname{argmax}_{\vec{c}}L(\vec{c})$ comprises all permutations of a single vector $\vec{c}^{~{}*}$;
  5. (v)
    There exists some $\Delta>0$ such that $L(\vec{c}^{~{}*})-\operatorname{argmax}_{\vec{c}\notin\mathcal{C}^{*}}L(\vec{c})\geq\Delta$;
  6. (vi)
    Let $\mathcal{C}_{n}\coloneqq\cup_{\vec{c}_{-k},\mathbf{y},\mathbf{X}}\{c\colon Q(c;\vec{c}_{-k},\mathbf{y},\mathbf{X})>0\}$. Assume $\mathcal{C}_{n}$ is a finite set of size at most $\exp(n^{1-\epsilon})$ for some $10<\epsilon<$ 1;
  7. (vii)
    There exists $\eta>0$ such that for each $\vec{c}_{-k}$, there exists $l\in\{1,2,\ldots,K\}$ such that $c_{l}^{*}=\operatorname{argmax}_{c}L((c,\vec{c}_{-k}))$, $c_{l}^{*}\notin\vec{c}_{-k}$, and $Q(c^{*}_{l};\vec{c}_{-k},\mathbf{y},\mathbf{X})\geq\eta$ almost surely for all $\mathbf{y},\mathbf{X}$ with $n$ large enough;
  8. (viii)
    There exists some $B$ such that for any $\vec{c}\in\mathcal{C}_{n}^{K}$, all distinct, we have $\|\vec{\theta}_{\vec{c}}^{*}\|_{2}\leq R$, where
    $\vec{\theta}_{\vec{c}}^{*}\coloneqq\operatorname{argmax}_{\vec{\theta}}\mathbb{E}_{(X,Y)\sim\nu}\{\log p(Y|X,\vec{\theta},\vec{c})\};$ (21)
  9. (ix)
    Let $\Phi_{\vec{c}}=(\phi_{c_{1}}(x),\phi_{c_{2}}(x),\ldots,\phi_{c_{K}}(x))$ for $x\sim\nu$. There exists some $\lambda_{\min}$ such that for any $\vec{c}\in\mathcal{C}_{n,K}$, all eigenvalues of the second moment matrix $\mathbb{E}[\Phi_{\vec{c}}\Phi_{\vec{c}}^{T}]$ are bounded from below by $\lambda_{\min}$. Here, $\mathcal{C}_{n,K}$ denotes the $K$-fold product of $\mathcal{C}_{n}$ but with any vectors containing duplicate concepts removed;
  10. (x)
    We modify the algorithm so that the random subset $S$ can only take one of $\exp(n^{1-\epsilon})$ possibilities $\mathcal{S}_{n}\coloneqq\{S_{1},S_{2},\ldots,S_{B,n}\}$.
Some of these assumptions can be relaxed and are made primarily for the sake of clarity of exposition. First, while we have focused on binary $Y$, the results readily extends to multiclass and continuous $Y$ as well. Second, the Gaussian prior assumption for $\vec{\theta}$ can be relaxed to one that simply assumes some smoothness and coverage properties. Note that for simplicity, we avoid discussing the constant term $\theta_{0}$.
Proof of Theorem 3.1.
We prove the theorem only for SS-MH-Update for simplicity, but it will be clear how to generalize it to Multi-SS-MH-Update.
Step 1: Taking a subsequence. Note that the Markov chain in question corresponds to the list $L$ returned in line 7 of Algorithm 1. Denote this using $(\tilde{\vec{c}}^{~{}t})_{t=1}^{\infty}$. Because of the cyclic nature of the updates, this chain is not time invariant and is hence difficult to work with. We argue that it suffices to study the temporal projection $(\vec{c}^{~{}t})_{t=1}^{\infty}$, where $\vec{c}_{t}=\tilde{\vec{c}}_{tK}$ for $t=0,1,2,\ldots$. This is because every stationary distribution for $(\tilde{\vec{c}}^{~{}t})_{t=1}^{\infty}$ is also stationary for $(\vec{c}^{~{}t})_{t=1}^{\infty}$.
Step 2: Reduction to connected components. Since $(\tilde{\vec{c}}^{~{}t})_{t=1}^{\infty}$ satisfies the detailed balance equations, it is a symmetric Markov chain, i.e. it has an equivalent representation as a flow on a network, in which the vertices are states and the edges correspond to nonzero transition probabilities. This property is inherited by $(\vec{c}^{~{}t})_{t=1}^{\infty}$. Consider a connected component $\mathcal{A}$ of the network. The Markov chain, when restricted to $\mathcal{A}$, is aperiodic, since there must exist a self-loop. Furthermore, it is finite according to Assumption H.1(vi). As such, $(\vec{c}^{~{}t})_{t=1}^{\infty}$ has a unique stationary distribution $\pi_{\mathcal{A}}$ supported on $\mathcal{A}$ (LevinPeresWilmer2006). Indeed, all stationary distributions of $(\vec{c}^{~{}t})_{t=1}^{\infty}$ are convex combinations of such distributions. Hence, it suffices to show the desired property for $\pi_{\mathcal{A}}$.
Step 3: Concentration. In order to analyze $\pi_{\mathcal{A}}$, we need quantitative control over the transition probabilities via concentration. For a given dataset $\mathcal{D}$ of size $n$, we condition on the event guaranteed by Proposition H.3.
Step 4: Convergence. We again focus on a single connected component $\mathcal{A}$ and the goal is to show that with high probability, $\pi_{\mathcal{A}}(\mathcal{C}^{*}\cap\mathcal{A})\geq 1-\delta$ for some $\delta\to 0$ uniformly. We may ignore the rest of the network. Let $a=|\mathcal{A}|$ and $b=|\mathcal{C}^{*}\cap\mathcal{A}|$. Since $\mathcal{A}$ itself is finite, we can enumerate its states starting with those in $\mathcal{C}^{*}\cap\mathcal{A}$. The stationary distribution can be written as a vector $v=(v_{1:b},v_{b+1:a})$, while the Markov chain itself can be represented as a transition matrix $P$, which we write in block form as
$P=\begin{bmatrix}P_{11}&P_{12}\\ P_{21}&P_{22}\end{bmatrix}.$ (22)
The stationarity equation gives us
$v_{b+1:a}(I-P_{22})=v_{1:b}P_{12},$ (23)
from which we obtain
$\|v_{b+1:a}\|_{2}=\left\|v_{1:b}P_{12}(I-P_{22})^{-1}\right\|_{2}\leq\frac{\|P_{12}\|_{op}}{1-\|P_{22}\|_{op}},$ (24)
so long as the denominator on the right hand side is nonzero. To see this, notice that any state $j$ for $b+1\leq j\leq a$ corresponds to a vector $\tilde{\vec{c}}^{~{}(0)}\notin\mathcal{C}^{*}$. However, using Assumption H.1(vii), there is a sequence $(\tilde{\vec{c}}^{~{}(0)},\tilde{\vec{c}}^{~{}(1)},\ldots,\tilde{\vec{c}}^{~{}(K-1)})$ such that for $k=1,\ldots,K$, we have (i) $\tilde{\vec{c}}^{~{}(k)}_{-k}=\tilde{\vec{c}}^{~{}(k-1)}_{-k}$, (ii) $Q(\tilde{c}_{k};\tilde{\vec{c}}^{~{}(k-1)}_{-k},\mathbf{y}_{S},\mathbf{X})\geq\eta$, (iii) $L(\tilde{\vec{c}}^{~{}(k)})\geq L(\tilde{\vec{c}}^{~{}(k-1)})$ with equality holding iff $\tilde{\vec{c}}^{~{}(k)}=\tilde{\vec{c}}^{~{}(k-1)}$. Combining (iii) with (32), we see that all acceptance ratios along this sequence are equal to 1, which means that the transition probability from state $j$ to states in $\mathcal{C}^{*}$ is at least
$\prod_{k=1}^{K}Q(\tilde{c}_{k};\tilde{\vec{c}}^{~{}(k-1)}_{-k},\mathbf{y}_{S},\mathbf{X})\geq\eta^{K}.$ (25)
The sum of entries in row $j-b$ in $P_{22}$ is the total probability of transition from state $j$ to other states within $\mathcal{A}\backslash\mathcal{C}^{*}$ and hence has value less than $1-\eta^{K}$. This holds for any row. Applying Lemma H.2, we thus get $\|P_{22}\|_{op}\leq 1-\eta^{K}$, which allows us to bound the denominator in (24). Next, every entry in $P_{12}$ is the probability of transition from a state in $\mathcal{C}^{*}$ to a state in $\mathcal{A}\backslash\mathcal{C}^{*}$. Using Step 3 and Assumption H.1(v), the acceptance probability $\alpha$ of such a transition satisfies
$\begin{split}\log\alpha&\leq\max_{\vec{c}\in\mathcal{A}\backslash\mathcal{C}^{*}}\log p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})-\log p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c}^{~{}*},\mathbf{X})\\ &\leq-n\Delta+O(n^{1-\epsilon^{\prime}}).\end{split}$ (26)
We thus have
$\|P_{12}\|_{op}\leq\sqrt{(a-b)b}\exp\left(-n\Delta(1+O(n^{-\epsilon^{\prime}}))\right),$ (27)
which allows us to bound the numerator in (24). Combining this with the earlier bound on the denominator gives
$\|v_{a+1:c}\|_{2}\leq\sqrt{(a-b)b}\eta^{-K}\exp\left(-n\Delta(1+O(n^{-\epsilon^{\prime}}))\right).$ (28)
We then compute
$\begin{split}\pi_{\mathcal{A}}(\mathcal{C}^{*}\cap\mathcal{A})&=\|v_{1:b}\|_{1}\\ &=1-\|v_{b+1:a}\|_{1}\\ &\geq 1-\sqrt{a-b}\|v_{a+1:c}\|_{2}\\ &\geq 1-\sqrt{b(a-b)^{2}}\eta^{-K}\exp\left(-n\Delta(1+O(n^{-\epsilon^{\prime}})\right),\end{split}$ (29)
whose error term we can bound as
$\begin{split}\sqrt{b(a-b)^{2}}\eta^{-K}\exp\left(-n\Delta(1+O(n^{-\epsilon^{\prime}})\right)&\lesssim|\mathcal{C}_{n}|\exp\left(-n\Delta(1+O(n^{-\epsilon^{\prime}})\right)\\ &\leq\exp(n^{1-\epsilon})\exp\left(-n\Delta(1+O(n^{-\epsilon^{\prime}})\right)\\ &\leq\exp(O(n^{-\epsilon})).\end{split}$ (30)
Finally, we note that all hidden constants can be chosen to be independent of $n$. ∎
Lemma H.2.
Let $M$ be a $m\times n$ matrix whose entries are non-negative, and suppose $\sum_{j=1}^{n}M_{ij}\leq\beta$ for $i=1,2,\ldots,m$. Then $\|M\|_{op}\leq\beta$.
Proof.
Let $D$ be the $m\times m$ diagonal matrix whose entries are given by $D_{ii}=\sum_{j=1}^{n}M_{ij}$. Then $D^{-1}M$ is a stochastic matrix, which has operator norm at most 1, while $\|D\|_{op}\leq\max_{1\leq i\leq n}\sum_{j=1}^{n}M_{ij}\leq\beta$. Putting these together, we get
$\|M\|_{op}=\|DD^{-1}M\|_{op}\leq\|D\|_{op}\|D^{-1}M\|_{op}\leq\beta,$ (31)
as we wanted. ∎
Proposition H.3 (Concentration of split-sample posterior).
Assuming the conditions in Assumption H.1, then for any $\epsilon^{\prime}<\epsilon/2$, we have
$\sup_{\vec{c}\in\mathcal{C}_{n,K},S\in\mathcal{S}_{n}}\left|\frac{1}{\lceil(1-\omega)n\rceil}\log p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})-L(\vec{c})\right|=O(n^{-\epsilon^{\prime}})$ (32)
with probability converging to 1.
Proof.
Let us fix $\vec{c}$ for now. We first write
$\begin{split}l_{i}(\vec{\theta})&\coloneqq\log p(Y=y_{i}|X=\mathbf{x}_{i},\vec{\theta},\vec{c})\\ &=y_{i}\vec{\theta}^{T}\Phi_{i\cdot}-\log\left(1+\exp(\vec{\theta}^{T}\Phi_{i\cdot})\right).\end{split}$ (33)
Let $\bar{l}(\vec{\theta})\coloneqq\mathbb{E}_{\mathbf{x}_{i},y_{i}}[l_{i}(\vec{\theta})]$ and set $\vec{\theta}^{*}=\operatorname{argmax}\bar{l}(\vec{\theta})$. For convenience, denote $\Phi=\Phi_{\vec{c}}$. Note that $\Phi$ has entries bounded between 0 and 1 and hence has a second moment matrix with maximum eigenvalue bounded above by $K$. Furthermore, it is a sub-Gaussian random vector with bounded sub-Gaussian norm. Using these properties, we may find a universal value $R^{\prime}$ such that $\mathbb{E}\left[\Phi\Phi^{T}\mathbf{1}(\|\Phi\|_{2}\leq R^{\prime})\right]\succeq\frac{1}{2}\mathbb{E}\left[\Phi\Phi^{T}\right]$. Consider the ball $B(2R)$ of radius $2R$. For any $\vec{\theta}\in B(2R)$, the Hessian of the population log likelihood hence satisfies
$\begin{split}H(\theta)&\coloneqq\nabla^{2}\bar{l}(\vec{\theta})\\ &=-\mathbb{E}\left[\frac{\exp(\Phi^{T}\vec{\theta})}{\left(1+\exp(\Phi^{T}\vec{\theta})\right)^{2}}\Phi\Phi^{T}\right]\\ &\preceq-\mathbb{E}\left[\frac{\exp(\Phi^{T}\vec{\theta})}{\left(1+\exp(\Phi^{T}\vec{\theta})\right)^{2}}\Phi\Phi^{T}\mathbf{1}(\|\Phi\|_{2}\leq M)\right]\\ &\preceq-\frac{1}{2}\inf_{\vec{\theta}\in B(2R),\Phi\in B(M)}\frac{\exp(\Phi^{T}\vec{\theta})}{\left(1+\exp(\Phi^{T}\vec{\theta})\right)^{2}}\mathbb{E}\left[\Phi\Phi^{T}\right]\\ &=\frac{e^{2R^{\prime}R}}{2(1+e^{R^{\prime}R})^{2}}\mathbb{E}\left[\Phi\Phi^{T}\right]\\ &\preceq-\frac{\lambda_{\min}e^{2R^{\prime}R}}{2(1+e^{R^{\prime}R})^{2}}I.\end{split}$ (34)
Using this lower bound on the curvature, we see that
$\bar{l}(\vec{\theta}^{*})-\sup_{\theta\in\partial B(2R)}\bar{l}(\vec{\theta})\geq\frac{\lambda_{\min}R^{2}e^{2R^{\prime}R}}{2(1+e^{R^{\prime}R})^{2}}.$ (35)
Next, we use the fact that $n^{-1/2}\sum_{i=1}^{n}\left(l_{i}(\vec{\theta})-\bar{l}(\vec{\theta})\right)$ is a sub-Gaussian process with increments bounded according to
$\left\|n^{-1/2}\sum_{i=1}^{n}\left(l_{i}(\vec{\theta})-\bar{l}(\vec{\theta})\right)-n^{-1/2}\sum_{i=1}^{n}\left(l_{i}(\vec{\theta}^{\prime})-\bar{l}(\vec{\theta}^{\prime})\right)\right\|_{\psi_{2}}\lesssim K\|\vec{\theta}-\vec{\theta}^{\prime}\|_{2}.$ (36)
We can thus use Talagrand’s comparison inequality (vershynin2018high) to get, for any $t>0$, that
$n^{-1/2}\sup_{\vec{\theta}\in B(2R)}\left|\sum_{i=1}^{n}\left(l_{i}(\vec{\theta})-\bar{l}(\vec{\theta})\right)\right|\lesssim R\sqrt{K}+Rt$ (37)
with probability at least $1-e^{-t^{2}}$. Take $t=n^{1/2-\epsilon^{\prime}}$ for any $\epsilon^{\prime}<\epsilon/2$. Note that $\hat{l}_{n}(\vec{\theta})\coloneqq n^{-1}\sum_{i=1}^{n}l_{i}(\vec{\theta})$ is concave, with
$\hat{l}_{n}(\vec{\theta}^{*})\geq\bar{l}(\vec{\theta}^{*})-O(n^{-\epsilon^{\prime}}),$ (38)
while
$\begin{split}\sup_{\vec{\theta}\in\partial B(2R)}\hat{l}_{n}(\vec{\theta})&\leq\sup_{\theta\in\partial B(2R)}\bar{l}(\vec{\theta})+O(n^{-\epsilon^{\prime}})\\ &\leq\bar{l}(\vec{\theta}^{*})+\frac{\lambda_{\min}R^{2}e^{2R^{\prime}R}}{2(1+e^{R^{\prime}R})^{2}}+O(n^{-\epsilon^{\prime}}).\end{split}$ (39)
As such, whenever $n$ is large enough, we see that $\hat{l}_{n}$ also attains its global maximum in $B(2R)$ and that this maximum value satisfies
$\left|\sup_{\vec{\theta}\in\mathbb{R}^{K}}\hat{l}_{n}(\vec{\theta})-\bar{l}(\vec{\theta}^{*})\right|=O(n^{-\epsilon^{\prime}}).$ (40)
Recall further that $\bar{l}(\vec{\theta}^{*})=L(\vec{c})$. We next want to integrate out $p(\mathbf{y}|\vec{\theta},\vec{c},\mathbf{X})=\exp(n\hat{l}_{n}(\vec{\theta}))$ with respect to the prior. Using (40), we get the upper bound
[…]
To get a lower bound, we use the fact that the Hessian of the population likelihood also has an lower bound
$H(\theta)\succeq-\|h(\theta)\|_{op}I\succeq-\sqrt{K}I.$ (42)
For any $0<r<R$, this gives a lower bound of
$\inf_{\vec{\theta}\in B(\vec{\theta}^{*},r)}L(\vec{\theta})\geq L(\vec{\theta}^{*})-\sqrt{K}r^{2},$ (43)
where $B(\vec{\theta}^{*},r)$ is the ball of radius $r$ centered at $\vec{\theta}^{*}$. Using (37) again, we get
$\inf_{\vec{\theta}\in B(\vec{\theta}^{*},r)}\hat{l}_{n}(\vec{\theta})\geq L(\vec{\theta}^{*})-\sqrt{K}r^{2}-O(n^{-\epsilon^{\prime}}).$ (44)
The prior density is bounded from below by $\frac{1}{(2\pi\gamma^{2})^{K/2}}\exp(-2R^{2}/\gamma^{2})$ on $B(2R)$. We therefore have
[…]
The maximum is achieved when we set $r=\frac{K^{1/4}}{(2n)^{1/2}}$. Combining (41) and (45) and taking logarithms, we get
$\log p(\mathbf{y}|\vec{c},\mathbf{X})=nL(\vec{c})+O(n^{1-\epsilon^{\prime}}).$ (46)
Performing the same calculation with $\mathbf{y}_{S}$ instead of $\mathbf{y}$ and taking differences, we get
$\left|\frac{1}{\lceil(1-\omega)n\rceil}\log p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})-L(\vec{c})\right|=O(n^{-\epsilon^{\prime}}).$ (47)
Finally, we take a union bound over all $\vec{c}\in\mathcal{C}_{n,K}$ and $S\in\mathcal{S}_{n}$, noting that all hidden constants can be chosen to be independent of $\vec{c}$. ∎
Appendix I Stationary Distribution Under Consistent Proposals
In this section, we motivate the multiple-try partial posterior Metropolis update (Algorithm 3) in more detail. Notably, we have the following proposition.
Proposition I.1.
Assuming equality holds in (2), the Markov chain defined by running Gibbs sampling (Algorithm 1) with SS-MH-Update or Multi-SS-MH-Update instead of MH-Update has the posterior $p(\vec{C}|\mathbf{y},\mathbf{X})$ as a stationary distribution.
Proof.
We first show this for SS-MH-Update. It suffices to show that the acceptance probability in Line 4 of Algorithm 2 is equal to that of a Metropolis-Hastings filter with $p(\vec{C}|\mathbf{y},\mathbf{X})$ as the target. We compute:
$\begin{split}p((c,\vec{c}_{-k})|\mathbf{X},\mathbf{y})&=\frac{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},(c,\vec{c}_{-k}),\mathbf{X})p(\mathbf{y}_{S},(c,\vec{c}_{-k})|\mathbf{X})}{p(\mathbf{y}|\mathbf{X})}\\ &=p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},(c,\vec{c}_{-k}),\mathbf{X})p(c|\vec{c}_{-k},\mathbf{X},\mathbf{y}_{S})\frac{p(\vec{c}_{-k},\mathbf{y}_{S}|\mathbf{X})}{p(\mathbf{y}|\mathbf{X})}\end{split}$ (48)
Hence,
$\begin{split}\frac{p((\check{c},\vec{c}_{-k})|\mathbf{X},\mathbf{y})Q(c;\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})}{p((c,\vec{c}_{-k})|\mathbf{y},\mathbf{X})Q(\check{c};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})}&=\frac{p((\check{c},\vec{c}_{-k})|\mathbf{X},\mathbf{y})p((c,\vec{c}_{-k})|\mathbf{X},\mathbf{y})}{p((c,\vec{c}_{-k})|\mathbf{y},\mathbf{X})p((\check{c},\vec{c}_{-k})|\mathbf{X},\mathbf{y})}\\ &=\frac{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},(\vec{c}_{-k},\check{c}),\mathbf{X})}{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})}\end{split}$ (49)
as we wanted.
To show the statement for Multi-SS-MH-Update, we use its equivalence, for a fixed $S$, to the modified Multiple-Try Metropolis-Hastings method described in Appendix A. In Appendix A, we prove that this method satisfies the detailed balance equations for the posterior. Since the detailed balance equations are linear in the transition probabilities, we can marginalize this over $S$ to show that the posterior is stationary with respect for Multi-SS-MH-Update, which draws $S$ uniformly at random. ∎

### v2

@@
 The standard training procedure for CBMs is to predefine a candidate set of human-interpretable concepts, extract their values from the training data, and identify a sparse subset as inputs to a transparent prediction model.
-However, such approaches are often hampered by the tradeoff between enumerating a sufficiently large set of concepts to include those that are truly relevant versus controlling the cost of obtaining concept extractions.
+However, such approaches are often hampered by the tradeoff between exploring a sufficiently large set of concepts versus controlling the cost of obtaining concept extractions, resulting in a large interpretability-accuracy tradeoff.
 This work investigates a novel approach that sidesteps these challenges: BC-LLM iteratively searches over a potentially infinite set of concepts within a Bayesian framework, in which Large Language Models (LLMs) serve as both a concept extraction mechanism and prior.
-BC-LLM is broadly applicable and multi-modal.
-Despite imperfections in LLMs, we prove that BC-LLM can provide rigorous statistical inference and uncertainty quantification.
-In experiments, it outperforms comparator methods including black-box models, converges more rapidly towards relevant concepts and away from spuriously correlated ones, and is more robust to out-of-distribution samples.
+Even though LLMs can be miscalibrated and hallucinate, we prove that BC-LLM can provide rigorous statistical inference and uncertainty quantification.
+Across image, text, and tabular datasets, BC-LLM outperforms interpretable baselines and even black-box models in certain settings, converges more rapidly towards relevant concepts, and is more robust to out-of-distribution samples.
+[^1] Code for running BC-LLM and reproducing results in the paper are available at https://github.com/jjfeng/bc-llm.
 1 Introduction
-Although machine learning (ML) algorithms have demonstrated remarkable predictive performance, many lack the interpretability and transparency necessary for human experts to audit their accuracy, fairness, and safety (Dwork et al.,, 2012; Rudin,, 2018).
-This has limited their adoption in high-stakes applications such as medicine (Thirunavukarasu et al.,, 2023) and settings where regulatory agencies require algorithms to be explainable (Meskó and Topol,, 2023).
-Recent works have explored Concept Bottleneck Models (CBMs) (Koh et al.,, 2020; Kim et al.,, 2023) as a potential solution: these methods leverage black-box algorithms to extract a small number of interpretable concepts, which are subsequently processed by a fully transparent tabular model to make predictions.
-This allows users to understand how exactly each concept contributes to the final prediction and, if necessary, modify the extracted concepts to fix model predictions.
-However, there are major limitations in the way CBMs are currently trained.
-They require human experts to identify a large pool of concepts a priori and annotate the concepts present in each observation.
-This heavy dependence on human experts is expensive and often impractical.
-More importantly, it is difficult to ensure this pool of concepts includes the truly relevant concepts.
-Consequently, resulting CBMs are often outperformed by their black-box counterparts (Yuksekgonul et al.,, 2022) and may even mislead users as to which concepts are truly relevant (Ramaswamy et al.,, 2023).
-Consider the task of predicting hospital readmission from patient notes.
-It is impossible to prespecify all possible related concepts.
-Not only are there many different concepts, but any single concept can also be refined.
-For instance, for the concept regarding the “patient’s smoking status,” there is the more general concept regarding “substance use,” the correlated concept of “whether a patient has quit smoking,” and many more.
-While iterative refinement of candidate concepts may help, it would make learning a CBM even more burdensome.
-LLMs have the potential to address these challenges.
-They are cheaper than human experts, can provide (relatively) high-quality concept annotations, and have sufficient world knowledge to hypothesize useful concepts.
-However, applying LLMs as a simple stand-in for a human expert does not fully address the challenges of concept discovery.
-First, LLM queries, though cheaper than human experts, still have a non-zero cost, which practically limits the number of concepts one can explore.
-Second, searching over concepts using LLMs introduces a “garden of forking paths” and may lead to the selection of spuriously correlated concepts (Gelman and Loken,, 2013).
-Finally, LLMs are imperfect query engines that can hallucinate, suggest bad candidate concepts, provide incorrect concept annotations, and may not even be self-consistent in their prior beliefs (Wang et al.,, 2023).
-As such, current methods that rely on LLMs to learn CBMs have struggled to achieve good performance, particularly in areas where the LLM’s prior knowledge is not well-aligned with the truth (Ludan et al.,, 2023; Benara et al.,, 2024).
-Figure 1: Overview of Bayesian inference for learning concept bottleneck models with LLM priors (BC-LLM).
-To address the aforementioned challenges, we show how wrapping LLMs within a Bayesian framework allows one to iteratively explore concepts in a way that is statistically principled and provides uncertainty quantification.
-An LLM can serve multiple purposes in this framework: it defines a prior over a potentially infinite set of candidate concepts, proposes concepts to explore during posterior sampling, and annotates observations with the candidate concepts.
-The procedure, which we refer to as BC-LLM, is designed to be computationally efficient; at each iteration, the LLM proposes a batch of candidate concepts so their respective annotations for each observation can be batched into a single LLM query.
-We prove that asymptotically, BC-LLM converges to the correct concepts even when the LLM prior is inconsistent or incompatible.
-Moreover, BC-LLM is highly generalizable and works across multiple modalities (text, images, and tabular data).
-Experiments across multiple datasets show that BC-LLM outperforms comparator methods and often even black-box models, converges more rapidly towards the true concepts and away from spuriously correlated ones, and is more robust to out-of-distribution (OOD) samples.
-Finally, when BC-LLM assisted a real-world hospital’s data science team to revise an existing tabular ML model, clinicians found the resulting CBM to be substantially more interpretable and actionable than the prior model.
+Although machine learning (ML) algorithms have demonstrated remarkable predictive performance, many lack the interpretability and transparency necessary for human experts to audit their accuracy, fairness, and safety [1, 2].
+This has limited their adoption in high-stakes applications such as medicine [3] and settings where regulatory agencies require algorithms to be explainable [4].
+Recent works have explored Concept Bottleneck Models (CBMs) [5, 6] as a potential solution: these methods leverage black-box algorithms to extract a small number of interpretable concepts, which are subsequently processed by a fully transparent tabular model to predict a target label.
+Thus, in principle, CBMs are much safer for use in high-stakes applications because humans can audit and, if necessary, modify the extracted concepts to fix predictions.
+Nevertheless, CBMs have yet to fully realize this promise.
+Early CBM methods were difficult to scale because they relied on human experts to specify and annotate concepts on training data.
+Recent works have proposed using LLMs to suggest and even annotate concepts [7, 8, 9, 10], as LLMs are cheaper, often have sufficient world knowledge to hypothesize useful concepts, and can provide (relatively) high-quality concept annotations [9, 10].
+This introduces new problems such as LLM hallucinations and inconsistencies but, more critically, does not resolve the fundamental problem that standard CBMs sacrifice a significant amount of accuracy to gain interpretability [11].
+Recent proposals try to address this by encoding concepts using “soft” continuous values, such as through embedding similarity [8], rather than “hard” binary values.
+However, “soft” CBMs can leak information about the label that would otherwise not have been available [12, 13, 14, 15, 16]; this risk is particularly severe when concept extraction is jointly trained with the final label predictor [15].
+[^2] For instance, a human would fill in 0 when asked “does this bird have a gray breast” if there is no gray breast, but a “soft” CBMs may extract 0.2 if it sees a blue breast or gray head.
+This provides additional hints for downstream prediction.
+Because the definitions of soft concepts are ambiguous, humans can no longer easily interpret, audit, or intervene on such CBMs.
+The aim of this work is to minimize the accuracy gap without sacrificing the interpretability of the extracted concepts.
+To this end, we revisit a known pain point of CBMs: Current training procedures require a pool of candidate concepts to be identified a priori, but ensuring this pool contains all relevant concepts is difficult, particularly in domains that are yet to be scientifically well-understood and where there are an infinite number of potentially relevant concepts [17, 10].
+For instance, for the task of predicting hospital readmission, there are an infinite number of factors affecting a patient’s risk.
+Even a single concept like “smoking status” has an infinite number of refinements, such as the correlated concept of “whether a patient has quit smoking,” the broader concept of “substance use,” and more.
+As shown in prior work, when the candidate pool misses important concepts, the resulting CBMs may be inaccurate and even mislead users as to which concepts are truly relevant [18].
+Rather than specifying concepts upfront, we investigate iterative refinement of concepts in a CBM with the assistance of an LLM.
+This updates a classical modeling paradigm for the LLM era: prediction models were traditionally designed through the collaboration between domain experts and data scientists, in which they iteratively refine/engineer features for simple tabular models.
+LLMs can help significantly accelerate this co-design process and scale up the number of iterations, but they are also imperfect query engines that can hallucinate, suggest bad candidate concepts, incorrectly annotate concepts, and may not even be self-consistent in their prior beliefs [19].
+To set up guardrails, we “wrap” the LLM within a Bayesian posterior inference procedure to explore concepts with the help of an LLM in a statistically principled manner.
+This approach, which we refer to as Bayesian Concept bottleneck models with LLM priors (BC-LLM), offers the following benefits:
+  • BC-LLM finds relevant concepts, even in settings where there is little to no prior knowledge: Unlike prior works that soley rely on a prespecified set of candidate concepts, BC-LLM explores a potentially infinite set of concepts data-adaptively and converge towards the true ones.
+  • BC-LLM is cost-effective: BC-LLM selectively annotates $O(KT)$ concepts, where $K$ is the number of concepts in the CBM and $T$ is the number of outer loops; we often use only $K\approx 5$ and $T\approx 5$.
+In contrast, existing methods annotate hundreds or thousands of prespecified concepts [8].
+  • BC-LLM corrects LLM mistakes in a statistically rigorous manner to provide calibrated uncertainty quantification: Mistakes in the LLM’s prior are fully allowed and even expected, as they are corrected in the Metropolis accept/reject step.
+Moreover, BC-LLM provides uncertainty quantification of the selected concepts, which improves performance in both in-distribution and out-of-distribution (OOD) settings.
+We prove that BC-LLM converges to the correct concepts, asymptotically, even when the LLM prior is poor or fails to be self-consistent.
+The Bayesian framework also naturally protects against overfitting and overconfidence in selected concepts.
+  • BC-LLM substantially improve the interpretability-accuracy tradeoff in real-world datasets: Experiments across multiple datasets and modalities (text, images, and tabular data) show that BC-LLM selects human-interpretable concepts while outperforming comparator methods, often even black-box models.
+Moreover, when assisting a hospital’s data science team to revise an existing tabular ML model, clinicians found BC-LLM to be substantially more interpretable and actionable.
+Figure 1: BC-LLM is initialized by having the LLM hypothesize the top concepts based on keyphrases extracted from each observation (Step 0).
+The concepts are then iteratively refined by dropping a concept (Step 1), querying the LLM for candidate replacements (Step 2), annotating each observation with the candidate concepts using the LLM (Step 3), and determining which, if any, of the candidate concepts to accept (Step 4).
 2 Related work
-Prior CBM methods train on observations with hand-annotated (Koh et al.,, 2020; Kim et al.,, 2023; Yuksekgonul et al.,, 2023) and, more recently, LLM-annotated concepts (Patel et al.,, 2023; McInerney et al.,, 2023; Benara et al.,, 2024) for a set of prespecified concepts by human experts.
-These methods are known to be highly sensitive to the selected set of concepts (Ramaswamy et al.,, 2023).
-More recent works have suggested using LLMs to prespecify the list of concepts instead (Oikarinen et al.,, 2023; Yang et al.,, 2023), but this requires the LLM to have accurate prior knowledge about the supervised learning task.
-Consequently, recent works have hypothesized that taking an iterative approach can more effectively find a set of relevant concepts (Ludan et al.,, 2023; Liu et al., 2024a, ).
+Prior CBM methods train on observations with hand-annotated concepts [5, 6, 20] and, more recently, concepts annotated by an LLM/VLM [21, 9, 10].
+These methods are known to be highly sensitive to the selected set of concepts [18].
+More recent works have suggested using LLMs to prespecify the list of candidate concepts instead [7, 8, 22, 23], but this requires the LLM to have accurate prior knowledge about the supervised learning task.
+We note that many of these works also make use of “soft concepts” and therefore suffer from information leakage [12].
+To find relevant concepts in a more data-driven manner, recent works have proposed taking an iterative approach [17, 24].
 These methods employ a boosting-like approach, where the LLM is provided a set of misclassified examples and asked to add/revise concepts to improve performance.
 However, the LLM can typically only review a few observations due to limits on context length, so the resulting concept proposals are far from optimal.
-An alternative strategy is to conduct post-hoc distillation of a black-box model into interpretable concepts (Kim et al.,, 2018; Yeh et al.,, 2020; Zhao et al.,, 2024; Jiang et al.,, 2023), which can be used to learn an interpretable prediction model.
-Nevertheless, prior work has shown that the generated explanations are often not faithful to the black-box model, leading to poorer performance from the interpretable model (Doshi-Velez and Kim,, 2017).
-Another related line of work uses LLMs to describe conceptual differences between pairs of datasets  (Zhong et al.,, 2023, 2022; Dunlap et al.,, 2024).
+An alternative strategy is to conduct post-hoc distillation of a black-box model into interpretable concepts [25, 26, 27, 28, 29], which can be used to learn an interpretable prediction model.
+Nevertheless, prior work has shown that the generated explanations are often not faithful to the black-box model, leading to poorer performance from the interpretable model [30].
+Another related line of work uses LLMs to describe conceptual differences between pairs of datasets  [31, 32, 33].
 However, existing methods are primarily suited for simple classification tasks, as they essentially learn a CBM with a single concept rather than a set of concepts.
-Finally, recent works have considered combining LLMs with Bayesian techniques (Yang et al.,, 2024; Liu et al., 2024b, ; Liu et al., 2024c, ).
-Some have suggested that In-Context Learning (ICL) can be viewed as LLMs conducting Bayesian inference (Xie et al.,, 2021; Ye et al.,, 2024), though recent work has proven that ICL is not fully consistent with Bayesian inference (Falck et al.,, 2024).
-To our knowledge, this work is the first to leverage LLMs in a Bayesian framework for learning CBMs and show that rigorous statistical inference is feasible despite imperfections in LLM reasoning.
-3 Method
+Finally, recent works have considered combining LLMs with Bayesian techniques [34, 35, 36].
+Some have suggested that In-Context Learning (ICL) can be viewed as LLMs conducting Bayesian inference [37, 38], though recent work has proven that ICL is not fully consistent with Bayesian inference [39].
+In contrast, this work studies how we can iteratively discover new concepts to include using LLMs using a Bayesian inference procedure, accounting for imperfections in LLM reasoning.
+3 Method: BC-LLM
 Consider a dataset $\mathcal{D}=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$, where $x_{i}$ are model inputs and $y_{i}$ are labels.
@@
 A concept $c$ defines a binary- or real-valued function $\phi_{c}$ of a model input, e.g.
-“Does the doctor note describe an elderly patient?” where $1$=yes and 00=no.
+“Does the doctor note describe an elderly patient?” where 11=yes and 0=no.
 A CBM selects a fixed number of $K$ concepts $\vec{c}=(c_{1},\ldots,c_{K})$ and fits a parametric model for $y$ using the extractions $(\phi_{c_{1}}(x),\ldots,\phi_{c_{K}}(x))$.
 For simplicity of presentation, our discussion will focus on logistic regression (LR)-based CBMs, although other choices are possible.
-In this case, the data is modeled by
-$p\left(Y=1|X=x,\vec{\theta},\vec{c}\right)=\sigma\left(\sum_{k=1}^{K}\theta_{k}\phi_{c_{k}}(x)+\theta_{0}\right)$ (1)
-where $\sigma$ is the sigmoid function and $\vec{\theta}=(\theta_{0},\cdots,\theta_{K})$ are the parameters.
+In this case, the data is modeled by $p\left(Y=1|X=x,\vec{\theta},\vec{c}\right)=\sigma\left(\sum_{k=1}^{K}\theta_{k}\phi_{c_{k}}(x)+\theta_{0}\right)$ where $\sigma$ is the sigmoid function and $\vec{\theta}=(\theta_{0},\cdots,\theta_{K})$ are the parameters.
 We refer to $\vec{c}$ as the “support” of the model.
@@
 3.1 Review: Bayesian variable selection
-Learning CBMs shares many similarities with the more standard problem of learning sparse models in high-dimensional settings, where one must select among a finite number of features that have already been tabulated (Ishwaran et al.,, 2008; Castillo et al.,, 2015; Biswas et al.,, 2022).
-Here, we review the Bayesian solution for this more standard setting.
-In later sections, we describe extensions to the CBM setting, where there are a potentially infinite number of concepts and a cost associated with extracting each.
-[…]
+Learning CBMs shares many similarities with the more standard problem of learning sparse models in high-dimensional settings, where one must select among a finite number of features that have already been tabulated [40, 41, 42].
+Here, we review the Bayesian solution for this more standard setting and then discuss how it may be extended to the CBM setting.
+Given priors on $\vec{c}$ and $\vec{\theta}$, the goal of Bayesian inference for sparse models in high-dimensional settings is to sample from the posterior $p(\vec{\theta},\vec{c}|\mathbf{X},\mathbf{y})$, which allows for uncertainty quantification of the true support and coefficients.
+Posterior samples also describe the uncertainty at a new point $x_{0}$ via the posterior predictive distribution $p(y_{0}|x_{0},\mathbf{y},\mathbf{X},\vec{c})=\iint p(y_{0}|x_{0},\vec{\theta},\vec{c})p(\vec{\theta},\vec{c}|\mathbf{y},\mathbf{X})d\vec{\theta}d\vec{c},$ which can be viewed as combining the posterior samples into an ensemble model.
+Factorizing the posterior per $p(\vec{\theta},\vec{c}|\mathbf{y},\mathbf{X})=p(\vec{\theta}|\vec{c},\mathbf{y},\mathbf{X})p(\vec{c}|\mathbf{y},\mathbf{X})$, the first term describes the posterior for the model parameters and the second describes the posterior over selected variables.
+As the former can be readily obtained by classical Bayesian inference for low-dimensional settings (e.g., posterior inference for LR coefficients), our discussion will focus on the latter.
 Inference for $p(\vec{c}|\mathbf{y},\mathbf{X})$ is typically achieved through Gibbs sampling.
-For each epoch, Gibbs rotates through indices $k=1,\ldots,K$, during which it replaces the $k$th concept in the current iterate $\vec{c}$ by drawing from the posterior distribution for $C_{k}$ conditional on the other concepts $\vec{c}_{-k}$.
-When one is unable to sample from the conditional posterior, Metropolis-within-Gibbs sampling can be used instead (Gilks et al.,, 1995).
-Rather than replacing the $k$th concept, it proposes a candidate concept $\check{c}_{k}$ given the other concepts per some distribution $Q(C_{k};\vec{C}_{-k}=\vec{c}_{-k})$ and accepts the candidate concept with probability $\alpha$.
+For each outer loop, Gibbs rotates through indices $k=1,\ldots,K$, during which it replaces the $k$th concept in the current iterate $\vec{c}$ by drawing from the posterior distribution for $C_{k}$ conditional on the other concepts $\vec{c}_{-k}$.
+When one is unable to sample from the conditional posterior, Metropolis-within-Gibbs can be used instead [43].
+Rather than immediately replacing the $k$th concept, it proposes a candidate concept $\check{c}_{k}$ given the other concepts per some distribution $Q(C_{k};\vec{C}_{-k}=\vec{c}_{-k})$ and accepts the candidate concept with probability $\alpha$.
 This acceptance probability $\alpha$, delineated in Algorithm 1, depends on the relative posterior probabilities of the candidate and existing concepts and is carefully designed so that the posterior distribution is stationary with respect to the sampling procedure.
@@
 12:  else return $c_{k}$
-3.2 Bayesian variable selection with LLMs
-Given the Gibbs procedures described above, there are various ways to integrate LLMs for learning CBMs. To begin, we describe two naïve approaches:
-Approach 1: Propose with LLM “prior.” Let the LLM define a conditional prior over concepts, i.e.
-$p(C_{k}|\vec{c}_{-k})$, and propose concepts from this conditional prior, i.e.
-$Q(C_{k};\vec{c}_{-k})=p(C_{k}|\vec{c}_{-k})$.
-As concepts are generated in a data-agnostic fashion, the LLM may produce many irrelevant concepts that have low acceptance probabilities.
-So while Approach 1 may provide valid inference, it is inefficient and costly.
-Approach 2: Propose with LLM “posterior.” Let the LLM conduct posterior inference for the $k$th concept given all the data and all but the $k$th concept through ICL and let the LLM propose from this conditional posterior, i.e.
-$Q(C_{k};\vec{c}_{-k})=p(C_{k}|\vec{c}_{-k},\mathbf{y},\mathbf{X})$.
-This simplifies Metropolis-within-Gibbs to Gibbs, which always accepts the proposed concept.
-As LLMs are not perfect Bayesian inference engines (Falck et al.,, 2024) and may be miscalibrated, this may lead to the inclusion of inappropriate concepts.
-While efficient, Approach 2 leads to biased sampling.
-Instead, we propose to take the middle road, which we argue gives the best of both worlds.
-At each iteration, we first randomly split the data to construct subset $S$ of size $\lfloor\omega n\rfloor$ and its complement $S^{c}$, for some fraction $\omega$ (e.g.
+We extend Bayesian variable selection to learn CBMs, where the key difference is that the number of potential concepts is effectively infinite, which makes it difficult to formulate an appropriate prior over concepts $p(\vec{C})$ and proposal transition kernel $Q(C_{k};\vec{C}_{-k})$.
+Intuitively, it seems that LLMs can help perform both these tasks, but how to do so correctly and efficiently is surprisingly subtle.
+To motivate our choice, we first describe two naïve options:
+Update 1: Propose with the LLM’s “prior.” Given the working concepts $\vec{C}_{-k}=\vec{c}_{-k}$, we can prompt the LLM to propose new concepts, without revealing any data, e.g.
+“Propose an additional concept for predicting readmissions, given the already-existing concepts $\vec{c}_{-k}$”.
+[^3] If one believes the LLM’s prior knowledge about the outcome of interest may be misleading, we can mask this information, e.g.
+“Propose an additional concept for predicting some label $Y$, given the already-existing concepts $\vec{c}_{-k}$.” This can be viewed as treating the proposal distribution $Q(C_{k};\vec{C}_{-k}=\vec{c}_{-k})$ as sampling from the LLM’s conditional prior distribution $p(C_{k}|\vec{c}_{-k})$.
+Consequently, the MH acceptance ratio simplifies to $\frac{p(\mathbf{y}|(\vec{c}_{-k},\check{c}_{k}),\mathbf{X})}{p(\mathbf{y}|\vec{c},\mathbf{X})}$, which allows us to override proposed concepts that lead to poor model performance.
+Without access to data however, the LLM is likely to propose many irrelevant concepts from its prior, resulting in low acceptance probabilities and slow sampling.
+So while Update 1 may provide valid inference, it is inefficient and costly.
+Update 2: Propose with the LLM’s “posterior.” Leveraging LLMs’ ability to perform ICL [37, 38], another idea is to include all the data when prompting the LLM to propose new concepts.
+That is, we ask the LLM itself to “conduct posterior inference” and interpret the LLM’s proposal distribution $Q(C_{k};\vec{C}_{-k}=\vec{c}_{-k})$ as sampling from a conditional posterior distribution $p(C_{k}|\vec{c}_{-k},\mathbf{y},\mathbf{X})$ for some implicitly defined concept prior.
+This approach is risky; while LLMs may be able to approximate posterior inference for small datasets, this approximation breaks down for larger sample sizes [39].
+Furthermore, we can no longer override poor concepts with data as the MH acceptance ratio simplifies to 1, i.e.
+the proposed concept is always accepted.
+So while Update 2 may generate concepts more efficiently, the sampling procedure may not converge to a good or even valid posterior distribution.
+These update options are at the end of two extremes, trading off between efficient sampling versus unbiased posterior inference. Our task then is to find a better balance.
+3.2 The Split-sample update: Propose with the LLM’s partial posterior
+Rather than relying on the LLM’s prior or posterior, a key idea in BC-LLM is to take the middle road by sampling from an LLM’s partial posterior, which we will show enjoys the best of both worlds.
+The proposal, which we refer to as Split-sample Metropolis update (Algorithm 2), uses the LLM to propose candidate concepts based on only a portion of the data and fixes mistakes in the LLM’s proposal using the remaining data in the accept-reject step.
+This addresses the drawbacks of the two naïve updating procedures in Section 3.1, as it simultaneously uses the data to generate more efficient proposals while overriding the LLM as needed.
+More specifically, the split-sample update first randomly splits the data to construct subset $S$ of size $\lfloor\omega n\rfloor$ and its complement $S^{c}$, for some fraction $\omega$ (e.g.
 half).
-Then the LLM is asked to propose a candidate concept based on $\vec{c}_{-k}$ and subset $S$, which can be viewed as approximately sampling from the “partial posterior”
-$Q(C_{k};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})\approx p(C_{k}|\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X}).$ (2)
-Assuming (2) holds with equality, the Metropolis acceptance probability simplifies to
-$\alpha\leftarrow\min\left\{\frac{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},(\vec{c}_{-k},\check{c}),\mathbf{X})}{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})},1\right\}.$
-This essentially compares the likelihood of the held-out data with respect to the candidate and existing concepts, much like sample-splitting in frequentist settings.
-If the CBM with the candidate concept is better, it is accepted; otherwise it is accepted with probability $\alpha$.
-This procedure, which we refer to as Split-sample Metropolis-within-Gibbs (Algorithm 2) can learn from the data to override erroneous priors and concepts from the LLM, unlike Approach 2.
-In addition, it is efficient because the LLM will propose relevant concepts based on some of (though not all of) the data, unlike Approach 1.
+The LLM is prompted to generate candidate concepts by combining its prior knowledge with information from the data in $S$.
+We can then interpret the LLM’s proposal distribution $Q(C_{k};\vec{C}_{-k}=\vec{c}_{-k})$ as the (conditional) partial posterior distribution $p(C_{k}|\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})$.
+Then in the accept-reject step, the MH acceptance probability simplifies to the partial Bayes factor $\frac{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},(\vec{c}_{-k},\check{c}),\mathbf{X})}{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})}$ [44].
+This compares the likelihood of the held-out data $S^{c}$ with respect to the candidate and existing concepts—much like sample-splitting in frequentist settings—and thus provides an opportunity to correct inconsistencies between the LLM’s proposal and the actual posterior distribution.
 Algorithm 2 Split-sample Metropolis update
-1:function SS-MH-Update($\vec{c}$, $k$)
+1:function $S$-MH-Update($\vec{c}$, $k$)
 2:  Sample subset $S$ of size $\lfloor\omega n\rfloor$.
@@
 6:  else return $c_{k}$
-In practice, rather than sampling a single candidate concept, we use a version of multiple-try Metropolis-Hastings (Algorithm 3) (Liu et al.,, 2000).
-Here, the LLM proposes a batch of candidate candidates, and then we sample a candidate from this batch based on their posterior probabilities and consider the sampled candidate for acceptance.
-This is more cost-effective as it lets us batch concept annotations, as described later.
 Algorithm 3 Multiple-try split-sample Metropolis update
-1:function Multi-SS-MH-Update($\vec{c}$, $k$)
+1:function Multi-$S$-MH-Update($\vec{c}$, $k$)
 2:  Sample subset $S$ of size $\lfloor\omega n\rfloor$.
@@
 8:  else return $c_{k}$
-Assuming that the LLM’s partial posterior across all iterations is consistent with some prior distribution $p(\vec{C})$, we can show that the Markov chain defined by this procedure has the posterior $p(\vec{c}|\mathbf{y},\mathbf{X})$ as its stationary distribution (see Proposition I.1 in the Appendix).
-More importantly, we can prove that the procedure still converges to the true concepts even in the likely scenario where there is no valid prior.
+In practice, rather than sampling a single candidate concept, we use a version of multiple-try Metropolis-Hastings (Algorithm 3) [45].
+Here, the LLM proposes a batch of candidates.
+We sample a candidate from this batch based on their posterior probabilities and consider the sampled candidate for acceptance.
+This is more cost-effective as it lets us batch concept annotations, as described later.
+Assuming that the LLM’s partial posterior across all iterations is consistent with some prior distribution $p(\vec{C})$, we can prove that the Markov chain defined by this procedure has the posterior $p(\vec{c}|\mathbf{y},\mathbf{X})$ as its stationary distribution (see Proposition H.1 in the Appendix).
+Since LLMs are known to be imperfect Bayesian inference engines [39], this assumption is at best only approximately satisfied.
+Nevertheless, we can prove that the procedure still converges to the true concepts even when this assumption does not hold:
 Theorem 3.1.
@@
 Let $L(\vec{c})\coloneqq\max_{\vec{\theta}}\mathbb{E}\{\log p(Y|X,\vec{\theta},\vec{c})\}$ and $\mathcal{C}^{*}\coloneqq\operatorname{argmax}_{\vec{c}}L(\vec{c}).$ For sample size $n$, let $\Pi_{n}$ denote the set of stationary distributions of the Markov chain defined by running Algorithm 1 with SS-MH-Update or Multi-SS-MH-Update instead of MH-Update.
-Assuming a Gaussian prior for $\vec{\theta}$ and under regularity conditions (see Assumption H.1 in the Supplement), then $\inf\{\pi(\mathcal{C}^{*})\colon\pi\in\Pi_{n}\}\to 1$ in probability as $n\to\infty$.
-3.3 BC-LLM
-We now describe the main steps in BC-LLM (Fig 1): the procedure is initialized in Step 0, and each iteration (Steps 1 to 4) samples a CBM.
-The samples then describe the posterior distribution over CBMs and form an ensemble prediction model (after removing those from the burn-in period).
-Depending on one’s use case, the search over concepts can be guided by the LLM’s prior knowledge about the target $Y$ by injecting this information into the prompts (e.g.
-“we are trying to predict readmission…”) or in an “unbiased” fashion by masking the meaning of $Y$ (e.g.
-“we are trying to predict some label $Y$…”).
-Example prompts and other implementation details are provided in the Appendix.
-Step 0: Initialization and keyphrase extraction
-We first use an LLM to extract “keyphrases” describing each observation $x$, which will later be used by the LLM to brainstorm candidate concepts.
-Here, we use “keyphrases” to refer to short phrases that describe the observation: keyphrases for images describe what’s in the image (e.g.
-blue eyes), while keyphrases for text data can be phrases that appear in the text or summarize the main ideas (e.g.
-diabetes) (Papagiannopoulou and Tsoumakas,, 2020).
-Unlike concepts, keyphrases do not define a function that outputs a value, though they can help one come up with a concept (e.g.
-“diabetes” is a keyphrase whereas the binary questions “Does the patient currently have diabetes?” or “Does the patient have a family history of diabetes?” are concepts).
-In addition, this step does not restrict the set of keyphrases one can extract to a prespecified list, which allows for a wide exploration of keyphrases.
-The tradeoff is that the extracted keyphrases may be, and likely are, an incomplete summary of each observation.
-To initialize the CBM sampling procedure, these keyphrases are converted into numeric vectors to fit a (penalized) LR model for the target $Y$, similar to the process for fitting a bag-of-words model.
-An LLM analyzes the top keyphrases in this LR model to generate an initial set $\vec{c}$ of $K$ concepts.
-The LLM also extracts these concepts from each observation.
-Step 1: Drop a concept & fit a keyphrase model
-Each iteration proposes candidate concepts based on some data subset $S$ to replace the $k$th concept for $k\in\{1,\cdots,K\}$.
-Given the context limits of LLMs, it is unlikely that an LLM can generate a candidate concept from the partial posterior distribution by reviewing each observation in $S$ individually.
-As such, we will generate a summary by highlighting the top keyphrases in $S$ for predicting $Y$, on top of the existing $K-1$ concepts $\vec{c}_{-k}$.
+Assuming a Gaussian prior for $\vec{\theta}$ and under regularity conditions (see Assumption G.1), then $\inf\{\pi(\mathcal{C}^{*})\colon\pi\in\Pi_{n}\}\to 1$ in probability as $n\to\infty$.
+3.3 Putting it together: BC-LLM
+We now put all the pieces together.
+Mathematically, BC-LLM runs Metropolis-within-Gibbs (Algorithm 1) with a split-sample update (Algorithm 2 or 3).
+We translate this into a step-by-step recipe (Fig 1): After initializing BC-LLM in Step 0, each iteration of Metropolis-within-Gibbs (Steps 1 to 4) performs a split-sample update to output a CBM.
+The CBMs form a posterior distribution, as well as an ensemble prediction model.
+Example prompts and implementation details are in the Appendix.
+Step 0: Initialization and keyphrase extraction We first use an LLM to extract “keyphrases” from each observation $x$, which will be used by the LLM to brainstorm candidate concepts.
+Here, “keyphrases” refer to short phrases: keyphrases for images describe what’s in the image (e.g.
+blue eyes), while keyphrases for text data can be direct quotes or summarizations (e.g.
+diabetes) [46].
+Keyphrases can be thought of as rough concepts, as they do not define a function but can help an LLM come up with a formal concept question (e.g.
+“diabetes” is a keyphrase whereas “Does the patient currently have diabetes?” is a concept).
+This brainstorming phase does not restrict the set of keyphrases the LLM can extract to allow for a wide exploration of concepts.
+The tradeoff is that the extracted keyphrases may be an incomplete summary of each observation.
+To initialize the CBM sampling procedure, these keyphrases are used to fit a (penalized) LR model for the target $Y$, similar to fitting a bag-of-words (BoW) model.
+An LLM then analyzes the most predictive keyphrases to generate an initial set of $K$ concepts and extracts these concepts from each observation.
+Step 1: Drop a concept & fit a keyphrase model Each iteration proposes candidate concepts based on some data subset $S$ to replace the $k$th concept for $k\in\{1,\cdots,K\}$.
+Given the context limits of LLMs, having an LLM review each observation in $S$ individually is ineffective.
+Instead, we generate a summary by highlighting the top keyphrases in $S$ for predicting $Y$, on top of the existing K1K-1 concepts $\vec{c}_{-k}$.
 For continuous outcomes, this can be accomplished by finding the top keyphrases associated with the residuals of a model that predicts $Y$ given $\vec{c}_{-k}$.
-As the residuals are noisy for binary/categorical outcomes, we fit a (multinomial) LR model of $Y$ instead, where the input features are the keyphrases (again encoded using a bag-of-words approach) and annotated concepts $\vec{c}_{-k}$.
-The parameters of this “keyphrase model” $f$ are the coefficients for the keyphrases and concepts, denoted $\beta_{W}$ and $\beta_{\vec{c}_{-k}}$, respectively.
-Given the high-dimensionality of the keyphrases, we penalize $\beta_{W}$ using some penalty function $J$ and estimate $f$ by solving
-$\min_{\beta_{\vec{c}_{-k}},\beta_{W}}\frac{1}{n}\sum_{i=1}^{n}\ell\left(y_{i},f\left(x_{i}|\beta_{\vec{c}_{-k}},\beta_{W}\right)\right)+\lambda J(\beta_{W}),$
-where $\ell$ is the logistic loss and penalty parameter $\lambda\geq 0$ is selected by cross-validation.
-We found the ridge penalty (i.e., $J(\beta_{W})=\|\beta_{W}\|_{2}^{2}$) to work well in practice, though other choices for $J$ may be used instead.
-Step 2: Query the LLM for candidate concepts Given the summary of the top keyphrases generated from Step 1, the LLM is asked to propose $M$ candidates for the $k$-th concept from the partial posterior $p(C_{k}|\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})$.
+For binary/categorical outcomes, we instead fit a (multinomial) LR model of $Y$, where the inputs are keyphrases encoded using BoW and annotated concepts $\vec{c}_{-k}$.
+Let the coefficients in this “keyphrase model” $f$ for keyphrases and concepts be denoted $\beta_{W}$ and $\beta_{\vec{c}_{-k}}$, respectively.
+We fit it by solving $\min_{\beta_{\vec{c}_{-k}},\beta_{W}}\frac{1}{n}\sum_{i=1}^{n}\ell\left(y_{i},f\left(x_{i}|\beta_{\vec{c}_{-k}},\beta_{W}\right)\right)+\lambda J(\beta_{W})$, where $\ell$ is the logistic loss, $\beta_{W}$ is penalized with some penalty function $J$, and penalty parameter $\lambda$ is selected by cross-validation.
+We found the ridge penalty to work well in practice but other penalties may be used instead.
+Step 2: Query the LLM for candidate concepts Given the summary of the top keyphrases generated from Step 1, the LLM is asked to propose $M$ candidates for the $k$-th concept.
 In addition, we ask the LLM to provide the value of the partial posterior probability for each candidate concept.
 For instance, if the top keyphrases include “chronic kidney disease” and “acute kidney injury,” the LLM may propose a concept like “Does the patient have a history of kidney disease?” with an assigned partial posterior probability of 0.4.
-Step 3: Concept annotation Given the batch of candidate concepts, we use the LLM to extract their values for each observation $x$, i.e.
+Step 3: Concept annotation Given candidate concepts, we use the LLM to extract their values using a zero-shot approach, i.e.
 $\phi_{c}(x)$ for $c\in\{\check{c}^{(1)}_{k},\ldots,\check{c}^{(M)}_{k}\}$.
 Because multiple-try Metropolis-within-Gibbs proposes a batch of candidate concepts, we can batch their extractions into a single LLM query for each observation.
-This makes BC-LLM more cost/query-efficient than ordinary Metropolis-within-Gibbs, which outputs only a single concept per iteration and thus cannot batch extractions.
+This makes BC-LLM more cost/query-efficient than ordinary Metropolis-within-Gibbs, which outputs only a single concept per iteration and thus cannot be batched.
+Note that letting the LLM occasionally output probabilities when it is unsure can be helpful, though it is critical to avoid information leakage (e.g.
+no joint training of concept extraction and label prediction).
 Step 4: Accept/reject step The final step is to compute the sampling probabilities for the candidate concepts and the acceptance ratio, which involve comparing the proposal probabilities for the existing and candidate concept sets $\vec{c}$ as well as their split-sample posteriors
-[…]
+$p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})=\int p(\mathbf{y}_{S^{c}}|\vec{\theta},\vec{c},\mathbf{X})p(\vec{\theta}|\mathbf{y}_{S},\vec{c},\mathbf{X})d\vec{\theta}.$ (1)
 To estimate the integral (1), one approach is to sample from the posterior on $\vec{\theta}$, but this can be computationally expensive since one would need to implement posterior sampling in inner and outer loops.
-Alternatively, in the case of LR-based CBMs, one can use a Laplace-like approximation for (3) (Lewis and Raftery,, 1997), which is discussed in more detail in the Appendix.
-This essentially approximates the partial and full posterior distributions with Gaussian distributions.
+Alternatively, in the case of LR-based CBMs, one can use a Laplace-like approximation for (1) [47], which is discussed in more detail in the Appendix.
 As the posterior distribution of LR parameters are asymptotically normal, this approximation becomes increasingly accurate as $n\rightarrow\infty$.
-3.4 Computational cost
-The key computational cost of BC-LLM is from querying the LLM.
-Running BC-LLM requires $O(nTK)$ queries, where $T$ is the number of epochs.
-$K$ is typically small by design, because large $K$ leads to less interpretable CBMs.
-One may be concerned by $T$, as standard Bayesian inference procedures typically draw thousands of posterior samples.
-Nevertheless, combined with a greedy warm-start procedure, we found that a small $T$ to be quite effective; our experiments go as low as $T=5$.
-Generally speaking, if the goal is solely prediction accuracy, a small $T$ may suffice.
-If the goal is uncertainty quantification for relevant concepts, a larger $T$ (e.g., 10) is helpful for more complete exploration of concepts.
-Figure 2: MIMIC evaluation: (a) Comparison of BC-LLM and existing methods in terms of performance and recovery of true concepts.
-Error bars indicate the 95% CI.
-(b) Dendrograms of concepts learned by BC-LLM with 100 and 800 training observations (left and right, respectively).
-Labels are shortened concept questions, which generally have the format “Does the note mention the patient having…?” Highlighted labels correspond to the true, color-coded concepts in Section 4.2.
-Numbers in parentheses describe the proportion of posterior samples with the concept.
-4 Experimental results
-4.1 Experimental setup
-We compare BC-LLM against One-pass summarization, where an LLM performs post-hoc distillation of the keyphrase model into a CBM (Pham et al.,, 2023), and Boosting, where an LLM iteratively adds concepts to a CBM by analyzing observations with large residuals (Ludan et al.,, 2023; Liu et al., 2024a, ).
-We also compare it to black-box and semi-interpretable models to assess whether BC-LLM can close the performance gap.
-When the true concepts are known, we also fit an Oracle CBM.
-Methods were evaluated with respect to predictive performance as measured by AUC and uncertainty quantification as measured by log likelihood (higher is better).
-When true concepts are known, we also quantify concept selection rates using “concept precision,” as defined by $\frac{1}{K}\sum_{k=1}^{K}p(\hat{C}_{k}\notin\vec{c}^{*}\mid\mathbf{y},\mathbf{X})$, and “concept recall,” as defined by $\frac{1}{K}\sum_{k=1}^{K}p(c_{k}^{*}\notin\hat{\vec{C}}\mid\mathbf{y},\mathbf{X})$, where $\hat{\vec{C}}$ is the posterior distribution over concept sets from BC-LLM or a single set of concepts learned by a non-Bayesian procedure.
-Quantifying a “correct” concept match requires going beyond exact string matching, as there are many possible refinements for a single concept.
-For instance, for the concept “Does the note mention the patient smoking in the present or the past?”, there are rephrasings (“Does this note mention the patient currently smoking or having smoked?”), polar opposites (“Does the note not mention the currently smoking or having previously smoked?”), or closely overlapping concepts (“Does the note mention the patient currently smokes?”).
-To quantify if a learned concept was correct, we used the following rule: if the absolute value of the Pearson correlation for LLM annotations of two concepts exceeds 50%, the two concepts are sufficiently similar.
-We manually confirmed that any pair of learned and true concepts that passed this rule were always closely related.
-Experiments used GPT-4o-mini (OpenAI,, 2023) and a protected health information (PHI)-compliant version of GPT-4o for real-world clinical notes in Section 4.4.
-Additional implementation details and experimental settings are in the Appendix.
-4.2 Simulation study: Learning CBMs for clinical notes from MIMIC
-We first compared BC-LLM against existing methods in terms of predictive performance and concept selection rates.
-To assess how well the methods perform even when the LLM has no prior knowledge about the target, we simulated $Y$ using a LR model with five Social Determinants of Health as inputs, which were annotated from real-world patient notes in MIMIC-IV (Johnson et al.,, 2023; Ahsan et al.,, 2021):
+Computational cost.
+BC-LLM performs $O(nTK)$ LLM queries, where $T$ is the number of outer loops.
+$K$ is typically chosen to be small, because a large $K$ yields less interpretable CBMs.
+Although standard Bayesian procedures typically draw thousands of posterior samples, we found that a small $T$ was quite effective in practice when combined with a greedy warm-start procedure; our experiments all use $T=5$.
+Generally speaking, a small $T$ suffices for achieving high prediction accuracy, a larger $T$ (e.g., 10) is helpful for quantifying the uncertainty of relevant concepts.
+Critically, running BC-LLM is significantly more cost-effective than standard CBMs, which require $O(nW)$ queries where $W$ is the number of pre-specified concepts and must be large (typically hundreds or thousands) to avoid missing relevant concepts [8].
+In our experiments, BC-LLM was quick to run, as one iteration takes less than a minute for text data and a few minutes for image data on a normal laptop.
+4 Experiments
+We evaluated BC-LLM in across three domains and modalities: classifying birds in images (Section 4.1), simulated outcomes from clinical notes (Section 4.2), and readmission risk in real-world clinical data (Section 4.3).
+The experiments below used GPT-4o-mini [48] and Section 4.3 used a protected health information-compliant version of GPT-4o for real-world clinical notes.
+The Appendix includes results using other LLMs, implementation details, and experimental settings.
+Baselines.
+We compared against CBMs fit using the standard approach where humans select and annotate concepts (Human+CBM) [5], CBMs where an LLM brainstorms concepts after seeing which keyphrases are most associated with the label but without any iteration (LLM+CBM) [49, 10], and CBMs fit using a boosting algorithm that iteratively adds concepts by having an LLM analyze observations with large residuals (Boosting LLM+CBM) [17, 24].
+We include black-box and semi-interpretable models to assess the tradeoff between interpretability and accuracy.
+For the bird-classification dataset, we also compare to a CBM that uses LLM-suggested concepts rather than keyword-based concepts; we use the 370 LLM-suggested concepts from [8] and extract the values of these concepts with an LLM (LLM+CBM (No keyphrases)).
+Evaluation metrics.
+Methods were evaluated with respect to predictive performance as measured by AUC and/or accuracy and uncertainty quantification as measured by Brier score.
+When true concepts are known, we also quantify concept selection rates using “concept precision,” as defined by $\frac{1}{K}\sum_{k=1}^{K}p(\hat{C}_{k}\in\vec{c}^{*}\mid\mathbf{y},\mathbf{X})$, and “concept recall,” as defined by $\frac{1}{K}\sum_{k=1}^{K}p(c_{k}^{*}\in\hat{\vec{C}}\mid\mathbf{y},\mathbf{X})$, where $\hat{\vec{C}}$ is the posterior distribution over concept sets from BC-LLM or a single set of concepts learned by a non-Bayesian procedure.
+Concept matches were verified manually (see Appendix for details).
+4.1 CBMs for classifying bird images
+We first evaluate how well BC-LLM can learn concepts that differentiate between bird species within the same family in the standard CUB-birds image dataset [50].
+Grouping the 200 bird species into their respective families, this resulted in 37 prediction tasks.
+We considered family-based prediction tasks because there likely exists a parsimonious CBM within each bird family, in contrast to the task of predicting between all 200 species.
+To assess how well BC-LLM performs in settings with limited prior knowledge, we do not tell the LLM that we are predicting bird species.
+That is, we replace the labels with one-hot encodings and simply tell the LLM to find concepts for predicting $Y$.
+Consequently, the LLM must search over a larger space of concepts.
+The data is split 50/50 between training and testing.
+The number of concepts $K$ learned for each task was set to the number of classes, but no smaller than 4 and no greater than 10.
+For the black-box comparator, we fine-tuned the last layer of ResNet50 pre-trained on ImageNetV2 [51, 52].
+Human+CBM was trained on the 312 human-annotated features available in the CUB-birds dataset.
+Table 1: Performance of CBMs and black-box models (ResNet) for classifying bird species.
+In-distribution OOD
+Method Accuracy ($\uparrow$) AUC ($\uparrow$) Brier ($\downarrow$) Entropy ($\uparrow$)
+BC-LLM 0.680 (0.614, 0.747) 0.874 (0.840, 0.907) 0.428 (0.357, 0.500) 0.865 (0.693, 1.036)
+LLM+CBM 0.640 (0.573, 0.707) 0.810 (0.768, 0.853) 0.452 (0.377, 0.528) 0.663 (0.474, 0.852)
+Boosting LLM+CBM 0.538 (0.463, 0.614) 0.722 (0.673, 0.772) 0.577 (0.499, 0.654) 0.842 (0.630, 1.054)
+Human+CBM 0.658 (0.591, 0.725) 0.835 (0.791, 0.879) 0.499 (0.414, 0.584) 0.758 (0.558, 0.959)
+LLM+CBM (No keyphrases) 0.555 (0.488, 0.623) 0.759 (0.713, 0.805) 0.651 (0.548, 0.754) 0.626 (0.495, 0.757)
+ResNet 0.664 (0.613, 0.716) 0.853 (0.821, 0.885) 0.457 (0.398, 0.516) 0.914 (0.748, 1.079)
+BC-LLM achieved higher accuracy and better calibration on average, compared to all the other fully interpretable CBMs (Table 1).
+Notably, soft CBMs no longer performed as well once the concepts were constrained to be fully interpretable.
+BC-LLM also outperformed CBMs trained on human collated and annotated features, potentially because it was able to consider additional concepts.
+More critically, BC-LLM outperformed ResNet50 in both accuracy and calibration, as ResNet50 easily overtrained on the small bird datasets.
+In contrast, BC-LLM could identify relevant concepts for a simple parametric model, leading to faster model convergence.
+We also assessed BC-LLM’s robustness in OOD settings, as the Bayesian framing should improve uncertainty quantification (Table 1).
+For each model, we used it to classify birds for a species it was not trained on.
+Compared to other CBM methods, BC-LLM was more appropriately unsure, yielding higher entropy across the output classes.
+We note that ResNet model had the highest entropy, though it has the advantage of not being limited to extracting a small set of interpretable concepts.
+To illustrate the benefit of using fully interpretable CBMs in OOD settings, Figure 2 (right) shows the difference when applying the Bunting bird CBM to an actual bunting bird versus an OOD sample.
+Whereas the CBM extracted zeros and ones for the actual bunting bird, it was unsure how to extract the concept “Does this image show vibrant feathers?” for the image of a dog wearing rainbow wings.
+The LLM thus outputted a probability of 0.5, citing that “the image contains vibrant colors but the wings do not actually contain feathers.” This shows how having an LLM output its reasoning when answering concrete concept questions can reveal situations that need human intervention.
+Finally, we visualize the posterior distribution over concepts using hierarchical clustering (Fig 2 left).
+Continuing with the Bunting bird example, we see that BC-LLM converges towards distinguishing concepts as training data increases and away from vague or non-specific concepts (e.g.
+“sharp beaks” apply to all bunting birds).
+Moreover, the number of unique concepts decreases with the amount of training data, because it has converged to concepts that truly distinguish the different species.
+Figure 2: Example of BC-LLM classifying Bunting birds.
+(Left) Left and right dendrograms are learned concepts when given 1/3 versus 3/3 of the training data, respectively.
+Labels are shortened concept questions, generally of the form “Does the image depict…?” Proportion of posterior samples with the concept are shown in parentheses.
+Highlighted labels are distinguishing bird features.
+(Right) Application of BC-LLM to an actual bunting bird (top) versus a dog pretending to be one (bottom).
+4.2 CBMs for classifying clinical notes with simulated outcomes
+Figure 3: MIMIC results: (Left) Comparison of BC-LLM and existing methods in terms of performance and recovery of true concepts with 95% CI.
+(Right) Dendrograms of concepts learned by BC-LLM with 100 and 800 observations (left and right, respectively).
+Labels are shortened concept questions generally of the form “Does the note mention the patient…?” Highlighted labels correspond to the true color-coded concepts in Section 4.2.
+To objectively measure how well BC-LLM can recover true concepts, we conducted a simulation study where the true concepts are known.
+To assess BC-LLM in settings where the LLM has no prior knowledge about the target, we simulated $Y$ using a LR model with five Social Determinants of Health as inputs, which were annotated from real-world patient notes in MIMIC-IV [53, 54]:
 1. Does the note mention the patient consuming alcohol in the present or the past?
@@
 5. Does the note imply the patient is retired?
-CBMs with $K=6$ were trained on 100, 200, 400, and 800 observations and evaluated on 500 held-out observations.
-This was done for 4 replicates, where the data were randomly selected from the full dataset of 7000 observations.
-An Oracle CBM was also fit as the concepts are known.
-BC-LLM outperforms the comparator methods both in predictive performance and uncertainty quantification (Figure 2a) and, with only 400 observations, performs as well as Oracle.
+CBMs with KK=6 were trained on 100 to 800 observations. Human(Oracle)+CBM uses true concepts.
+BC-LLM outperforms the comparator methods both in predictive performance and calibration (Figure 3a) and, with only 400 observations, performs as well as Human(Oracle)+CBM.
 The performance gap between BC-LLM and the comparator methods widens with more training observations because BC-LLM can iteratively refine its concepts.
-Although Boosting is also an iterative procedure, it struggles to hypothesize relevant concepts because it reviews only a few misclassified observations each iteration and cannot revise concepts added in previous iterations.
+Although Boosting LLM+CBM is also an iterative procedure, it struggles to hypothesize relevant concepts because it reviews only a few misclassified observations each iteration and cannot revise concepts added in previous iterations.
 Additionally, BC-LLM was better at recovering the true concepts compared to existing methods.
-Table 1: AUCs of CBMs and black-box models (ResNet) for classifying bird subcategories.
-Bird # classes BC-LLM One-pass Boosting ResNet
-Albatross 3 0.837 (0.771, 0.894) 0.833 (0.771, 0.890) 0.775 (0.709, 0.842) 0.837 (0.767, 0.897)
-Auklet 4 0.884 (0.830, 0.928) 0.865 (0.813, 0.914) 0.733 (0.671, 0.792) 0.831 (0.767, 0.893)
-Blackbird 4 0.869 (0.822, 0.912) 0.719 (0.671, 0.771) 0.749 (0.689, 0.801) 0.857 (0.811, 0.900)
-Bunting 3 0.932 (0.880, 0.972) 0.629 (0.585, 0.672) 0.681 (0.593, 0.762) 0.919 (0.869, 0.962)
-Flycatcher 7 0.862 (0.835, 0.887) 0.800 (0.766, 0.833) 0.784 (0.753, 0.813) 0.817 (0.781, 0.849)
-Sparrow 21 0.803 (0.785, 0.819) 0.723 (0.704, 0.739) 0.618 (0.597, 0.640) 0.812 (0.794, 0.830)
-Tern 7 0.817 (0.784, 0.849) 0.705 (0.667, 0.740) 0.609 (0.567, 0.651) 0.714 (0.674, 0.751)
-The posterior distribution over concepts can be visualized using hierarchical clustering (Figure 2b).
-Comparing concepts learned given 100 versus 800 observations, we find that the posterior learned by BC-LLM notably contracts towards the true concepts with increasing data.
-4.3 Learning to classify birds
-To demonstrate the multimodality of BC-LLM, we evaluate how well BC-LLM and other CBM methods can differentiate between bird categories in the CUB-birds image dataset (Welinder et al.,, 2010).
-[^1] We classify bird categories rather than all subtypes because there is likely a parsimonious CBM for classifying between categories but not between all subtypes.
-While the exact set of $K$ concepts needed to classify bird categories are unknown, we know certain features that are certainly spurious (e.g.
-features in the background are irrelevant).
-Similar to the prior experiment, we do not tell the LLM that we are predicting bird types.
-By replacing the bird labels with one-hot encodings, the supervised learning task is more difficult, as the LLM cannot rely on prior knowledge about birds and must conduct an unbiased exploration over concepts.
-The data is split 50/50 between training and testing (so 30 bird images per category are available for training).
-Due to cost limitations, seven bird classification tasks were considered.
-The number of concepts $K$ learned for each task was set to the number of classes, but no smaller than 4 and no greater than 10.
-For the black-box comparator, we fine-tuned the last layer of ResNet50 pre-trained on ImageNetV2 (He et al.,, 2016; Recht et al.,, 2019).
-BC-LLM outperforms all the other CBM learning procedures for every bird type (Table 1).
-Moreover, it outperforms ResNet50 model in all but one bird.
-Whereas ResNet50 easily overtrained on the small bird datasets, BC-LLM identified relevant concepts as inputs to a simple parametric model, leading to faster model convergence.
-As an example, we show concepts learned by BC-LLM for classifying Bunting birds in Figure 3b.
-The concepts learned by BC-LLM tended to be more relevant than those from comparator methods.
-For instance, given all the training data, One-pass summarization checks for a “bird sitting sideways”, “bird on a perch”, and “bird feeding or foraging,” whereas BC-LLM assigned bird position a low posterior weight.
-Moreover, visualizing the learned concepts helps us understand the learning procedure itself.
-For instance, we see that BC-LLM learns to downweight spurious features, i.e.
-concepts relating to backgrounds or bird positions, with accumulating data.
-We also assess how robust BC-LLM is in out-of-distribution (OOD) settings, as the Bayesian framing should improve uncertainty quantification.
-We apply the Bunting CBMs to classify images of the first author’s dog (Fig 3c).
-Compared to other CBM methods, BC-LLM was more appropriately unsure which Bunting bird to classify the dog as.
-Interestingly, when the dog is wearing a rainbow unicorn costume (C2), One-pass and Boosting overconfidently classify the dog as the colorful Bunting category (A3) with probabilities greater than 95%.
-Likewise, when these CBMs were used to classify images of finches (also small colorful birds), the average entropy of BC-LLM was again highest.
-Method C1 C2 Finches (e.g., C3)
-BC-LLM 0.91 0.90 0.70 (0.53 -0.88)
-One-pass 0.06 0.20 0.27 (0.00-0.66)
-Boosting 0.00 0.00 0.15 (0.00-0.57)
-Figure 3: Bunting bird classification across categories A1-A3.
-(b) Highlighted labels in the dendrogram correspond to relevant features.
-Labels are shortened concept questions, generally in the format “Does the image depict…?” Numbers in parentheses describe the proportion of posterior samples with the concept.
-Left and right dendrograms correspond to learned posteriors for 1/3 and 3/3 of the training data, respectively.
-(c) To test robustness in OOD settings, the Bunting bird CBMs are asked to classify two images of the first author’s dog and 30 images of finches.
-Entropies of the model predictions (a measure of model uncertainty) are shown for all the CBM methods.
-95% CI is shown for the Finch dataset.
-Method AUC (95% CI) Log Lik (95% CI)
-BC-LLM 0.64 (0.58, 0.70) -0.45 (-0.50, -0.40)
-One-pass 0.59 (0.52, 0.65) -0.46 (-0.51, -0.41)
-Boosting 0.59 (0.52, 0.66) -0.46 (-0.51, -0.41)
-Bag-of-words 0.52 (0.46, 0.58) -0.47 (-0.52, -0.41)
+Visualizing the posterior distribution using hierarchical clustering (Fig 3 right), the learned posterior notably contracts towards the true concepts with increasing data as the number of training observations increases from 100 to 800.
+Moreover, BC-LLM remains appropriately unsure about certain concepts that are highly correlated.
+For instance, it states many variations of the smoking concept that are essentially impossible to distinguish between, which is both expected and desired.
+4.3 Augmenting a tabular model with clinical notes
+Here we demonstrate how BC-LLM can help ML developers integrate different data modalities to improve model performance.
+The data science team at hospital [Anonymous] has both tabular data and unstructured notes from the electronic health record (EHR).
+The team has so far fit a model using the tabular data to predict readmission risk for heart failure patients and wants to assess if clinical notes contain additional concepts that are predictive.
+To address this, we extend BC-LLM to take as inputs (i) the risk prediction from the original tabular model and (ii) $K=4$ concepts extracted from the clinical notes, which can be viewed as revising the existing model by adding features.
+The CBMs were trained on 1000 patients and evaluated on 500 held-out patients.
+The original tabular model, trained on lab and flowsheet values, achieved an AUC of 0.60.
+BC-LLM revised this model to perform substantially better, achieving an AUC of 0.64 (Fig 4 left).
+Moreover, it outperforms comparator methods with respect to both AUC and brier score.
+To assess the interpretability of the learned concepts, four clinicians were asked to rate concepts in the CBMs as well as top features in the original tabular model by how predictive they were, from 1==low to 3==high clinical relevance.
+[^4] Tabular features were rewritten as questions so they could not be distinguished from learned concepts.
+See Appendix for annotation instructions.
+Concepts from BC-LLM received an average rating of 2.5, whereas concepts from the other methods were rated $\leq$2.1 (Fig 4 right).
+When unblinded to the concepts, clinicians commented on a number of advantages to using BC-LLM.
+First, many of the learned concepts were known to be causally relevant but were difficult to engineer solely from tabular data.
+For example, information for determining “substance dependence” is rarely documented in the tabular data but often noted in clinical notes.
+Second, BC-LLM also suggested new features to engineer that the data science team had not previously considered.
+Most interestingly, some of the learned concepts pointed to interventions that the hospital could try to implement, such as ensuring follow-up appointments were scheduled and opting for medication regimes that were easier to follow.
+In contrast, the top features in the original tabular model were difficult to interpret and not actionable.
+Method AUC (95% CI) Brier (95% CI)
+BC-LLM 0.64 (0.58, 0.70) 0.14 (0.12, 0.62)
+LLM+CBM 0.59 (0.52, 0.65) 0.29 (0.25, 0.33)
+Boosting LLM+CBM 0.59 (0.52, 0.66) 0.14 (0.12, 0.16)
+Bag-of-words 0.52 (0.46, 0.58) 0.29 (0.25, 0.34)
+Refer to caption Refer to caption
 Figure 4: Learning to augment a readmission risk prediction model for heart failure patients.
-Labels are shortened versions of concept questions of the format “Does the note mention the patient having…?”
-4.4 Augmenting an existing tabular model with real-world clinical notes
-Finally, we demonstrate how BC-LLM can help ML developers integrate different data modalities to improve model performance.
-In particular, the data science team at the Zuckerberg San Francisco General Hospital (ZSFG) has both tabular data and unstructured clinical notes from the electronic health record (EHR).
-The team has so far fit a model using the tabular data to predict 30-day unplanned readmission risk for patients with congestive heart failure (CHF).
-Using CBMs, the team wants to assess if clinical notes contain valuable information to extract as additional features.
-Moreover, the learned model should be interpretable and easy to audit.
-To address this, we extend BC-LLM by having the CBM take as inputs (i) the risk prediction from the original tabular model and (ii) $K$ concepts extracted from the clinical notes.
-This can be viewed as revising the existing model by including additional features.
-We set to $K=4$.
-The CBMs were trained on 1000 patients and evaluated on 500 held-out patients.
-Data in this experiment contained PHI, for which IRB approval was obtained.
-The original tabular model, trained on lab and flowsheet values, achieved an AUC of 0.60.
-BC-LLM revised this model to perform substantially better, achieving an AUC of 0.64 (Figure 4).
-Moreover, it outperforms comparator methods with respect to both AUC and log likelihood.
-When the CBMs were shown to four clinicians, [^2] All clinicians work or have worked in ZSFG.
-Three are in the hospital’s data science team, one of whom (LZ) is a coauthor on this paper.
-The fourth leads a separate clinical data science research group.
-They were all interviewed separately.
-they found that the CBM learned by BC-LLM to be substantially more interpretable and meaningful than the existing tabular model.
-The clinicians from the hospital’s data science team noted that they had not considered engineering many of the learned concepts on the list.
-All clinicians noted that many of the learned concepts are known to be causally relevant and would have been difficult to learn from existing tabular data.
-For example, “substance dependence” is rarely documented directly in the tabular data, and information relevant to determining its value is often missing.
-Finally, the clinicians highly valued the fact that many of the learned concepts were clinically actionable, such as helping patients better control their diabetes or opting for medication regimes that are easier for patients to follow.
+Dendrogram labels are shortened questions of the format “Does the note mention the patient having…?”.
+Highlighted concepts received scores from clinicians as being highly predictive (scores 2.5+).
+Average clinician ratings for concepts/features from the different methods are shown on the right.
 5 Discussion
-BC-LLM is a new method for learning CBMs that iteratively proposes concepts using an LLM within a Bayesian framework, which allows for rigorous uncertainty quantification despite LLMs being prone to error and hallucinations.
-The method is highly general: it is compatible with various data modalities (text, images, and tabular data) and can be extended beyond the settings of binary and multiclass classification.
-The empirical results show that BC-LLM outperforms existing methods, learns to recover relevant concepts and downweight spurious ones, and can suggest useful features for ML developers to engineer.
-Code for running BC-LLM and reproducing results in the paper are available at https://github.com/jjfeng/bc-llm.
+BC-LLM is a new approach to learning CBMs that iteratively proposes concepts using an LLM within a Bayesian framework, which allows for rigorous uncertainty quantification despite LLMs being prone to error and hallucinations.
+Because it explores and suggests concepts in a data-adaptive manner, BC-LLM is particularly well-suited for settings with limited prior knowledge about which concepts are relevant or where the number of potentially relevant concepts is infinite.
+The method is compatible with various data modalities (text, images, and tabular data) and can be extended beyond the settings of binary and multiclass classification.
+The empirical results show that BC-LLM outperforms existing methods, even black-box models in certain settings.
+Future work.
+BC-LLM is currently designed for learning highly interpretable CBMs where the number of concepts $K$ is typically no more than $20$.
+While BC-LLM can be applied to learn even more concepts, future directions can consider further speeding up posterior inference, such as through mini-batching of observations or concepts.
+Impact Statement.
+The interpretability provided by BC-LLM additionally enables better understanding of the underlying algorithm and facilitates human oversight, which may improve the safety of AI algorithms and AI-based decision-making.
+The use of LLMs incurs significant computational cost and corresponding environmental impacts; however, BC-LLM is much more computationally efficient than comparable LLM-based CBMs, potentially reducing overall environmental impacts.
 Acknowledgements
-JF was funded through a Patient-Centered Outcomes Research Institute® (PCORI®) Award (ME-2022C1-25619).
-YT gratefully acknowledges support from NUS Start-up Grant A-8000448-00-00 and MOE AcRF Tier 1 Grant A-8002498-00-00.
+This work was funded through a Patient-Centered Outcomes Research Institute® (PCORI®) Award (ME-2022C1-25619).
+The views presented in this work are solely the responsibility of the author(s) and do not necessarily represent the views of the PCORI®, its Board of Governors or Methodology Committee.
 The authors thank the UCSF AI Tiger Team, Academic Research Services, Research Information Technology, and the Chancellor’s Task Force for Generative AI for their technical support related to the use of Versa API gateway.
-The authors thank Yifan Mai, Seth Goldman, Andrew Bishara, and Jim Marks for their invaluable feedback on the work.
-The views presented in this work are solely the responsibility of the author(s) and do not necessarily represent the views of the PCORI®, its Board of Governors or Methodology Committee.
+The authors thank Patrick Vossler, Yifan Mai, Seth Goldman, Andrew Bishara, Jim Marks, Julian Hong, Aaron Kornblith, Nicholas Petrick, and Gene Pennello for their invaluable feedback on the work.
 Appendix
@@
 In this section, we provide intuition for the Multiple-Try Metropolis-Hastings partial posterior method (Algorithm 3).
-To draw connections with the original Multiple-Try Metropolis-Hastings method (Liu et al.,, 2000), we first work in a completely abstract setting.
+To draw connections with the original Multiple-Try Metropolis-Hastings method [45], we first work in a completely abstract setting.
 Given a state space $\Omega$, stationary distribution $\pi$, and proposal transition kernel $T(y;x)$, define weights $w(x,y)=\pi(x)T(y;x)$.
@@
 Classical version.
-We briefly describe Liu et al., (2000)’s original method. Suppose the current state is $x$.
+We briefly describe [45]’s original method. Suppose the current state is $x$.
   1. 1.
@@
     $\beta(x,y;\mathbf{z},\mathbf{z}^{*})=\min\left\{\frac{\sum_{i=1}^{M}w(z_{i},x)}{\sum_{i=1}^{M}w(z_{i}^{*},y)},1\right\}.$ (3)
-One can show that this Markov chain satisfies the detailed balance equations when marginalizing out the intermediate states $z_{1},z_{2},\ldots,z_{M},z_{1}^{*},\ldots,z_{M}$ (see Theorem 1 in Liu et al., (2000)).
+One can show that this Markov chain satisfies the detailed balance equations when marginalizing out the intermediate states $z_{1},z_{2},\ldots,z_{M},z_{1}^{*},\ldots,z_{M}$ (see Theorem 1 in [45]).
 We would like, however, to minimize the number of proposals necessary, so it is natural to ask whether the algorithm can be modified such that the proposals needed for the backward transitions, $z_{1}^{*},\ldots,z_{M-1}^{*}$, can be omitted.
@@
 To see this, we compute
-[…]
+$\begin{split}\pi(x)\Phi(x,y)&=k\pi(x)T(y;x)\int\prod_{i=2}^{M}T(z_{i};x)q(y;\mathbf{z})\alpha(x,y;\mathbf{z}_{-1})d\mathbf{z}_{-1}\\ &=k\int\min\left\{\pi(y)T(x;y)\prod_{i=2}^{M}T(z_{i};y)q(x;\mathbf{z}^{*}),\pi(x)T(y;x)\prod_{i=2}^{M}T(z_{i};x)q(y;\mathbf{z})\right\}d\mathbf{z}_{-1}\\ &=\pi(y)\Phi(y,x).\qed\end{split}$ (7)
 Connections between modified and classical versions.
@@
 $\begin{split}\min\left\{\frac{\pi(y)\prod_{i=1}^{M}T(z_{i};y)q(x;\mathbf{z}^{*})}{\pi(x)\prod_{i=1}^{M}T(z_{i}^{*};x)q(y;\mathbf{z})},1\right\}&=\min\left\{\frac{\pi(y)\prod_{i=1}^{M}T(z_{i};y)\frac{w(x,y)}{\sum_{i=1}^{M}w(z_{i}^{*},y)}}{\pi(x)\prod_{i=1}^{M}T(z_{i}^{*};x)\frac{w(y,x)}{\sum_{i=1}^{M}w(z_{i},x)}},1\right\}\\ &=\min\left\{\frac{\sum_{i\neq\check{m}}w(z_{i},x)+w(y,x)}{\sum_{i\neq\check{m}}w(z_{i},y)+w(x,y)}\cdot\prod_{i=2}^{M}\frac{T(z_{i};y)}{T(z_{i};x)},1\right\}.\end{split}$ (8)
-If $T(-,-)$ is invariant in the first argument (as in our application to BC-LLM), then this last formula is exactly equal to $\beta(x,y;\mathbf{z},\mathbf{z}^{*})$, the acceptance probability for the classical version, but if we were to use the same points for both the non-realized forward and backward proposals.
+If T(,)T(-,-) is invariant in the first argument (as in our application to BC-LLM), then this last formula is exactly equal to $\beta(x,y;\mathbf{z},\mathbf{z}^{*})$, the acceptance probability for the classical version, but if we were to use the same points for both the non-realized forward and backward proposals.
 Equivalence with Multi-SS-MH-Update.
@@
   • Warm-start and Burn-in: Since Gibbs sampling can be slow to converge, we precede it with a warm-start, which we obtain by updating concepts greedily.
-That is, we select the concept that maximizes $\operatorname{argmax}p(\gamma|\vec{c}_{-k},\mathbf{y}_{S},\boldsymbol{\Phi})$, instead of sampling from the distribution.
+That is, we select the concept that maximizes $\operatorname{argmax}p(\gamma|\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})$, instead of sampling from the distribution.
 In experiments, we run warm-start for one epoch and stored the last 20 iterates as posterior samples; the rest of the samples were treated as burn-in.
@@
 10) for more complete exploration of concepts.
-Appendix C Example prompt for keyphrase extraction (Step 0)
+Appendix C Example prompts
+Here we provide example prompts that one may use with BC-LLM.
+The prompt should vary with how much prior knowledge one has about the prediction target as well as how much information one would like to reveal to the LLM about the prediction target.
+C.1 Example prompt for keyphrase extraction (Step 0)
+Example prompt for extracting keyphrases from images, where we reveal limited information to the LLM about the prediction target:
+Given an image, your task is to brainstorm a set of descriptors that will help classify the image to its corresponding label Y.
+For the provided image, list as many descriptors about it.
+For each descriptor, also list as many descriptors that mean the same thing or generalizations of the descriptor.
+All descriptors, synonyms, and generalizations cannot be more than two words.
+Output characteristics that this image possess, such as "round wings" or "red head." Do not output general categories of descriptors like "wing color" or "head shape." Output at least 10 keyphrases.
+Example prompt for extracting keyphrases from clinical notes, where we do not reveal information about the prediction target but we provide ideas on patient characteristics that one should extract:
 Here is a clinical note: $\langle$ note $\rangle$
@@
 All descriptors, synonyms, and generalizations cannot be more than two words.
-Output as a JSON in the following format.
-Do not output more than 30 entries in the JSON:
-Appendix D Example prompt for proposing concepts (Step 2)
-The goal is to come up with a concept bottleneck model (CBM) that only extracts 5 meta-concepts from patient notes to predict some outcome Y with maximum accuracy.
+Output as a JSON in the following format...
+C.2 Example prompt for proposing concepts (Step 2)
+Example prompt for having the LLM brainstorm candidate concepts based on top keyphrases in the keyphrase model, where we do not reveal information about the prediction target and only reveal that the dataset contains patient notes:
+The goal is to come up with a concept bottleneck model (CBM) that only extracts 3 meta-concepts from patient notes to predict some outcome Y with maximum accuracy.
 A meta-concept is a binary feature extractor defined by a yes/no question.
-We have 4 meta-concepts so far:
+We have 2 meta-concepts so far:
   1. 1.
-    Does the note mention the patient having social support?
   2. 2.
-    Does the note mention the patient having stable blood pressure?
-  3. 3.
-    Does the note mention the patient not experiencing respiratory distress?
-  4. 4.
-    Does the note mention the patient experiencing substance abuse or dependence?
-To come up with the 5th meta-concept, I have done the following: I first fit a CBM on the 4 existing meta-concepts.
-Then to figure out how to improve this 4-concept CBM, I first asked an LLM to extract a list of concepts that are present in each note, and then fit a linear regression model on the extracted concepts to predict the residuals of the 4-concept CBM.
+To come up with the 3rd meta-concept, I have done the following: I first fit a CBM on the 2 existing meta-concepts.
+Then to figure out how to improve this 3-concept CBM, I first asked an LLM to extract a list of concepts that are present in each note, and then fit a linear regression model on the extracted concepts to predict the residuals of the 3-concept CBM.
 These are the top extracted concepts in the resulting residual model, in descending order of importance:
-$\langle$ top words from residual model $\rangle$
-To interpret this residual model, a general rule of thumb is that an extracted concept with a large positive coefficient means that the outcome is positively associated with the concept being mentioned in the note, and an extracted concept with a negative coefficient means that the outcome is negatively associated with the concept being mentioned in the note.
-Remember that notes typically discuss the current state as well as the past medical history of a patient, so a concept being mentioned is in reference to the present OR past state of the patient.
-Given the residual model, create cohesive candidates for the 5th meta-concept.
+$\langle$ top keyphrases from keyphrase model $\rangle$
+Given the residual model, create cohesive candidates for the 3rd meta-concept.
 Be systematic and consider all the listed concepts in the residual model.
@@
 Suggestions for generating candidate meta-concepts: Do not propose meta-concepts that are simply a union of two different concepts (e.g.
-“Does the note mention this patient experiencing stomach pain or being female?” is not allowed), questions with answers that are almost always a yes (e.g.
-the answer to “Does the note mention this patient being sick?” is almost always yes), or questions where the yes/no options are not clearly defined (e.g.
-“Does the note mention this patient experiencing difficulty?” is not clearly defined because difficulty may mean financial difficulty, physical difficulties, etc).
+‘‘Does the note mention this patient experiencing stomach pain or being female?" is not allowed), questions with answers that are almost always a yes (e.g.
+the answer to ‘‘Does the note mention this patient being sick?" is almost always yes), or questions where the yes/no options are not clearly defined (e.g.
+‘‘Does the note mention this patient experiencing difficulty?" is not clearly defined because difficulty may mean financial difficulty, physical difficulties, etc).
 Do not propose meta-concepts where you would expect over 95% agreement or disagreement with the 4 existing meta-concepts (e.g.
-“Does the note mention the patient having high blood pressure?” overlaps too much with “Does the note mention the patient having hypertension?”).
-Finally, summarize all the generated candidates for the 5th meta-concept in a JSON.
-Merge any candidate meta-concepts that are essentially the same (where you would expect over 95% agreement) or essentially opposites (you would expect over 95% disagreement).
-In the JSON, include a list of comma-separated list of phrases that mean the same or opposite of each candidate meta-concept.
-Appendix E Example prompt for annotations for candidate concepts (Step 3)
-A simple concept extraction procedure is to obtain binary or categorical concept extractions.
-In line with findings in Kim et al., (2023), we found that BC-LLM performed better when the LLM quantified its uncertainty when extracting concepts.
-So if the LLM was unsure about the concept’s value, it was allowed to return a probability.
-As such, we use prompts like the following:
+‘‘Does the note mention the patient having high blood pressure?" overlaps too much with ‘‘Does the note mention the patient having hypertension?").
+Finally, summarize all the generated candidate concepts in a JSON.
+C.3 Example prompt for annotations for candidate concepts (Step 3)
+A simple concept extraction procedure is to obtain binary or categorical concept extractions. We use prompts like this:
 You will be given a clinical note.
@@
 clinical note: $\langle$ note $\rangle$
-Appendix F Experiment details
-Note on computation time: When running BC-LLM with GPT-4o-mini-2024-07-18 and GPT-4o, each iteration completed within a few minutes.
-F.1 MIMIC
+Note that in cases where the LLM was unsure about the concept’s value, it was allowed to return a probability, which prior works have found to be helpful [6].
+In general, the LLM outputted 1’s and 0’s.
+Probabilities were outputted occasionally, which as highlighted in the main manuscript, can be audited and checked by a human.
+Appendix D Additional Results
+D.1 MIMIC
+We evaluate BC-LLM against baseline methods using Cohere’s Command-R LLM using the same settings as discussed in Section 4.2. Performance can be found in Figure 5.
+Figure 5: Comparison of BC-LLM and existing methods in terms of performance and recovery of true concepts. Error bars indicate the 95% CI.
+D.2 CUB-Birds dataset
+We evaluate BC-LLM against baseline methods using Anthropic’s Claude 3.5 Haiku using the same settings as discussed in Section 4.1. Performance can be found in Table 2.
+Method Accuracy ($\uparrow$) AUC ($\uparrow$) Brier ($\downarrow$)
+BC-LLM 0.665 (0.593, 0.736) 0.849 (0.811, 0.888) 0.443 (0.367, 0.519)
+LLM+CBM 0.583 (0.517, 0.648) 0.771 (0.722, 0.821) 0.507 (0.434, 0.580)
+Boosting LLM+CBM 0.640 (0.568, 0.712) 0.811 (0.769, 0.852) 0.552 (0.455, 0.649)
+Table 2: Results for the CUB-Birds dataset using Claude 3.5 Haiku
+Appendix E Experiment details
+E.1 Evaluating correct concept matches
+Quantifying a “correct” concept match requires going beyond exact string matching, as there are many possible refinements for a single concept.
+For instance, for the concept “Does the note mention the patient smoke at present?”, there are rephrasings (“Does this note mention if the patient currently smokes?”), polar opposites (“Does the note not mention the currently smoking?”), or closely overlapping concepts (“Does the note mention the patient’s smoking history?”).
+To quantify if a learned concept was correct, we used the following rule: if the absolute value of the Pearson correlation for LLM annotations of two concepts exceeds 50%, the two concepts are sufficiently similar.
+We manually confirmed that any pair of learned and true concepts that passed this rule were always closely related.
+E.2 MIMIC
 The entire dataset consists of 7043 observations of which we trained on 100, 200, 400, and 800 randomly selected observations and evaluated 500 held-out observations.
@@
 The label $Y$ was generated per a LR model, in which
+$\log{\frac{\Pr(Y=1|X)}{\Pr(Y=0|X)}}=$ $4*\mathbbm{1}\{\text{Does the note imply the patient is unemployed?}\}$
+$+4*\mathbbm{1}\{\text{Does the note imply the patient is retired?}\}$
 […]
@@
 […]
-[…]
-[…]
-One-pass summarization and BC-LLM were run to fit CBMs with $K=6$ concepts. BC-LLM was run for $T=5$ epochs. Boosting ran for 10 iterations, in which at most 1 new candidate was added each iteration.
-F.2 CUB-Birds dataset
-To train a bird classifier for $R$ subtypes, One-pass summarization and BC-LLM were run to fit CBMs with $K=\min(10,\max(4,R))$ concepts.
+LLM+CBM summarization and BC-LLM were run to fit CBMs with $K=6$ concepts.
+BC-LLM was run for $T=5$ epochs.
+Boosting LLM+CBM ran for 10 iterations, in which at most 1 new candidate was added each iteration.
+E.3 CUB-Birds dataset
+To train a bird classifier for $R$ subtypes, LLM+CBM summarization and BC-LLM were run to fit CBMs with $K=\min(10,\max(4,R))$ concepts.
 BC-LLM was run for a maximum of $T=5$ epochs or 25 iterations, whichever was reached first.
-Boosting ran for 10 iterations, in which at most 1 new candidate was added each iteration.
-F.3 Clinical notes and tabular data from ZSFG
+Boosting LLM+CBM ran for 10 iterations, in which at most 1 new candidate was added each iteration.
+For the black-box comparator, we used a ResNet50 model pre-trained on ImageNetV2, which ensures ResNet’s training set does not overlap with the CUB-Birds dataset.
+E.4 Clinical notes and tabular data from Hospital Anonymous
+Data in this experiment contained PHI, for which IRB approval was obtained.
 To predict readmission risk, the models were trained to analyze the sections “Brief history leading to hospitalization” and the “Summary of hospitalization” in the discharge summary for each patient.
-One-pass summarization and BC-LLM were run to fit CBMs with $K=4$ concepts.
+LLM+CBM summarization and BC-LLM were run to fit CBMs with $K=4$ concepts.
 BC-LLM was run for $T=5$ epochs.
-Boosting ran for 10 iterations, in which at most 1 new candidate was added each iteration.
-For the black-box comparator, we used a ResNet50 model pre-trained on ImageNetV2, which ensures ResNet’s training set does not overlap with the CUB-Birds dataset.
-For the OOD experiment, the Bunting CBM was applied to 30 images of finches.
-Appendix G Laplace Approximation of Split-Sample Posterior
+Boosting LLM+CBM ran for 10 iterations, in which at most 1 new candidate was added each iteration.
+E.5 Clinician survey
+To conduct the survey for clinical relevance of features and concepts, the following instructions were given to clinicians:
+Please rate the features learned by BC-LLM and its comparator methods.
+In the attached CSV, we have listed 25 or so features from various methods, where we mask which method has learned which feature.
+Please fill in the column "How clinically relevant is the candidate concept with determining a patient’s readmission risk?
+Enter values = 1: low, 2:medium, 3: high." Please try to use the full range of scores, rather than assigning all the features the same score.
+Table 3 shows the table that clinicians were asked to fill in. The ordering of features was randomly shuffled to obfuscate which method had learned which concepts.
+Table 3: Clinicians were asked to fill in the following table with their ratings of the clinical relevance of each feature/concept for predicting a patient’s readmission risk.
+Candidate concept Score (1-3)
+Does the patient have multiple medications prescribed?
+Does the patient have a history of hypertension (HTN)?
+Is this patient experiencing chest pain?
+What is the patient’s weight?
+Does the patient have heart failure with reduced ejection fraction?
+What are the patient’s lab results for BNP?
+Does the patient have substance dependence?
+Is the patient stable at discharge?
+Does the patient have a history of falls?
+How many emergency department encounters does this patient have?
+What are the patient’s lab results for Lactate Dehydrogenase?
+Does the patient have uncontrolled diabetes?
+Does the patient have a history of frequent hospital admissions?
+How many encounters does this patient have?
+Does the patient have a history of diabetes mellitus type 2 (DM2)?
+Does the patient have complex medical history?
+Does the patient have a history of sepsis?
+What are the patient’s lab results for Creatinine?
+Does the patient have a history of drug or substance use disorder?
+What is the patient’s value for Expiratory Positive Airway Pressure (EPAP)?
+Does the patient have had any missed outpatient appointments?
+What are the patient’s lab results for Glucose?
+Are the patient’s outpatient follow-up appointments scheduled?
+Appendix F Laplace Approximation of Split-Sample Posterior
 In this section, we describe how to perform Laplace approximation of the split-sample posterior $p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})$ when the likelihood model is logistic and the prior on the coefficient vector $\vec{\theta}$ is a standard normal $\mathcal{N}(0,\gamma^{2}I)$.
@@
 $\begin{split}&p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})\\ &=\frac{p(\mathbf{y}|\vec{c},\mathbf{X})}{p(\mathbf{y}_{S}|\vec{c},\mathbf{X})}\\ &\approx\exp\left(\sum_{i=1}^{n}\left(y_{i}\vec{\theta}_{\operatorname{MAP}}^{T}\Phi_{i\cdot}-\log(1+\exp(\vec{\theta}_{\operatorname{MAP}}^{T}\Phi_{i\cdot}))\right)-\sum_{i\in S}\left(y_{i}\vec{\theta}_{\operatorname{MAP},S}^{T}\Phi_{i\cdot}-\log(1+\exp(\vec{\theta}_{\operatorname{MAP},S}^{T}\Phi_{i\cdot}))\right)\right)\\ &\quad\cdot\exp\left(\frac{\|\vec{\theta}_{\operatorname{MAP},S}\|_{2}^{2}-\|\vec{\theta}_{\operatorname{MAP}}\|_{2}^{2}}{2\gamma^{2}}\right)\left(\frac{\det(H_{S}(\vec{\theta}_{\operatorname{MAP},S}))}{\det(H(\vec{\theta}_{\operatorname{MAP}}))}\right)^{1/2}.\end{split}$ (17)
-Remark G.1.
-The MAP estimate $\vec{\theta}_{\operatorname{MAP}}$ can be computed by a call to scikit-learn’s LogisticRegression() model, setting C $=2\gamma^{2}$.
-Remark G.2.
+Remark F.1.
+The MAP estimate $\vec{\theta}_{\operatorname{MAP}}$ can be computed by a call to scikit-learn’s LogisticRegression() model, setting C $2\gamma^{2}$.
+Remark F.2.
 In the current implementation of our procedure, we use a slightly different method to perform Laplace approximation, which instead approximates the integral in (1).
-Appendix H Proof of Theorem 3.1
+Appendix G Proof of Theorem 3.1
 Recall the definition
@@
 Note that the definition does not require that the model class is correctly specified (i.e.
-there does not have to be some $\vec{c}$ and $\vec{\theta}$ so that (1) holds).
+there does not have to be some $\vec{c}$ and $\vec{\theta}$ so that $p\left(Y=1|X=x,\vec{\theta},\vec{c}\right)=\sigma\left(\sum_{k=1}^{K}\theta_{k}\phi_{c_{k}}(x)+\theta_{0}\right)$).
 Before presenting the proof, we first detail the assumptions we make.
-Assumption H.1.
-  1. (i)
+Assumption G.1.
+  1. 1.
     $Y\in\{0,1\}$ is a binary response;
-  2. (ii)
+  2. 2.
     The data $\mathcal{D}$ is generated i.i.d. from some distribution $\nu$;
-  3. (iii)
+  3. 3.
     The prior on $\vec{\theta}$ is $\mathcal{N}(0,\gamma^{2}I)$;
-  4. (iv)
+  4. 4.
     $\mathcal{C}^{*}\coloneqq\operatorname{argmax}_{\vec{c}}L(\vec{c})$ comprises all permutations of a single vector $\vec{c}^{~*}$;
-  5. (v)
+  5. 5.
     There exists some $\Delta>0$ such that $L(\vec{c}^{~*})-\operatorname{argmax}_{\vec{c}\notin\mathcal{C}^{*}}L(\vec{c})\geq\Delta$;
-  6. (vi)
+  6. 6.
     Let $\mathcal{C}_{n}\coloneqq\cup_{\vec{c}_{-k},\mathbf{y},\mathbf{X}}\{c\colon Q(c;\vec{c}_{-k},\mathbf{y},\mathbf{X})>0\}$.
-Assume $\mathcal{C}_{n}$ is a finite set of size at most $\exp(n^{1-\epsilon})$ for some $10<\epsilon<$ 1;
-  7. (vii)
+Assume $\mathcal{C}_{n}$ is a finite set of size at most $\exp(n^{1-\epsilon})$ for some $0<\epsilon<1$;
+  7. 7.
     There exists $\eta>0$ such that for each $\vec{c}_{-k}$, there exists $l\in\{1,2,\ldots,K\}$ such that $c_{l}^{*}=\operatorname{argmax}_{c}L((c,\vec{c}_{-k}))$, $c_{l}^{*}\notin\vec{c}_{-k}$, and $Q(c^{*}_{l};\vec{c}_{-k},\mathbf{y},\mathbf{X})\geq\eta$ almost surely for all $\mathbf{y},\mathbf{X}$ with $n$ large enough;
-  8. (viii)
+  8. 8.
     There exists some $B$ such that for any $\vec{c}\in\mathcal{C}_{n}^{K}$, all distinct, we have $\|\vec{\theta}_{\vec{c}}^{*}\|_{2}\leq R$, where
     $\vec{\theta}_{\vec{c}}^{*}\coloneqq\operatorname{argmax}_{\vec{\theta}}\mathbb{E}_{(X,Y)\sim\nu}\{\log p(Y|X,\vec{\theta},\vec{c})\};$ (19)
-  9. (ix)
+  9. 9.
     Let $\Phi_{\vec{c}}=(\phi_{c_{1}}(x),\phi_{c_{2}}(x),\ldots,\phi_{c_{K}}(x))$ for $x\sim\nu$.
@@
 Here, $\mathcal{C}_{n,K}$ denotes the $K$-fold product of $\mathcal{C}_{n}$ but with any vectors containing duplicate concepts removed;
-  10. (x)
+  10. 10.
     We modify the algorithm so that the random subset $S$ can only take one of $\exp(n^{1-\epsilon})$ possibilities $\mathcal{S}_{n}\coloneqq\{S_{1},S_{2},\ldots,S_{B,n}\}$.
@@
 The Markov chain, when restricted to $\mathcal{A}$, is aperiodic, since there must exist a self-loop.
-Furthermore, it is finite according to Assumption H.1(vi).
-As such, $(\vec{c}^{~{}t})_{t=1}^{\infty}$ has a unique stationary distribution $\pi_{\mathcal{A}}$ supported on $\mathcal{A}$ (LevinPeresWilmer2006).
+Furthermore, it is finite according to Assumption G.1(vi).
+As such, $(\vec{c}^{~t})_{t=1}^{\infty}$ has a unique stationary distribution $\pi_{\mathcal{A}}$ supported on $\mathcal{A}$ [55].
 Indeed, all stationary distributions of $(\vec{c}^{~t})_{t=1}^{\infty}$ are convex combinations of such distributions.
@@
 In order to analyze $\pi_{\mathcal{A}}$, we need quantitative control over the transition probabilities via concentration.
-For a given dataset $\mathcal{D}$ of size $n$, we condition on the event guaranteed by Proposition H.3.
+For a given dataset $\mathcal{D}$ of size $n$, we condition on the event guaranteed by Proposition G.3.
 Step 4: Convergence.
@@
 To see this, notice that any state $j$ for $b+1\leq j\leq a$ corresponds to a vector $\tilde{\vec{c}}^{~(0)}\notin\mathcal{C}^{*}$.
-However, using Assumption H.1(vii), there is a sequence $(\tilde{\vec{c}}^{~{}(0)},\tilde{\vec{c}}^{~{}(1)},\ldots,\tilde{\vec{c}}^{~{}(K-1)})$ such that for $k=1,\ldots,K$, we have (i) $\tilde{\vec{c}}^{~{}(k)}_{-k}=\tilde{\vec{c}}^{~{}(k-1)}_{-k}$, (ii) $Q(\tilde{c}_{k};\tilde{\vec{c}}^{~{}(k-1)}_{-k},\mathbf{y}_{S},\mathbf{X})\geq\eta$, (iii) $L(\tilde{\vec{c}}^{~{}(k)})\geq L(\tilde{\vec{c}}^{~{}(k-1)})$ with equality holding iff $\tilde{\vec{c}}^{~{}(k)}=\tilde{\vec{c}}^{~{}(k-1)}$.
+However, using Assumption G.1(vii), there is a sequence $(\tilde{\vec{c}}^{~(0)},\tilde{\vec{c}}^{~(1)},\ldots,\tilde{\vec{c}}^{~(K-1)})$ such that for $k=1,\ldots,K$, we have (i) $\tilde{\vec{c}}^{~(k)}_{-k}=\tilde{\vec{c}}^{~(k-1)}_{-k}$, (ii) $Q(\tilde{c}_{k};\tilde{\vec{c}}^{~(k-1)}_{-k},\mathbf{y}_{S},\mathbf{X})\geq\eta$, (iii) $L(\tilde{\vec{c}}^{~(k)})\geq L(\tilde{\vec{c}}^{~(k-1)})$ with equality holding iff $\tilde{\vec{c}}^{~(k)}=\tilde{\vec{c}}^{~(k-1)}$.
 Combining (iii) with (30), we see that all acceptance ratios along this sequence are equal to 1, which means that the transition probability from state $j$ to states in $\mathcal{C}^{*}$ is at least
 $\prod_{k=1}^{K}Q(\tilde{c}_{k};\tilde{\vec{c}}^{~(k-1)}_{-k},\mathbf{y}_{S},\mathbf{X})\geq\eta^{K}.$ (23)
-The sum of entries in row $j-b$ in $P_{22}$ is the total probability of transition from state $j$ to other states within $\mathcal{A}\backslash\mathcal{C}^{*}$ and hence has value less than $1-\eta^{K}$.
+The sum of entries in row jbj-b in $P_{22}$ is the total probability of transition from state $j$ to other states within $\mathcal{A}\backslash\mathcal{C}^{*}$ and hence has value less than $1-\eta^{K}$.
 This holds for any row.
-Applying Lemma H.2, we thus get $\|P_{22}\|_{op}\leq 1-\eta^{K}$, which allows us to bound the denominator in (24).
+Applying Lemma G.2, we thus get $\|P_{22}\|_{op}\leq 1-\eta^{K}$, which allows us to bound the denominator in (22).
 Next, every entry in $P_{12}$ is the probability of transition from a state in $\mathcal{C}^{*}$ to a state in $\mathcal{A}\backslash\mathcal{C}^{*}$.
-Using Step 3 and Assumption H.1(v), the acceptance probability $\alpha$ of such a transition satisfies
+Using Step 3 and Assumption G.1(v), the acceptance probability $\alpha$ of such a transition satisfies
 $\begin{split}\log\alpha&\leq\max_{\vec{c}\in\mathcal{A}\backslash\mathcal{C}^{*}}\log p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})-\log p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c}^{~*},\mathbf{X})\\ &\leq-n\Delta+O(n^{1-\epsilon^{\prime}}).\end{split}$ (24)
@@
 Finally, we note that all hidden constants can be chosen to be independent of $n$. ∎
-Lemma H.2.
+Lemma G.2.
 Let $M$ be a $m\times n$ matrix whose entries are non-negative, and suppose $\sum_{j=1}^{n}M_{ij}\leq\beta$ for $i=1,2,\ldots,m$. Then $\|M\|_{op}\leq\beta$.
@@
 as we wanted. ∎
-Proposition H.3 (Concentration of split-sample posterior).
-Assuming the conditions in Assumption H.1, then for any $\epsilon^{\prime}<\epsilon/2$, we have
+Proposition G.3 (Concentration of split-sample posterior).
+Assuming the conditions in Assumption G.1, then for any $\epsilon^{\prime}<\epsilon/2$, we have
 $\sup_{\vec{c}\in\mathcal{C}_{n,K},S\in\mathcal{S}_{n}}\left|\frac{1}{\lceil(1-\omega)n\rceil}\log p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})-L(\vec{c})\right|=O(n^{-\epsilon^{\prime}})$ (30)
@@
 $\left\|n^{-1/2}\sum_{i=1}^{n}\left(l_{i}(\vec{\theta})-\bar{l}(\vec{\theta})\right)-n^{-1/2}\sum_{i=1}^{n}\left(l_{i}(\vec{\theta}^{\prime})-\bar{l}(\vec{\theta}^{\prime})\right)\right\|_{\psi_{2}}\lesssim K\|\vec{\theta}-\vec{\theta}^{\prime}\|_{2}.$ (34)
-We can thus use Talagrand’s comparison inequality (vershynin2018high) to get, for any $t>0$, that
+We can thus use Talagrand’s comparison inequality [56] to get, for any $t>0$, that
 $n^{-1/2}\sup_{\vec{\theta}\in B(2R)}\left|\sum_{i=1}^{n}\left(l_{i}(\vec{\theta})-\bar{l}(\vec{\theta})\right)\right|\lesssim R\sqrt{K}+Rt$ (35)
@@
 Using (38), we get the upper bound
-[…]
+$\begin{split}p(\mathbf{y}|\vec{c},\mathbf{X})&=\int p(\mathbf{y}|\vec{\theta},\vec{c},\mathbf{X})\frac{1}{(2\pi\gamma^{2})^{K/2}}\exp(-\|\vec{\theta}\|^{2}/2\gamma^{2})d\vec{\theta}\\ &\leq\exp(nL(\vec{c})+O(n^{1-\epsilon^{\prime}})).\end{split}$ (39)
 To get a lower bound, we use the fact that the Hessian of the population likelihood also has an lower bound
@@
 The prior density is bounded from below by $\frac{1}{(2\pi\gamma^{2})^{K/2}}\exp(-2R^{2}/\gamma^{2})$ on $B(2R)$. We therefore have
-[…]
+$\int p(\mathbf{y}|\vec{\theta},\vec{c},\mathbf{X})\frac{1}{(2\pi\gamma^{2})^{K/2}}\exp(-\|\vec{\theta}\|^{2}/2\gamma^{2})d\vec{\theta}\geq\frac{r^{K}\operatorname{Vol}(B(1))}{(2\pi\gamma^{2})^{K/2}}\exp(-2R^{2}/\gamma^{2})\exp\left(n(L(\vec{\theta}^{*})-\sqrt{K}r^{2}-O(n^{-\epsilon^{\prime}}))\right).$ (43)
 The maximum is achieved when we set $r=\frac{K^{1/4}}{(2n)^{1/2}}$. Combining (39) and (43) and taking logarithms, we get
@@
 Finally, we take a union bound over all $\vec{c}\in\mathcal{C}_{n,K}$ and $S\in\mathcal{S}_{n}$, noting that all hidden constants can be chosen to be independent of $\vec{c}$. ∎
-Appendix I Stationary Distribution Under Consistent Proposals
+Appendix H Stationary Distribution Under Consistent Proposals
 In this section, we motivate the multiple-try partial posterior Metropolis update (Algorithm 3) in more detail. Notably, we have the following proposition.
-Proposition I.1.
-Assuming equality holds in (2), the Markov chain defined by running Gibbs sampling (Algorithm 1) with SS-MH-Update or Multi-SS-MH-Update instead of MH-Update has the posterior $p(\vec{C}|\mathbf{y},\mathbf{X})$ as a stationary distribution.
+Proposition H.1.
+Suppose the LLM proposes from the conditional partial posterior distribution, i.e.
+$Q(C_{k};\vec{C}_{-k}=\vec{c}_{-k})=p(C_{k}|\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})$.
+Then the Markov chain defined by running Gibbs sampling (Algorithm 1) with SS-MH-Update or Multi-SS-MH-Update instead of MH-Update has the posterior $p(\vec{C}|\mathbf{y},\mathbf{X})$ as a stationary distribution.
 Proof.
@@
 $\begin{split}\frac{p((\check{c},\vec{c}_{-k})|\mathbf{X},\mathbf{y})Q(c;\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})}{p((c,\vec{c}_{-k})|\mathbf{y},\mathbf{X})Q(\check{c};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})}&=\frac{p((\check{c},\vec{c}_{-k})|\mathbf{X},\mathbf{y})p((c,\vec{c}_{-k})|\mathbf{X},\mathbf{y})}{p((c,\vec{c}_{-k})|\mathbf{y},\mathbf{X})p((\check{c},\vec{c}_{-k})|\mathbf{X},\mathbf{y})}\\ &=\frac{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},(\vec{c}_{-k},\check{c}),\mathbf{X})}{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})}\end{split}$ (47)
-as we wanted.
-To show the statement for Multi-SS-MH-Update, we use its equivalence, for a fixed $S$, to the modified Multiple-Try Metropolis-Hastings method described in Appendix A.
-In Appendix A, we prove that this method satisfies the detailed balance equations for the posterior.
-Since the detailed balance equations are linear in the transition probabilities, we can marginalize this over $S$ to show that the posterior is stationary with respect for Multi-SS-MH-Update, which draws $S$ uniformly at random.
-∎
+as we wanted. ∎

### v3

@@
 1 Introduction
-Although machine learning (ML) algorithms have demonstrated remarkable predictive performance, many lack the interpretability and transparency necessary for human experts to audit their accuracy, fairness, and safety [1, 2].
-This has limited their adoption in high-stakes applications such as medicine [3] and settings where regulatory agencies require algorithms to be explainable [4].
-Recent works have explored Concept Bottleneck Models (CBMs) [5, 6] as a potential solution: these methods leverage black-box algorithms to extract a small number of interpretable concepts, which are subsequently processed by a fully transparent tabular model to predict a target label.
+Although machine learning (ML) algorithms have demonstrated remarkable predictive performance, many lack the interpretability and transparency necessary for human experts to audit their accuracy, fairness, and safety [dwork2012fairness, rudin2018please].
+This has limited their adoption in high-stakes applications such as medicine [thirunavukarasu2023large] and settings where regulatory agencies require algorithms to be explainable [mesko2023imperative].
+Recent works have explored Concept Bottleneck Models (CBMs) [koh2020concept, Kim2023-vu] as a potential solution: these methods leverage black-box algorithms to extract a small number of interpretable concepts, which are subsequently processed by a fully transparent tabular model to predict a target label.
 Thus, in principle, CBMs are much safer for use in high-stakes applications because humans can audit and, if necessary, modify the extracted concepts to fix predictions.
@@
 Early CBM methods were difficult to scale because they relied on human experts to specify and annotate concepts on training data.
-Recent works have proposed using LLMs to suggest and even annotate concepts [7, 8, 9, 10], as LLMs are cheaper, often have sufficient world knowledge to hypothesize useful concepts, and can provide (relatively) high-quality concept annotations [9, 10].
-This introduces new problems such as LLM hallucinations and inconsistencies but, more critically, does not resolve the fundamental problem that standard CBMs sacrifice a significant amount of accuracy to gain interpretability [11].
-Recent proposals try to address this by encoding concepts using “soft” continuous values, such as through embedding similarity [8], rather than “hard” binary values.
-However, “soft” CBMs can leak information about the label that would otherwise not have been available [12, 13, 14, 15, 16]; this risk is particularly severe when concept extraction is jointly trained with the final label predictor [15].
+Recent works have proposed using LLMs to suggest and even annotate concepts [menon2022visual, Oikarinen2023-ee, McInerney2023-ab, Benara2024-lw], as LLMs are cheaper, often have sufficient world knowledge to hypothesize useful concepts, and can provide (relatively) high-quality concept annotations [McInerney2023-ab, Benara2024-lw].
+This introduces new problems such as LLM hallucinations and inconsistencies but, more critically, does not resolve the fundamental problem that standard CBMs sacrifice a significant amount of accuracy to gain interpretability [yuksekgonul2022post].
+Recent proposals try to address this by encoding concepts using “soft” continuous values, such as through embedding similarity [Oikarinen2023-ee], rather than “hard” binary values.
+However, “soft” CBMs can leak information about the label that would otherwise not have been available [mahinpei2021promises, margeloiu2021concept, havasi2022addressing, ragkousis2024tree, makonnen2025measuring]; this risk is particularly severe when concept extraction is jointly trained with the final label predictor [ragkousis2024tree].
 [^2] For instance, a human would fill in 0 when asked “does this bird have a gray breast” if there is no gray breast, but a “soft” CBMs may extract 0.2 if it sees a blue breast or gray head.
@@
 The aim of this work is to minimize the accuracy gap without sacrificing the interpretability of the extracted concepts.
-To this end, we revisit a known pain point of CBMs: Current training procedures require a pool of candidate concepts to be identified a priori, but ensuring this pool contains all relevant concepts is difficult, particularly in domains that are yet to be scientifically well-understood and where there are an infinite number of potentially relevant concepts [17, 10].
+To this end, we revisit a known pain point of CBMs: Current training procedures require a pool of candidate concepts to be identified a priori, but ensuring this pool contains all relevant concepts is difficult, particularly in domains that are yet to be scientifically well-understood and where there are an infinite number of potentially relevant concepts [ludan2023interpretable, Benara2024-lw].
 For instance, for the task of predicting hospital readmission, there are an infinite number of factors affecting a patient’s risk.
 Even a single concept like “smoking status” has an infinite number of refinements, such as the correlated concept of “whether a patient has quit smoking,” the broader concept of “substance use,” and more.
-As shown in prior work, when the candidate pool misses important concepts, the resulting CBMs may be inaccurate and even mislead users as to which concepts are truly relevant [18].
+As shown in prior work, when the candidate pool misses important concepts, the resulting CBMs may be inaccurate and even mislead users as to which concepts are truly relevant [Ramaswamy2023-qn].
 Rather than specifying concepts upfront, we investigate iterative refinement of concepts in a CBM with the assistance of an LLM.
 This updates a classical modeling paradigm for the LLM era: prediction models were traditionally designed through the collaboration between domain experts and data scientists, in which they iteratively refine/engineer features for simple tabular models.
-LLMs can help significantly accelerate this co-design process and scale up the number of iterations, but they are also imperfect query engines that can hallucinate, suggest bad candidate concepts, incorrectly annotate concepts, and may not even be self-consistent in their prior beliefs [19].
+LLMs can help significantly accelerate this co-design process and scale up the number of iterations, but they are also imperfect query engines that can hallucinate, suggest bad candidate concepts, incorrectly annotate concepts, and may not even be self-consistent in their prior beliefs [Wang2023-ma].
 To set up guardrails, we “wrap” the LLM within a Bayesian posterior inference procedure to explore concepts with the help of an LLM in a statistically principled manner.
@@
   • BC-LLM finds relevant concepts, even in settings where there is little to no prior knowledge: Unlike prior works that soley rely on a prespecified set of candidate concepts, BC-LLM explores a potentially infinite set of concepts data-adaptively and converge towards the true ones.
-  • BC-LLM is cost-effective: BC-LLM selectively annotates $O(KT)$ concepts, where $K$ is the number of concepts in the CBM and $T$ is the number of outer loops; we often use only $K\approx 5$ and $T\approx 5$.
-In contrast, existing methods annotate hundreds or thousands of prespecified concepts [8].
+  • BC-LLM substantially improves the interpretability-accuracy tradeoff in real-world datasets: Experiments across multiple datasets and modalities (text, images, and tabular data) show that BC-LLM selects human-interpretable concepts while outperforming comparator methods, often even black-box models.
+Moreover, when assisting a hospital’s data science team to revise an existing tabular ML model, clinicians found BC-LLM to be substantially more interpretable and actionable.
   • BC-LLM corrects LLM mistakes in a statistically rigorous manner to provide calibrated uncertainty quantification: Mistakes in the LLM’s prior are fully allowed and even expected, as they are corrected in the Metropolis accept/reject step.
@@
 The Bayesian framework also naturally protects against overfitting and overconfidence in selected concepts.
-  • BC-LLM substantially improve the interpretability-accuracy tradeoff in real-world datasets: Experiments across multiple datasets and modalities (text, images, and tabular data) show that BC-LLM selects human-interpretable concepts while outperforming comparator methods, often even black-box models.
-Moreover, when assisting a hospital’s data science team to revise an existing tabular ML model, clinicians found BC-LLM to be substantially more interpretable and actionable.
+  • BC-LLM is cost-effective: BC-LLM selectively annotates $O(KT)$ concepts, where $K$ is the number of concepts in the CBM and $T$ is the number of outer loops; we often use only $K\approx 5$ and $T\approx 5$.
+In contrast, existing methods annotate hundreds or thousands of prespecified concepts [Oikarinen2023-ee].
 Figure 1: BC-LLM is initialized by having the LLM hypothesize the top concepts based on keyphrases extracted from each observation (Step 0).
@@
 2 Related work
-Prior CBM methods train on observations with hand-annotated concepts [5, 6, 20] and, more recently, concepts annotated by an LLM/VLM [21, 9, 10].
-These methods are known to be highly sensitive to the selected set of concepts [18].
-More recent works have suggested using LLMs to prespecify the list of candidate concepts instead [7, 8, 22, 23], but this requires the LLM to have accurate prior knowledge about the supervised learning task.
-We note that many of these works also make use of “soft concepts” and therefore suffer from information leakage [12].
-To find relevant concepts in a more data-driven manner, recent works have proposed taking an iterative approach [17, 24].
+Prior CBM methods train on observations with hand-annotated concepts [koh2020concept, Kim2023-vu, Yuksekgonul2023-um] and, more recently, concepts annotated by an LLM/VLM [Patel2023-dp, McInerney2023-ab, Benara2024-lw].
+These methods are known to be highly sensitive to the selected set of concepts [Ramaswamy2023-qn].
+More recent works have suggested using LLMs to prespecify the list of candidate concepts instead [menon2022visual, Oikarinen2023-ee, Yang2023-oh, panousis2023sparse, singh2025evaluating, xu2025graph, lam2024concept], but this requires the LLM to have accurate prior knowledge about the supervised learning task.
+We note that many of these works also make use of “soft concepts” and therefore suffer from information leakage [mahinpei2021promises, margeloiu2021concept, havasi2022addressing, ragkousis2024tree, makonnen2025measuring].
+To find relevant concepts in a more data-driven manner, recent works have proposed taking an iterative approach [ludan2023interpretable, Liu2024-kq, sun2024general, kim2024constructing].
 These methods employ a boosting-like approach, where the LLM is provided a set of misclassified examples and asked to add/revise concepts to improve performance.
-However, the LLM can typically only review a few observations due to limits on context length, so the resulting concept proposals are far from optimal.
-An alternative strategy is to conduct post-hoc distillation of a black-box model into interpretable concepts [25, 26, 27, 28, 29], which can be used to learn an interpretable prediction model.
-Nevertheless, prior work has shown that the generated explanations are often not faithful to the black-box model, leading to poorer performance from the interpretable model [30].
-Another related line of work uses LLMs to describe conceptual differences between pairs of datasets  [31, 32, 33].
+However, the LLM often struggles to identify relevant patterns from the presented examples, so the resulting concept proposals are far from optimal.
+An alternative strategy is to conduct post-hoc distillation of a black-box model into interpretable concepts [Kim2018-ri, yeh2020completeness, Zhao2024-tj, Jiang2023-tx, singh2023augmenting], which can be used to learn an interpretable prediction model.
+Nevertheless, prior work has shown that the generated explanations are often not faithful to the black-box model, leading to poorer performance from the interpretable model [Doshi-Velez2017-uj].
+Another related line of work uses LLMs to describe conceptual differences between pairs of datasets  [Zhong2023-rc, zhong2022describing, dunlap2024describing] or to directly describe a black-box function using natural-language concepts [kopf2025capturing, singh2023explaining, bills2023language, hoang2024llm].
 However, existing methods are primarily suited for simple classification tasks, as they essentially learn a CBM with a single concept rather than a set of concepts.
-Finally, recent works have considered combining LLMs with Bayesian techniques [34, 35, 36].
-Some have suggested that In-Context Learning (ICL) can be viewed as LLMs conducting Bayesian inference [37, 38], though recent work has proven that ICL is not fully consistent with Bayesian inference [39].
+Finally, recent works have considered combining LLMs with Bayesian techniques [yang2024bayesian, liu2024checkpoint, liu2024large].
+Some have suggested that In-Context Learning (ICL) can be viewed as LLMs conducting Bayesian inference [Xie2021-gi, ye2024pre], though recent work has proven that ICL is not fully consistent with Bayesian inference [falck2024context].
 In contrast, this work studies how we can iteratively discover new concepts to include using LLMs using a Bayesian inference procedure, accounting for imperfections in LLM reasoning.
@@
 3.1 Review: Bayesian variable selection
-Learning CBMs shares many similarities with the more standard problem of learning sparse models in high-dimensional settings, where one must select among a finite number of features that have already been tabulated [40, 41, 42].
+Learning CBMs shares many similarities with the more standard problem of learning sparse models in high-dimensional settings, where one must select among a finite number of features that have already been tabulated [ishwaran2008random, Castillo2015-pl, biswas2022scalable].
 Here, we review the Bayesian solution for this more standard setting and then discuss how it may be extended to the CBM setting.
@@
 For each outer loop, Gibbs rotates through indices $k=1,\ldots,K$, during which it replaces the $k$th concept in the current iterate $\vec{c}$ by drawing from the posterior distribution for $C_{k}$ conditional on the other concepts $\vec{c}_{-k}$.
-When one is unable to sample from the conditional posterior, Metropolis-within-Gibbs can be used instead [43].
+When one is unable to sample from the conditional posterior, Metropolis-within-Gibbs can be used instead [Gilks1995-fw].
 Rather than immediately replacing the $k$th concept, it proposes a candidate concept $\check{c}_{k}$ given the other concepts per some distribution $Q(C_{k};\vec{C}_{-k}=\vec{c}_{-k})$ and accepts the candidate concept with probability $\alpha$.
@@
 So while Update 1 may provide valid inference, it is inefficient and costly.
-Update 2: Propose with the LLM’s “posterior.” Leveraging LLMs’ ability to perform ICL [37, 38], another idea is to include all the data when prompting the LLM to propose new concepts.
+Update 2: Propose with the LLM’s “posterior.” Leveraging LLMs’ ability to perform ICL [Xie2021-gi, ye2024pre], another idea is to include all the data when prompting the LLM to propose new concepts.
 That is, we ask the LLM itself to “conduct posterior inference” and interpret the LLM’s proposal distribution $Q(C_{k};\vec{C}_{-k}=\vec{c}_{-k})$ as sampling from a conditional posterior distribution $p(C_{k}|\vec{c}_{-k},\mathbf{y},\mathbf{X})$ for some implicitly defined concept prior.
-This approach is risky; while LLMs may be able to approximate posterior inference for small datasets, this approximation breaks down for larger sample sizes [39].
+This approach is risky; while LLMs may be able to approximate posterior inference for small datasets, this approximation breaks down for larger sample sizes [falck2024context].
 Furthermore, we can no longer override poor concepts with data as the MH acceptance ratio simplifies to 1, i.e.
@@
 We can then interpret the LLM’s proposal distribution $Q(C_{k};\vec{C}_{-k}=\vec{c}_{-k})$ as the (conditional) partial posterior distribution $p(C_{k}|\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})$.
-Then in the accept-reject step, the MH acceptance probability simplifies to the partial Bayes factor $\frac{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},(\vec{c}_{-k},\check{c}),\mathbf{X})}{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})}$ [44].
+Then in the accept-reject step, the MH acceptance probability simplifies to the partial Bayes factor $\frac{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},(\vec{c}_{-k},\check{c}),\mathbf{X})}{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})}$ [o1995fractional].
 This compares the likelihood of the held-out data $S^{c}$ with respect to the candidate and existing concepts—much like sample-splitting in frequentist settings—and thus provides an opportunity to correct inconsistencies between the LLM’s proposal and the actual posterior distribution.
@@
 8:  else return $c_{k}$
-In practice, rather than sampling a single candidate concept, we use a version of multiple-try Metropolis-Hastings (Algorithm 3) [45].
+In practice, rather than sampling a single candidate concept, we use a version of multiple-try Metropolis-Hastings (Algorithm 3) [liu2000multiple].
 Here, the LLM proposes a batch of candidates.
@@
 Assuming that the LLM’s partial posterior across all iterations is consistent with some prior distribution $p(\vec{C})$, we can prove that the Markov chain defined by this procedure has the posterior $p(\vec{c}|\mathbf{y},\mathbf{X})$ as its stationary distribution (see Proposition H.1 in the Appendix).
-Since LLMs are known to be imperfect Bayesian inference engines [39], this assumption is at best only approximately satisfied.
+Since LLMs are known to be imperfect Bayesian inference engines [falck2024context], this assumption is at best only approximately satisfied.
 Nevertheless, we can prove that the procedure still converges to the true concepts even when this assumption does not hold:
@@
 blue eyes), while keyphrases for text data can be direct quotes or summarizations (e.g.
-diabetes) [46].
+diabetes) [Papagiannopoulou2020-vl].
 Keyphrases can be thought of as rough concepts, as they do not define a function but can help an LLM come up with a formal concept question (e.g.
@@
 no joint training of concept extraction and label prediction).
-Step 4: Accept/reject step The final step is to compute the sampling probabilities for the candidate concepts and the acceptance ratio, which involve comparing the proposal probabilities for the existing and candidate concept sets $\vec{c}$ as well as their split-sample posteriors
-$p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})=\int p(\mathbf{y}_{S^{c}}|\vec{\theta},\vec{c},\mathbf{X})p(\vec{\theta}|\mathbf{y}_{S},\vec{c},\mathbf{X})d\vec{\theta}.$ (1)
-To estimate the integral (1), one approach is to sample from the posterior on $\vec{\theta}$, but this can be computationally expensive since one would need to implement posterior sampling in inner and outer loops.
-Alternatively, in the case of LR-based CBMs, one can use a Laplace-like approximation for (1) [47], which is discussed in more detail in the Appendix.
+Step 4: Accept/reject step The final step is to compute the sampling probabilities for the candidate concepts and the acceptance ratio, which involve comparing the proposal probabilities for the existing and candidate concept sets $\vec{c}$ as well as their split-sample posteriors $p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})=\int p(\mathbf{y}_{S^{c}}|\vec{\theta},\vec{c},\mathbf{X})p(\vec{\theta}|\mathbf{y}_{S},\vec{c},\mathbf{X})d\vec{\theta}.$ To estimate the integral, one approach is to sample from the posterior on $\vec{\theta}$, but this can be computationally expensive since one would need to implement posterior sampling in inner and outer loops.
+Alternatively, in the case of LR-based CBMs, one can use a Laplace-like approximation [Lewis1997-wl] (see Appendix).
 As the posterior distribution of LR parameters are asymptotically normal, this approximation becomes increasingly accurate as $n\rightarrow\infty$.
@@
 Generally speaking, a small $T$ suffices for achieving high prediction accuracy, a larger $T$ (e.g., 10) is helpful for quantifying the uncertainty of relevant concepts.
-Critically, running BC-LLM is significantly more cost-effective than standard CBMs, which require $O(nW)$ queries where $W$ is the number of pre-specified concepts and must be large (typically hundreds or thousands) to avoid missing relevant concepts [8].
-In our experiments, BC-LLM was quick to run, as one iteration takes less than a minute for text data and a few minutes for image data on a normal laptop.
+Critically, running BC-LLM is significantly more cost-effective than standard CBMs, which require $O(nW)$ queries where $W$ is the number of pre-specified concepts and must be large (typically hundreds or thousands) to avoid missing relevant concepts [Oikarinen2023-ee].
+In our experiments, BC-LLM was quick to run, as one iteration takes less than a few minutes on a normal laptop.
+[check timings a bit more carefully]
 4 Experiments
 We evaluated BC-LLM in across three domains and modalities: classifying birds in images (Section 4.1), simulated outcomes from clinical notes (Section 4.2), and readmission risk in real-world clinical data (Section 4.3).
-The experiments below used GPT-4o-mini [48] and Section 4.3 used a protected health information-compliant version of GPT-4o for real-world clinical notes.
+The experiments below used GPT-4o-mini [openai2023gpt4] and Section 4.3 used a protected health information-compliant version of GPT-4o for real-world clinical notes.
 The Appendix includes results using other LLMs, implementation details, and experimental settings.
 Baselines.
-We compared against CBMs fit using the standard approach where humans select and annotate concepts (Human+CBM) [5], CBMs where an LLM brainstorms concepts after seeing which keyphrases are most associated with the label but without any iteration (LLM+CBM) [49, 10], and CBMs fit using a boosting algorithm that iteratively adds concepts by having an LLM analyze observations with large residuals (Boosting LLM+CBM) [17, 24].
+We compared against CBMs fit using the standard approach where humans select and annotate concepts (Human+CBM) [koh2020concept], CBMs where an LLM brainstorms concepts after seeing which keyphrases are most associated with the label but without any iteration (LLM+CBM) [pham2023topicgpt, Benara2024-lw], and CBMs fit using a boosting algorithm that iteratively adds concepts by having an LLM analyze observations with large residuals (Boosting LLM+CBM) [ludan2023interpretable, Liu2024-kq].
 We include black-box and semi-interpretable models to assess the tradeoff between interpretability and accuracy.
-For the bird-classification dataset, we also compare to a CBM that uses LLM-suggested concepts rather than keyword-based concepts; we use the 370 LLM-suggested concepts from [8] and extract the values of these concepts with an LLM (LLM+CBM (No keyphrases)).
+For the bird-classification dataset, we also compare to a CBM that uses LLM-suggested concepts rather than keyword-based concepts; we use the 370 LLM-suggested concepts from [Oikarinen2023-ee] and extract the values of these concepts with an LLM (LLM+CBM (No keyphrases)).
 Evaluation metrics.
@@
 4.1 CBMs for classifying bird images
-We first evaluate how well BC-LLM can learn concepts that differentiate between bird species within the same family in the standard CUB-birds image dataset [50].
+We first evaluate how well BC-LLM can learn concepts that differentiate between bird species within the same family in the standard CUB-birds image dataset [welinder2010caltech].
 Grouping the 200 bird species into their respective families, this resulted in 37 prediction tasks.
@@
 The number of concepts $K$ learned for each task was set to the number of classes, but no smaller than 4 and no greater than 10.
-For the black-box comparator, we fine-tuned the last layer of ResNet50 pre-trained on ImageNetV2 [51, 52].
+For the black-box comparator, we fine-tuned the last layer of ResNet50 pre-trained on ImageNetV2 [He2016-rc, recht2019imagenet].
 Human+CBM was trained on the 312 human-annotated features available in the CUB-birds dataset.
@@
 To objectively measure how well BC-LLM can recover true concepts, we conducted a simulation study where the true concepts are known.
-To assess BC-LLM in settings where the LLM has no prior knowledge about the target, we simulated $Y$ using a LR model with five Social Determinants of Health as inputs, which were annotated from real-world patient notes in MIMIC-IV [53, 54]:
+To assess BC-LLM in settings where the LLM has no prior knowledge about the target, we simulated $Y$ using a LR model with five Social Determinants of Health as inputs, which were annotated from real-world patient notes in MIMIC-IV [johnson2023mimic, Ahsan2021-hx]:
 1. Does the note mention the patient consuming alcohol in the present or the past?
@@
 Here we demonstrate how BC-LLM can help ML developers integrate different data modalities to improve model performance.
-The data science team at hospital [Anonymous] has both tabular data and unstructured notes from the electronic health record (EHR).
+The data science team at Zuckerberg San Francisco General Hospital has both tabular data and unstructured notes from the electronic health record (EHR).
 The team has so far fit a model using the tabular data to predict readmission risk for heart failure patients and wants to assess if clinical notes contain additional concepts that are predictive.
@@
 The use of LLMs incurs significant computational cost and corresponding environmental impacts; however, BC-LLM is much more computationally efficient than comparable LLM-based CBMs, potentially reducing overall environmental impacts.
-Acknowledgements
+Acknowledgments
 This work was funded through a Patient-Centered Outcomes Research Institute® (PCORI®) Award (ME-2022C1-25619).
@@
 The authors thank the UCSF AI Tiger Team, Academic Research Services, Research Information Technology, and the Chancellor’s Task Force for Generative AI for their technical support related to the use of Versa API gateway.
-The authors thank Patrick Vossler, Yifan Mai, Seth Goldman, Andrew Bishara, Jim Marks, Julian Hong, Aaron Kornblith, Nicholas Petrick, and Gene Pennello for their invaluable feedback on the work.
+The authors thank Patrick Vossler, Yifan Mai, Seth Goldman, Andrew Bishara, James Marks, Julian Hong, Aaron Kornblith, Nicholas Petrick, and Gene Pennello for their invaluable feedback on the work.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: We carefully ensure the claims made in our abstract and introduction and clearly defined and justified through our methodology (Section 3) and experimental results (Section 4).
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: The discussion section mentions that the current method is primarily suitable for learning smaller, highly interpretable CBMs. Future work can further scale up the procedure.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate ”Limitations” section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: Assumptions are stated in Section 3 and full proofs for theorems can be found in the Appendix
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: Experiment details are found in Section 4 and reproducible code is found at our github.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24.
+Justification: A link to the source code is provided, with specification of all data, dependencies, instructions to reproduce results.
+The only experiment that cannot be reproduced is Section 4.3 due to the use of PHI data.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: Experimental details can be found in the Appendix and Section 4
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: Confidence intervals for performance measures are provided in all figures/tables in Section 4
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer ”Yes” if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: The main manuscript discusses compute time and resources.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: All ethical guidelines were followed.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: Broader impacts are discussed in the Discussion section at the end of the paper.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A] .
+  54. Justification: Our research does not release any new data or models
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: All assests in the paper are cited in the references section.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: The new method introduced is detailed in Section 3 and is implemented in our repository with instructions for use.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [Yes]
+  69. Justification: Instructions given to clinicians for rating concepts are provided in the Appendix.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [Yes]
+  74. Justification: IRB approval was obtained for analyzing real-world clinical notes from the hospital, as mentioned in the Appendix.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [Yes]
+  79. Justification: Our method detailed in Section 3 uses LLMs to propose and refine concepts for CBMs.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
 Appendix
@@
 In this section, we provide intuition for the Multiple-Try Metropolis-Hastings partial posterior method (Algorithm 3).
-To draw connections with the original Multiple-Try Metropolis-Hastings method [45], we first work in a completely abstract setting.
+To draw connections with the original Multiple-Try Metropolis-Hastings method [liu2000multiple], we first work in a completely abstract setting.
 Given a state space $\Omega$, stationary distribution $\pi$, and proposal transition kernel $T(y;x)$, define weights $w(x,y)=\pi(x)T(y;x)$.
@@
 Classical version.
-We briefly describe [45]’s original method. Suppose the current state is $x$.
+We briefly describe [liu2000multiple]’s original method. Suppose the current state is $x$.
   1. 1.
@@
     $\beta(x,y;\mathbf{z},\mathbf{z}^{*})=\min\left\{\frac{\sum_{i=1}^{M}w(z_{i},x)}{\sum_{i=1}^{M}w(z_{i}^{*},y)},1\right\}.$ (2)
-One can show that this Markov chain satisfies the detailed balance equations when marginalizing out the intermediate states $z_{1},z_{2},\ldots,z_{M},z_{1}^{*},\ldots,z_{M}$ (see Theorem 1 in [45]).
+One can show that this Markov chain satisfies the detailed balance equations when marginalizing out the intermediate states $z_{1},z_{2},\ldots,z_{M},z_{1}^{*},\ldots,z_{M}$ (see Theorem 1 in [liu2000multiple]).
 We would like, however, to minimize the number of proposals necessary, so it is natural to ask whether the algorithm can be modified such that the proposals needed for the backward transitions, $z_{1}^{*},\ldots,z_{M-1}^{*}$, can be omitted.
@@
 clinical note: $\langle$ note $\rangle$
-Note that in cases where the LLM was unsure about the concept’s value, it was allowed to return a probability, which prior works have found to be helpful [6].
+Note that in cases where the LLM was unsure about the concept’s value, it was allowed to return a probability, which prior works have found to be helpful [Kim2023-vu].
 In general, the LLM outputted 1’s and 0’s.
@@
 D.1 MIMIC
-We evaluate BC-LLM against baseline methods using Cohere’s Command-R LLM using the same settings as discussed in Section 4.2. Performance can be found in Figure 5.
-Figure 5: Comparison of BC-LLM and existing methods in terms of performance and recovery of true concepts. Error bars indicate the 95% CI.
+To assess the robustness of BC-LLM, we include the following variations.
+First, we test the robustness of BC-LLM to prompt phrasing.
+In particular, we feed in a prompt that is more vague than that used in the main manuscript during Step 0 of BC-LLM by replacing the original sentence “Output a list of descriptors that summarizes the patient case (such as aspects on demographics, diagnoses, SDOH, etc).“ with “Output a list of descriptors that summarizes the patient case.“ Second, we test the robustness of the results to the choice of LLM by rerunning the same experiment using Cohere’s Command-R LLM, rather than GPT-4o-mini.
+Results are shown in Figure 5.
+First, we see that BC-LLM is quite robust to prompt phrasing, as the results with this simpler prompt are similar to the more detailed prompt.
+Second, we see that performance of all the methods—BC-LLM and the comparator methods—are lower using Command-R.
+Nevertheless, we see that the relative performance between the methods is the same, with BC-LLM performing the best.
+[UPDATE FIGURE – BC-LLM should also be labelled with cohere vs prompt ablation]
+Figure 5: Additional results running BC-LLM and comparator methods on the MIMIC dataset, evaluated in terms of performance and recovery of true concepts. Error bars indicate the 95% CI.
 D.2 CUB-Birds dataset
-We evaluate BC-LLM against baseline methods using Anthropic’s Claude 3.5 Haiku using the same settings as discussed in Section 4.1. Performance can be found in Table 2.
+Table 2 reports results from two additional sets of experiments on the CUB-Birds dataset.
+First, to evaluate robustness of BC-LLM to the choice of LLM, we report the performance of BC-LLM and comparator methods when Anthropic’s Claude 3.5 Haiku is used instead, with all other settings the same as that discussed in Section 4.1.
+Second, to facilitate comparisons to other CBM results on the CUB-Birds dataset, we evaluate BC-LLM on the dataset’s original task of classifying all 200 bird species.
+Nevertheless, we emphasize that this is not the intended use case of BC-LLM, as its aim is to learn fully interpretable CBM methods with $K\leq 20$.
+Due to computation time, we run BC-LLM for one epoch rather than five.
+[technically it is half an epoch…]
 Method Accuracy ($\uparrow$) AUC ($\uparrow$) Brier ($\downarrow$)
+Using Claude 3.5 Haiku
 BC-LLM 0.665 (0.593, 0.736) 0.849 (0.811, 0.888) 0.443 (0.367, 0.519)
@@
 Boosting LLM+CBM 0.640 (0.568, 0.712) 0.811 (0.769, 0.852) 0.552 (0.455, 0.649)
-Table 2: Results for the CUB-Birds dataset using Claude 3.5 Haiku
+Classifying between all 200 bird species
+BC-LLM 0.562 (0.487, 0.644) 0.598 (0.523, 0.665)
+LLM+CBM
+Boosting LLM+CBM 0.005 (0.003, 0.007) 0.500 (0.500, 0.500) 1.012 (1.011, 1.012)
+Table 2: Additional results for the CUB-Birds dataset
+D.3 Other image datasets
+To evaluate BC-LLM on other real-world image datasets, Table 3 includes two additional imaging datasets:
+  • Functional Map of the World (fMoW) [Christie2018-dm]: RGB satellite images classified into 62 building/land use categories.
+Models were trained on satellite images from the US, with 100 images sampled from each category.
+CBMs were trained to have 60 concepts.
+Models were evaluated in terms of in-distribution performance (i.e.
+images from the USA) as well as out-of-distribution (OOD) performance when evaluated on satellite images from China.
+  • Imagenette [imagenette-xz]: 10 classes from ImageNet. CBMs were trained to have 10 concepts.
+Method Accuracy ($\uparrow$) AUC ($\uparrow$) Brier ($\downarrow$)
+fMoW: USA images
+BC-LLM 0.357 (0.340, 0.375) 0.904 (0.898, 0.910) 0.780 (0.766, 0.792)
+LLM+CBM 0.311 (0.295, 0.327) 0.878 (0.871, 0.885) 0.892 (0.872, 0.914)
+Boosting LLM+CBM 0.118 (0.105, 0.130) 0.757 (0.748, 0.764) 1.029 (1.018, 1.039)
+fMoW: OOD-China images
+BC-LLM 0.265 (0.246, 0.285) [???] 0.853 (0.839, 0.868)
+LLM+CBM 0.223 (0.205, 0.242) [???] 1.023 (1.00, 1.045)
+Boosting LLM+CBM 0.084 (0.072, 0.098) [??] 1.065 (1.054, 1.077)
+Imagenette
+BC-LLM 0.987 (0.980, 0.993) 0.999 (0.997, 1.000) 0.072 (0.062, 0.085)
+LLM+CBM 0.902 (0.883, 0.919) 0.988 (0.986, 0.990) 0.125 (0.107, 0.142)
+Boosting LLM+CBM 0.639 (0.610, 0.670) 0.899 (0.888, 0.909) 0.448 (0.422, 0.479)
+Table 3: Results on additional imaging datasets
 Appendix E Experiment details
@@
 For the black-box comparator, we used a ResNet50 model pre-trained on ImageNetV2, which ensures ResNet’s training set does not overlap with the CUB-Birds dataset.
-E.4 Clinical notes and tabular data from Hospital Anonymous
+E.4 Clinical notes and tabular data from the Zuckerberg San Francisco General Hospital
 Data in this experiment contained PHI, for which IRB approval was obtained.
@@
 Remark F.2.
-In the current implementation of our procedure, we use a slightly different method to perform Laplace approximation, which instead approximates the integral in (1).
+In the current implementation of our procedure, we use a slightly different method to perform Laplace approximation, which instead approximates the integral in (LABEL:eq:conditional_lkhd_decomposition).
 Appendix G Proof of Theorem 3.1
@@
 Furthermore, it is finite according to Assumption G.1(vi).
-As such, $(\vec{c}^{~t})_{t=1}^{\infty}$ has a unique stationary distribution $\pi_{\mathcal{A}}$ supported on $\mathcal{A}$ [55].
+As such, $(\vec{c}^{~t})_{t=1}^{\infty}$ has a unique stationary distribution $\pi_{\mathcal{A}}$ supported on $\mathcal{A}$ [LevinPeresWilmer2006].
 Indeed, all stationary distributions of $(\vec{c}^{~t})_{t=1}^{\infty}$ are convex combinations of such distributions.
@@
 $\left\|n^{-1/2}\sum_{i=1}^{n}\left(l_{i}(\vec{\theta})-\bar{l}(\vec{\theta})\right)-n^{-1/2}\sum_{i=1}^{n}\left(l_{i}(\vec{\theta}^{\prime})-\bar{l}(\vec{\theta}^{\prime})\right)\right\|_{\psi_{2}}\lesssim K\|\vec{\theta}-\vec{\theta}^{\prime}\|_{2}.$ (33)
-We can thus use Talagrand’s comparison inequality [56] to get, for any $t>0$, that
+We can thus use Talagrand’s comparison inequality [vershynin2018high] to get, for any $t>0$, that
 $n^{-1/2}\sup_{\vec{\theta}\in B(2R)}\left|\sum_{i=1}^{n}\left(l_{i}(\vec{\theta})-\bar{l}(\vec{\theta})\right)\right|\lesssim R\sqrt{K}+Rt$ (34)
@@
 $\begin{split}\frac{p((\check{c},\vec{c}_{-k})|\mathbf{X},\mathbf{y})Q(c;\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})}{p((c,\vec{c}_{-k})|\mathbf{y},\mathbf{X})Q(\check{c};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})}&=\frac{p((\check{c},\vec{c}_{-k})|\mathbf{X},\mathbf{y})p((c,\vec{c}_{-k})|\mathbf{X},\mathbf{y})}{p((c,\vec{c}_{-k})|\mathbf{y},\mathbf{X})p((\check{c},\vec{c}_{-k})|\mathbf{X},\mathbf{y})}\\ &=\frac{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},(\vec{c}_{-k},\check{c}),\mathbf{X})}{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})}\end{split}$ (46)
-as we wanted. ∎
+as we wanted.
+To show the statement for Multi-SS-MH-Update, we use its equivalence, for a fixed $S$, to the modified Multiple-Try Metropolis-Hastings method described in Appendix A.
+In Appendix A, we prove that this method satisfies the detailed balance equations for the posterior.
+Since the detailed balance equations are linear in the transition probabilities, we can marginalize this over $S$ to show that the posterior is stationary with respect for Multi-SS-MH-Update, which draws $S$ uniformly at random.
+∎